### 19 Mar 2026 v7

# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

Abstract
We formalize the orientation boundary for first-order step-duplicating recursors. At the schema, no measure in twelve formalized direct classes can strictly orient the duplicating step uniformly. The classes range from additive and transparent-compositional (excluded unconditionally) through affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, max-plus, and tracked vector / pair families (excluded under stated pump or growth hypotheses). Two cross-cutting meta-results tighten the stack: a scalar-projection meta-theorem for the matrix-side barriers and a symbolic variable-condition barrier yielding a KBO-style impossibility corollary for the concrete witness system studied later. To our knowledge, this is the first mechanized direct impossibility result for named termination-measure classes on a fixed terminating system. All internal proofs are formalized in Lean 4.
The boundary is explanatory, not only negative. An escape trichotomy for the explicit direct universe, extended further by scalar-projection representability on the matrix side, a transparency-essentiality theorem with a concrete witness, a dependency-pair escape characterization, a generalized polynomial boundary theorem isolating failure of frozen base dominance, computable barrier-witness extractors, ablations, and SCC synchronization theorems classify how successful methods cross the boundary. Every method that orients the duplicating step does so by importing structural assumptions outside the formalized barrier classes.
The concrete witness calculus KO7 carries a two-layer formal development. For the guarded fragment we internally certify strong normalization, confluence, a certified normalizer, decidable reachability to safe normal-form targets, an exact DM-component ordinal calibration at $\omega^{\omega}$, and a triple-lex certificate image below $\omega^{\omega}\cdot 2$. For the full unguarded system we separately prove root termination by a nonlinear polynomial witness and by a KO7-specialized MPO, lift the nonlinear witness to context-closed strong normalization with an explicit derivation-length bound, and connect the Lean development to TTT2/CeTA through a checked TPDB export and verifier bridge.
The Lean 4 artifact is available at https://github.com/MosesRahnama/OperatorKO7.
1 Introduction
Consider a first-order rewrite rule of the form
$\mathrm{recur}(b,\,s,\,\mathrm{succ}(n))\;\to\;\mathrm{wrap}(s,\,\mathrm{recur}(b,\,s,\,n))$
where the step argument $s$ appears once on the left and twice on the right. Any direct aggregative termination measure that aggregates subterm contributions and respects the wrapper constructor must account for both copies of $s$ on the right-hand side. Once the measured value of $s$ exceeds the fixed successor-side contribution available in the chosen class, the extra copy prevents a uniform strict decrease. Modular methods (dependency pairs with subterm criterion) escape by projecting to the recursion counter and discarding the duplicated mass.
The central question is: which classes of termination measures can orient this rule, and which cannot? We call the dividing line between these classes the orientation boundary. On one side sit measure classes that provably fail; on the other, methods that succeed by importing structural assumptions outside those classes.
This paper formalizes both sides of this boundary. We prove impossibility theorems at the schema for twelve axiomatized direct measure classes (enumerated below), classify the escape mechanisms, and validate the boundary on a concrete system, KO7, carrying a full certification chain. These twelve classes are not ad hoc: they are the direct scalar and tracked-tuple fragments of the broader tuple-interpretation taxonomy discussed later in Section 4.
Prior work
The observation that additive and KBO-style direct measures cannot orient duplicating rules is classical: KBO cannot orient any duplicating rule because of the variable condition [3], and the dependency pair framework was introduced to handle systems where simplification orderings fail [1]. Middeldorp and Zantema [18] established the hierarchy between simple termination and full termination. Our contribution is to axiomatize a broader transparency-constrained compositional class together with a separate affine/linear constructor-local class, and then prove barrier theorems, an escape characterization, a transparency-essentiality theorem, and a mechanized ablation for those classes. To our knowledge, no prior work formalizes impossibility theorems for named termination-measure classes on a fixed terminating system: every member of a specified class fails on that system. Mitterwallner, Middeldorp, and Thiemann [19, 27] formalized in Isabelle/HOL the undecidability of orientation by linear polynomial interpretations and by KBO with subterm coefficients. That is a result about the decision problem across all inputs. Their work does not produce standalone impossibility theorems for named systems; ours does. The broader formalization ecosystem (IsaFoR/CeTA, CoLoR) certifies positive termination proofs but not impossibility theorems for measure classes.
The closest neighboring lines of work therefore answer different questions. The LICS/AFP undecidability line asks whether a method class is algorithmically decidable across all inputs; certificate ecosystems such as IsaFoR/CeTA and CoLoR ask whether a successful external termination argument can be reconstructed and checked. Our claim is narrower and different: for one fixed terminating system, we prove object-level impossibility theorems for explicit direct measure families, explain the escape routes, and package constructive witnesses and classifiers around that boundary. This difference is not only mathematical but architectural. The artifact is not another generic certificate checker. It internalizes a negative boundary theory for a fixed terminating system and then reconnects that Lean development back to standard TRS tooling through a checked TPDB export and verifier bridge.
Correspondingly, the paper does not claim that KO7 defeats termination methods in general. The no-go theorems concern twelve direct measure families and an adjacent direct symbolic barrier. Modular methods and fixed-signature path-order arguments remain on the successful side of the boundary, and the artifact proves exactly that through the DP escape, the nonlinear full-step witness $W$, and the KO7-specialized MPO.
Contributions
All results below are machine-checked in Lean 4 unless otherwise noted.
A. The orientation barrier (primary contribution). The main theorem package is a twelve-class direct boundary for the duplicating step, with the latter ten classes blocked under their stated pump or growth hypotheses. The classes fall into three blocks: additive and transparency-constrained compositional measures; affine and nonlinear scalar extensions (restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus); and four tracked vector/pair families (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). Two structural results sit above that stack: a scalar-projection meta-theorem (Theorem 4.13) unifying the projection-based matrix arguments, and a symbolic variable-condition barrier (Theorem 4.14) together with an explicit KO7 KBO corollary (Corollary 4.15) for direct KBO-style comparators. Meta/PumpedBarrierClasses.lean supplies unconditional strengthened-subclass corollaries for the affine, restricted-quadratic, bounded cross-term quadratic, tracked-primary pair, bounded multilinear, max-plus, and projection-based componentwise matrix families. The affine contradiction cutoff is also shown exact for a canonical family in Meta/AffineThresholdSharpness.lean. Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean provide concrete violating triples for the additive and transparent-compositional classes, a packaged affine contradiction once a pump term is supplied, and additional certified extractors for the restricted-quadratic, max-plus, and projected matrix-side pumped subclasses.
B. Escape characterization and precision. A DP escape clause (Theorem 4.21), a transparency-essentiality theorem with a concrete witness (Theorem 4.19), an explicit DP base-order boundary note (Proposition 4.22), a generalized degree-bounded polynomial barrier with a necessary base-dominance failure condition (Theorem 4.6), and an escape trichotomy over an explicit direct universe, extended on the matrix side by projected representability (Theorem 4.30), turn the no-go side into an explicit escape theory. Two ablations and two SCC synchronization theorems sharpen the boundary: removing step-duplication dissolves the barrier; adding a wrapper-collapse rule does not; delayed and multiplicity-preserving SCC duplication both inherit the obstruction (§4.6).
C. Certified implementation and artifact engineering. The artifact does more than state theorems. For the guarded SafeStep fragment of KO7 it certifies strong normalization, root-step confluence, a certified normalizer, decidable reachability to safe normal-form targets, partial context-closure strong normalization and confluence, an exact DM-component ordinal calibration at $\omega^{\omega}$, and a triple-lex upper-bound calibration below $\omega^{\omega}\!\cdot\!2$ (Theorems 5.95.11). For the full unguarded Step relation it formalizes two independent root-step termination proofs (nonlinear polynomial and KO7-specialized MPO), a Lean proof of full context-closed strong normalization by lifting the nonlinear witness through contexts, and a concrete derivation-length bound for contextual reductions. The support layer is executable: Meta/BarrierClass_Classifier.lean, Meta/BarrierWitness.lean, and Meta/SynthesisOracle.lean expose a classifier / witness / oracle pipeline for direct-orientation attempts, while Meta/TPDB_Export.lean and lakeexeverifyTpdbExport provide a checked bridge to the external TTT2/CeTA artifact. TTT2 independently reproduces the context-closed termination result; CeTA certifies the external proof.
Organization
[…]
2 Background
Standard abstract reduction and term rewriting notions apply [3, 25]. A term is in normal form if no rule applies. Strong normalization (SN) means that no infinite reduction exists. Confluence means that for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma: SN + local confluence $\Rightarrow$ confluence [21], yielding unique normal forms.
[…]
3.1 The step-duplicating schema
Definition 3.1 (Step-duplicating schema).
A step-duplicating schema consists of a term carrier $T$ with designated operations
$\mathrm{base}\in T,\qquad\mathrm{succ}:T\to T,\qquad\mathrm{wrap}:T\times T\to T,\qquad\mathrm{recur}:T^{3}\to T,$
and a step relation containing the duplicating rule
$\mathrm{recur}(b,s,\mathrm{succ}(n))\to\mathrm{wrap}(s,\mathrm{recur}(b,s,n)).$
Definition 3.2 (Direct compositional measures).
Two compositional axiom systems, both mapping terms to $\mathbb{N}$:
  1. 1.
    Tier 1 (additive). A fixed weight per constructor; the measure of a compound term is the constructor’s weight plus the sum of subterm measures. Hypothesis: wrap contributes at least 1.
  2. 2.
    Tier 2 (abstract, with wrapper-subterm axioms). Each constructor has an arbitrary combining function on subterm measures, subject to:
    $c_{\mathrm{wrap}}(x,y)>x\qquad\text{and}\qquad c_{\mathrm{wrap}}(x,y)>y.$
    No growth property is assumed for $\mathrm{succ}$. The Tier 2 impossibility additionally requires $\mathrm{succ}$-transparency at $\mathrm{base}$: $c_{\mathrm{succ}}(c_{\mathrm{base}})=c_{\mathrm{base}}$.
A measure orients a rule instance if the measured value strictly decreases across that instance.
Separately, Theorem 4.2 isolates an affine/linear constructor-local fragment:
$M(\mathrm{succ}(t))=\alpha_{s}+\beta_{s}\,M(t),\qquad M(\mathrm{wrap}(x,y))=\alpha_{w}+\beta_{w}\,M(x)+\gamma_{w}\,M(y),$
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n),$
with no cross terms. Its barrier uses an explicit unbounded-pump hypothesis rather than only the local axioms of Definition 3.2.
The schema and barrier theorems are formulated for a single self-recursive duplicating rule. Mutually recursive step-duplicating systems, where two or more functions duplicate each other’s arguments across distinct rules, are not analyzed here and may require separate barrier arguments.
Remark 3.3 (Classic duplicating TRS as schema instance).
The schema also captures standard parameterized duplicating rules from first-order rewriting. For example, the rule
$f(x,\,s(y))\;\to\;g(x,\,f(x,y))$
is an exact instance: take $\mathrm{base}=b$ as a dummy parameter, let $\mathrm{succ}=s$ act on the second argument, set $\mathrm{wrap}(a,r)=g(a,r)$, and define $\mathrm{recur}(b,a,n)=f(a,n)$. Then the schema step
$\mathrm{recur}(b,a,\mathrm{succ}\,n)\;\to\;\mathrm{wrap}(a,\mathrm{recur}(b,a,n))$
becomes precisely the displayed rule. The Tier 1 and Tier 2 barrier theorems therefore apply directly to this familiar duplicating shape. Thus the formalized barriers are not artifacts of KO7 but of the underlying self-recursive duplication pattern.
3.2 KO7
KO7 is a first-order ground term calculus with 7 constructors and 8 rules.
Kernel signature
$t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
Table 1: KO7’s 8 kernel rules. R_rec_succ (*) duplicates the step argument $s$.
Rule Head Arity Shape Dup?
R_merge_void_left merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R_merge_void_right merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R_merge_cancel merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R_int_delta integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R_eq_refl eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
In full Step, R_eq_refl and R_eq_diff overlap at $\mathrm{eqW}\,a\,a$. The SafeStep fragment adds per-rule guards: $\delta$-flag guards on merge and rec-zero rules serve termination; $\kappa^{M}(a)=\varnothing$ on R_eq_refl and $a\neq b$ on R_eq_diff serve confluence.
Relations
Relation Scope SN Confluence Normalizer
Step Full, root Yes (mech.) No (witness) n/a
SafeStep Guarded, root Yes (mech.) Yes, root ARS Yes (mech.)
SafeStepCtx Partial ctx. Yes (mech.) Yes; exact Newman equivalence n/a
The primary motivation for introducing SafeStep is confluence rather than root termination. The chosen guards are also coordinated with the computable termination certificate: deltaFlag and $\kappa^{M}$ appear both in the guarded rules and in the triple-lexicographic decrease argument. Later sections prove that the full root relation Step is terminating, but it already fails local confluence at the concrete term $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$:
$\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}\to\mathrm{void}\qquad\text{and}\qquad\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}\to\mathrm{integrate}(\mathrm{merge}\;\mathrm{void}\;\mathrm{void}).$
Both targets are full-Step normal forms, so they do not join. The guarded relation SafeStep is introduced to block this overlap and recover unique normal forms for the root ARS.
The constructor app has no reduction rules. When rec_succ fires, the duplicated $s$ inside $\mathrm{app}(s,\mathrm{rec}\Delta(b,s,n))$ cannot feed back as a recursion counter: no rule has app at the root of its LHS. The duplicated payload does not itself become a new root recursive control position.
[…]
The impossibility and DP escape depend only on four constructors and two rules:
$\begin{array}[]{rcl}t&::=&\mathrm{void}\mid\delta\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\Delta\,t\,t\,t\\[4.0pt] \mathrm{rec}\Delta\,b\,s\,\mathrm{void}&\to&b\\ \mathrm{rec}\Delta\,b\,s\,(\delta\,n)&\to&\mathrm{app}\,s\,(\mathrm{rec}\Delta\,b\,s\,n)\end{array}$
[…]
4 The orientation barrier
The orientation barrier arises under four structural conditions, all satisfied by KO7:
  1. 1.
    No binding operators. First-order rewriting; no $\lambda$-abstraction, no free variables. Semantic methods (reducibility candidates, logical relations) rely on a typed or higher-order structure to supply the well-founded induction; KO7’s untyped first-order setting does not provide this structure.
  2. 2.
    Tree semantics (no sharing). All state is in the term tree; no external tape, graph-reduction sharing [3], or shared pointers. The step argument must be syntactically duplicated. This condition is part of the theorem scope: graph rewriting, shared environments, or explicit let-style sharing can change the duplication accounting and therefore fall outside the present barrier argument.
  3. 3.
    No axioms inside the object language. No equational theories or imported well-ordering principles are built into the rewrite language itself. (The rule $\mathrm{merge}(t,t)\to t$ is a directed reduction rule, not an algebraic identity.) The wrapper-collapse ablation in Section 4.6 should therefore be read only as a syntactic surrogate for object-level collapse, not as rewriting modulo an ambient equational theory.
  4. 4.
    Unrestricted step-duplicating recursor. A rule of the form $\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n))$ with $s$ unrestricted. Our ablation (§4.6) shows that removing duplication dissolves the barrier.
Simple typing by itself is not an escape from the barrier. The artifact includes a simply-typed first-order recursor fragment with separate result, step, and counter sorts, and proves that the additive barrier still survives there. It also proves that the affine barrier survives under an explicit typed step-pump hypothesis, namely, that the step sort still contains arbitrarily large closed terms. The point is structural: typing can block the old untyped pump patterns, but it does not dissolve the obstruction once a typed unbounded step family remains available. Machine-checked in Meta/TypedBarrierSurvival.lean as no_additive_orients_typed_recSucc and no_affine_orients_typed_recSucc_of_stepPump.
Under these conditions, the barrier stack has a clear internal structure rather than a flat catalog. Theorems 4.1 and 4.2 establish the baseline: additive, transparency-constrained compositional, and affine constructor-local measures already fail. Theorems 4.3 to 4.7 then widen the scalar side in stages, asking how much extra nonlinear growth or alternative aggregation can be added before the freeze-and-pump argument breaks; each theorem identifies the surviving obstruction. Theorems 4.9 to 4.13 form the vector / pair block, ending with a single scalar-projection meta-theorem that states the common matrix-side proof shape once. Theorem 4.14 and Corollary 4.15 close the direct symbolic side. This organization matters because the escape results later mirror it: successful methods work only by importing structure outside one of these blocks. Table 2 summarizes the resulting boundary.
This twelve-class stack should also be read against the modern interpretation-method taxonomy rather than as an ad hoc list. Tuple interpretations provide a unifying framework subsuming polynomial, matrix, arctic, and related interpretation families [29]. Most of our scalar and tracked vector / pair classes are direct syntactic fragments of that broader framework: the additive, affine, quadratic, multilinear, generalized bounded-polynomial, and max-plus barriers are scalar or one-coordinate fragments, while the fixed-dimension, lexicographic, balanced mixed-coordinate, and weighted scalar-projection theorems are tracked-tuple fragments with restricted comparison interfaces. The extra pump and dominance hypotheses in our statements are exactly the explicit growth conditions under which those direct fragments can be tested against the step-duplicating schema. The escape side then has a clean interpretation: DP projection and the full-step witnesses do not refute tuple-interpretation methods wholesale; they succeed by leaving this direct boundary through projection or a stronger global proof shape.
Table 2: Orientation boundary map for KO7’s step-duplicating recursor.
Bucket KO7 status
Provably blocked classes Twelve direct families are ruled out by theorem: additive and transparent-compositional measures; affine and nonlinear scalar extensions (restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus); and tracked vector/pair extensions (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). The additive and transparent-compositional classes are excluded unconditionally; the remaining ten are excluded under their stated pump or growth hypotheses (Theorems 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.7, 4.9, 4.10, 4.11, 4.12).
[…]
Transformed / projection escape Dependency-pair projection succeeds by discarding duplicated payload and thereby violating wrapper-subterm sensitivity (Theorem 4.21).
Direct out-of-class global orienters The full root relation is oriented internally by the nonlinear polynomial witness and the KO7-specialized MPO, both outside the formalized direct barrier classes (Propositions 4.26, 4.27).
4.1 Barrier at the schema
Theorem 4.1 (Barrier at the schema).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema (Definition 3.1). Then:
  1. 1.
    No Tier 1 measure with positive wrap-weight strictly orients the duplicating step uniformly.
  2. 2.
    No Tier 2 measure with the wrap-subterm axioms and successor transparency at $\mathrm{base}$ strictly orients the duplicating step uniformly.
  3. 3.
    Neither class globally orients any relation containing the duplicating step.
Machine-checked in Meta/StepDuplicatingSchema.lean. Computable certificate extractors in Meta/BarrierWitness.lean (additive_witness, compositional_witness, affine_witness) return a concrete violating triple $(b,s,n)$ with a bundled proof of orientation failure: directly from the measure data in the additive and transparent-compositional cases, and from the measure together with a supplied pump term above the required bound in the affine case.
Proof sketch. For Tier 1, set $b=n=\mathrm{base}$ and $s=\mathrm{wrapIter}(w_{\mathrm{succ}})$ (iterated wrapper). By additivity, $M(s)\geq w_{\mathrm{succ}}$. The rewrite right-hand side $\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base}))$ contributes $w_{\mathrm{wrap}}+M(s)$ beyond the shared recursive subterm, while the left-hand side $\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base}))$ contributes only $w_{\mathrm{succ}}$. Since $w_{\mathrm{wrap}}\geq 1$ and $M(s)\geq w_{\mathrm{succ}}$, the right-hand side is at least as large, contradicting strict orientation. For Tier 2, set $b=s=n=\mathrm{base}$. By transparency, $c_{\mathrm{succ}}(c_{\mathrm{base}})=c_{\mathrm{base}}$, so $M(\mathrm{recur}(\mathrm{base},\mathrm{base},\mathrm{succ}(\mathrm{base})))=c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}})$. The right-hand side $c_{\mathrm{wrap}}(c_{\mathrm{base}},c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}}))$ exceeds the left-hand side by the wrapper-subterm axiom $c_{\mathrm{wrap}}(x,y)>y$. $\square$
4.2 Affine / linear constructor-local measures
Theorem 4.2 (Affine / linear barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy
$M(\mathrm{succ}(t))=\alpha_{s}+\beta_{s}\,M(t),$
$M(\mathrm{wrap}(x,y))=\alpha_{w}+\beta_{w}\,M(x)+\gamma_{w}\,M(y),$
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n),$
with $\beta_{w},\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, i.e.
$\forall k\in\mathbb{N}\;\exists t\in T.\;M(t)\geq k,$
then $M$ does not strictly orient the duplicating step uniformly. In particular, the same conclusion follows if either
  1. 1.
    $\alpha_{s}\geq 1$ and $\beta_{s}\geq 1$ (positive successor pump), or
  2. 2.
    $\alpha_{w}+\gamma_{w}\,M(\mathrm{base})\geq 1$ (positive wrap/base pump).
Machine-checked in Meta/StepDuplicatingSchema.lean.
Proof sketch. Fix $b=n=\mathrm{base}$ and pump $s$. Write
$C_{\mathrm{src}}:=\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\delta_{r}\,M(\mathrm{succ}(\mathrm{base})),\qquad C_{\mathrm{tgt}}:=\alpha_{w}+\gamma_{w}\!\bigl(\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\delta_{r}\,M(\mathrm{base})\bigr),$
so the two sides of the duplicating rule become
$M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))=C_{\mathrm{src}}+\gamma_{r}\,M(s),$
$M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))=C_{\mathrm{tgt}}+(\beta_{w}+\gamma_{w}\gamma_{r})\,M(s).$
Hence the pumped-variable coefficient on the target exceeds the source coefficient by
$(\beta_{w}+\gamma_{w}\gamma_{r})-\gamma_{r}=\beta_{w}+(\gamma_{w}-1)\gamma_{r}\geq 1.$
Once $M(s)$ is larger than the fixed constant gap between $C_{\mathrm{src}}$ and $C_{\mathrm{tgt}}$, the target exceeds the source, contradicting strict orientation. The later scalar barrier theorems reuse this same coefficient comparison after freezing their additional nonlinear terms at the base point. $\square$
Sharpness note. The contradiction bound used in Theorem 4.2 is not merely a proof artifact. In Meta/AffineThresholdSharpness.lean we define a canonical affine family on $\mathbb{N}$ with
$M(\mathrm{base})=0,\qquad M(\mathrm{succ}(t))=1,\qquad M(\mathrm{wrap}(x,y))=x+y,\qquad M(\mathrm{recur}(b,s,n))=\delta_{r}\,n,$
for which the generic affine pump bound is exactly $\delta_{r}$, and the distinguished root-step instance satisfies
$M(\mathrm{wrap}(k,\mathrm{recur}(\mathrm{base},k,\mathrm{base})))<M(\mathrm{recur}(\mathrm{base},k,\mathrm{succ}(\mathrm{base})))\quad\Longleftrightarrow\quad k<\delta_{r}.$
Thus the generic affine contradiction bound is exact for this family. Machine-checked as affineThresholdMeasure_exact_cutoff.
Scope note. Theorems 4.1 and 4.2 cover the four-role schema and three axiomatized measure classes. They do not cover nonlinear polynomial interpretations with unrestricted cross-term coupling, generic DP frameworks, or duplicating rule shapes not reduced to a single wrapper. The following extensions address restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, max-plus, and tracked or projected vector / pair classes within the same schema; the full scope discussion is in §4.8.
Theorem 4.3 (Restricted quadratic counter barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy the same affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$ as in Theorem 4.2, and let the recursor take the restricted degree-$2$ form
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n)+\varepsilon_{r}\,M(n)^{2},$
with $\beta_{w},\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump, exactly as in Theorem 4.2. Machine-checked in Meta/QuadraticBarrier.lean as no_quadratic_counter_orients_dup_step_of_unbounded together with the successor- and wrap-pump corollaries.
Proof sketch. Fix $b=n=\mathrm{base}$ and pump the step argument $s$. The rewrite source is
$M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))={}$ $\underbrace{\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\delta_{r}\,M(\mathrm{succ}(\mathrm{base}))}_{\text{constant in }M(s)}$
$+\underbrace{\varepsilon_{r}\,M(\mathrm{succ}(\mathrm{base}))^{2}}_{\text{constant in }M(s)}+\gamma_{r}\,M(s).$
The rewrite target is
$M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))={}$ $\alpha_{w}+\beta_{w}\,M(s)$
$+\gamma_{w}(\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\gamma_{r}\,M(s)$
$\qquad+\delta_{r}\,M(\mathrm{base})+\varepsilon_{r}\,M(\mathrm{base})^{2}).$
[…]
Theorem 4.4 (Bounded cross-term quadratic barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy the same affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$ as in Theorem 4.2, and let the recursor take the degree-$2$ form
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n)+\varepsilon_{r}\,M(n)^{2}+\zeta_{r}\,M(s)M(n).$
Assume $M$ has unbounded attainable values, and that the cross-term coupling satisfies the bounded-regime condition
$\gamma_{r}+\zeta_{r}\,M(\mathrm{succ}(\mathrm{base}))+1\;\leq\;\beta_{w}+\gamma_{w}\bigl(\gamma_{r}+\zeta_{r}\,M(\mathrm{base})\bigr),$
formalized in Meta/QuadraticCrossTermBarrier.lean as CrossTermBoundedAtBase. Then $M$ does not strictly orient the duplicating step uniformly. In particular, the same conclusion follows from either a positive successor pump or a positive wrap/base pump together with this bounded-coupling condition. Machine-checked in Meta/QuadraticCrossTermBarrier.lean as no_cross_quadratic_orients_dup_step_of_unbounded and the corresponding KO7 specializations.
Non-vacuous example. The bounded-coupling condition is not an ad hoc exclusion of the known nonlinear witness. For example, if $M(\mathrm{base})=0$, $M(\mathrm{succ}(t))=1+M(t)$, $M(\mathrm{wrap}(x,y))=1+2M(x)+M(y)$, and
$M(\mathrm{recur}(b,s,n))=M(b)+M(n)+M(n)^{2}+M(s)M(n),$
then $\zeta_{r}=1$ gives genuine step-counter coupling, yet CrossTermBoundedAtBase still holds because the wrapper gain on the duplicated step argument dominates the one-step cross-term increment at the base point.
Proof sketch. Fix $b=n=\mathrm{base}$ and pump the step argument $s$. The source coefficient in the pump variable becomes $\gamma_{r}+\zeta_{r}\,M(\mathrm{succ}(\mathrm{base}))$, while the target coefficient becomes $\beta_{w}+\gamma_{w}(\gamma_{r}+\zeta_{r}\,M(\mathrm{base}))$. The bounded-coupling hypothesis says the target coefficient still dominates the source coefficient by at least one. Once the pumped step argument exceeds the frozen constant part, the target overtakes the source and strict decrease is impossible. This pushes the barrier one step beyond the pure counter-square class: some step-counter coupling is still excluded, provided the wrapper gain remains dominant at the base point. $\square$
Theorem 4.5 (Bounded multilinear barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy the same affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$ as in Theorem 4.2, and let the recursor take the form
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}M(b)+\gamma_{r}M(s)+\delta_{r}M(n)+\sum_{m\in\mathcal{T}}m(M(b),M(s),M(n)),$
where $\mathcal{T}$ is a finite table of multilinear monomials, each variable appearing at most once per monomial. Assume $M$ has unbounded attainable values and that, after freezing $b=\mathrm{base}$ and comparing the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$, the total coefficient of the pumped step variable on the source side is still dominated by the wrapper-side coefficient. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump under the same dominance condition. Machine-checked in Meta/MultilinearBarrier.lean as no_multilinear_orients_dup_step_of_unbounded together with the successor-pump, wrap-pump, and KO7 corollaries.
Proof sketch. Freeze $b=\mathrm{base}$ and pump the step argument $s$. Each multilinear monomial becomes affine in $M(s)$ once the other arguments are frozen, so the entire finite monomial table collapses to a constant part plus a frozen coefficient times $M(s)$ on both the source and target sides. The explicit dominance condition says the wrapper-side coefficient still exceeds the frozen source coefficient by at least one. Once $M(s)$ exceeds the frozen constant part, the target overtakes the source and strict decrease becomes impossible. This sharpens the nonlinear boundary: a large finite multilinear family still fails as long as its pumped-variable contribution remains frozen and dominated in this sense. $\square$
Theorem 4.6 (Generalized degree-bounded polynomial barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy constructor-local affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$, and let the recursor be given by a finite polynomial table in the three tracked scalar arguments $M(b)$, $M(s)$, and $M(n)$, with arbitrary natural exponents allowed in each monomial. Assume $M$ has unbounded attainable values and that, after freezing $b=\mathrm{base}$ and comparing the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$, the frozen target-side polynomial eventually dominates the frozen source-side polynomial as the pumped step value grows. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a certified successor pump or a certified wrap pump under the same frozen dominance condition. Equivalently, any successful orienter inside this bounded direct-polynomial family must violate the frozen base-dominance condition. Machine-checked in Meta/PolynomialBarrierGeneral.lean as no_polynomial_orients_dup_step_of_unbounded, no_global_step_orientation_polynomial_of_succ_pump, and polynomial_escape_requires_failure_of_base_dominance.
Proof sketch. Freeze the non-pumped arguments at the base point and compare the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$. Each monomial then becomes a one-variable polynomial in the pumped step value, possibly with repeated occurrences of that variable. The explicit dominance hypothesis says that beyond some threshold the frozen target-side polynomial overtakes the frozen source-side polynomial. Pump the step value beyond that threshold. The target is then at least the source, so strict decrease fails. This extends the freeze-and-pump argument beyond the multilinear regime: repeated variables are allowed, but only if the resulting one-variable growth remains dominated at the frozen base point. $\square$
Theorem 4.7 (Schema-level max-plus barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy constructor-local max-plus equations of the form
$M(\mathrm{succ}(t))=\alpha_{s}+M(t),$
$M(\mathrm{wrap}(x,y))=\alpha_{w}+\max(\beta_{w}+M(x),\gamma_{w}+M(y)),$
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\max(\beta_{r}+M(b),\max(\gamma_{r}+M(s),\delta_{r}+M(n))).$
If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor drift or a positive left-wrapper drift. Machine-checked in Meta/MaxBarrier.lean as no_max_orients_dup_step_of_unbounded together with the successor-pump, wrap-pump, and KO7 corollaries.
Proof sketch. Pump the step argument until the step branch $\gamma_{r}+M(s)$ dominates both the base branch and the counter branch inside the recursor. At that point the source becomes $\alpha_{r}+\gamma_{r}+M(s)$, while the target wraps the same visible branch again through a max-plus wrapper. Because the wrapper keeps a visible copy of the pumped branch, the target is at least the source plus one of the wrapper contributions, so strict decrease fails. This adds a different aggregation regime to the barrier stack: the obstruction persists even when aggregation is by $\max$ rather than by addition. $\square$
Corollary 4.8 (Arctic primary-projection barrier).
Any direct arctic-style interpretation with a distinguished finite primary projection satisfying the max-plus interface of Theorem 4.7 is blocked by the same argument. In particular, the KO7 root relation admits no such arctic primary-projection orienter under the corresponding unbounded or certified pump hypotheses. Machine-checked in Meta/ArcticBarrier.lean as no_arctic_primary_orients_dup_step_of_unbounded and the KO7 corollary no_global_step_orientation_arctic_primary_of_unbounded.
Proof sketch. Forget the arctic carrier and read only the distinguished finite primary coordinate. By hypothesis that projection already satisfies the max-plus interface, so Theorem 4.7 applies to it directly. The arctic corollary is therefore a tool-facing instantiation of the scalar max barrier, not a separate max-algebra proof. $\square$
The same projection pattern extends further to a broader tropical primary-projection corollary, formalized in Meta/TropicalBarrier.lean. We treat that module as an artifact-level continuation of the max-plus / arctic line rather than as a thirteenth direct family in the main theorem count.
Theorem 4.9 (Fixed-dimension componentwise affine barrier).
Fix any finite dimension $d\geq 1$. Let $M:T\to\mathbb{N}^{d}$ be a vector-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume one tracked coordinate of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that tracked coordinate. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, the same conclusion follows if the tracked coordinate admits either a positive successor pump or a positive wrap/base pump. Machine-checked in Meta/MatrixBarrierD.lean as no_matrixD_orients_dup_step_of_componentwise_pump and the corresponding KO7 specialization. The earlier dimension-$2$ development is retained in Meta/MatrixBarrier2.lean as the concrete low-dimensional worked instance.
Proof sketch. Strict componentwise orientation implies strict decrease in every coordinate, hence in the tracked one. Project the vector-valued measure to that tracked coordinate and apply Theorem 4.2. This yields a contradiction immediately. The result remains intentionally narrow: it covers arbitrary fixed dimension under strict componentwise order, but still only through one tracked affine coordinate, not through arbitrary mixed matrix interactions. $\square$
Theorem 4.10 (Dimension-$2$ lexicographic affine barrier).
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered lexicographically by the first component. Assume the first component of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that first component. Then the duplicating step cannot be strictly oriented lexicographically by $M$. In particular, the same conclusion follows if the first component admits either a positive successor pump or a positive wrap/base pump. Machine-checked in Meta/MatrixBarrierLex.lean as no_matrix2_lex_orients_dup_step_of_unbounded_primary and the corresponding KO7 specializations.
Proof sketch. Lexicographic decrease implies that the first component of the target is at most the first component of the source. Pump the tracked first component one unit beyond the scalar affine bound. The first component of the duplicating-step target then exceeds the first component of the source, so lexicographic decrease is impossible regardless of what the second component does. This remains a tracked-primary result, not a theorem about arbitrary mixed matrix interpretations. $\square$
Theorem 4.11 (Balanced mixed-coordinate dimension-$2$ barrier).
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume each constructor acts by a full mixed $2\times 2$ linear map on the two coordinates, and that the two column sums of each such map agree. Then the aggregate coordinate sum $M_{1}+M_{2}$ is an affine scalar measure. If that aggregate sum has an unbounded pump and the aggregate wrapper coefficients are positive, then the duplicating step cannot be strictly oriented componentwise by $M$. Machine-checked in Meta/MatrixBarrierMix2.lean as no_matrixMix2_orients_dup_step_of_sum_pump and the corresponding KO7 specialization.
Proof sketch. Balanced column sums make the aggregate quantity $M_{1}(t)+M_{2}(t)$ behave like a scalar affine interpretation, even though each coordinate may mix both inputs through off-diagonal coefficients. Strict componentwise decrease forces strict decrease of this aggregate sum, so Theorem 4.2 applies to the aggregate projection. This is the first matrix-style barrier in the artifact whose proof is not a single-coordinate projection theorem, but it still stops short of arbitrary mixed matrix interpretations. $\square$
Theorem 4.12 (Weighted functional componentwise affine barrier).
Fix any finite dimension $d\geq 1$. Let $M:T\to\mathbb{N}^{d}$ be a vector-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume there is a fixed nonzero natural weight vector $\lambda\in\mathbb{N}^{d}$ such that the scalar projection
$\lambda\cdot M(t):=\sum_{i}\lambda_{i}M_{i}(t)$
satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that weighted projection. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, singleton weights recover the tracked-coordinate argument, while the all-ones vector recovers aggregate-sum arguments. Machine-checked in Meta/MatrixBarrierFunctional.lean as no_matrixFunctional_orients_dup_step_of_componentwise_pump together with the successor-pump and wrap-pump corollaries and the KO7 specialization.
Proof sketch. Strict componentwise decrease forces strict decrease of every nonzero natural-weighted scalar projection. Apply this to $\lambda\cdot M$. By hypothesis, that projection already satisfies the scalar affine interface of Theorem 4.2. The contradiction therefore reduces to the scalar affine barrier after one projection step. This unifies the tracked-coordinate proof shape and the balanced aggregate-sum proof shape under one matrix-side principle: once a fixed scalar projection carries the affine interface, strict componentwise orientation is impossible. $\square$
Theorem 4.13 (Scalar-projection meta-barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $\mu:T\to A$ be any direct orienter into a codomain $A$, let $R$ be the chosen strict order on $A$, and let $\pi:A\to\mathbb{N}$ be a tracked scalar projection. If every strict $R$-decrease forces strict decrease of $\pi$, and if strict decrease of $\pi\circ\mu$ on the duplicating step is already ruled out by a scalar barrier theorem, then $R$ itself cannot orient the duplicating step uniformly. Machine-checked in Meta/ScalarProjectionBarrier.lean as no_orients_dup_step_of_scalar_projection, with concrete rederivations for tracked componentwise pairs, weighted functional matrices, and balanced mixed-coordinate pairs.
Proof sketch. Suppose $R$ oriented the duplicating step uniformly. By hypothesis, every such $R$-decrease forces strict decrease of the projected scalar $\pi$. Then the projected scalar family $\pi\circ\mu$ would also orient the duplicating step uniformly, contradicting the already available scalar barrier theorem. The value of the theorem is conceptual: it states once the common proof shape shared by the tracked-component, weighted-functional, and balanced mixed-coordinate matrix barriers. $\square$
Theorem 4.14 (Symbolic variable-condition barrier).
Let $\succ$ be any strict symbolic comparator on the abstract rule schema
$\mathrm{recur}(b,s,\mathrm{succ}(n))\to\mathrm{wrap}(s,\mathrm{recur}(b,s,n))$
such that $\succ$ satisfies the standard variable condition: if $x\succ y$, then every variable occurs in $y$ at most as often as in $x$. Then $\succ$ cannot orient the duplicating step from left to right. Machine-checked in Meta/SymbolicComparatorBarrier.lean as not_orients_dup_rule.
Proof sketch. The variable $s$ occurs once on the source side and twice on the target side. Any order satisfying the variable condition must therefore reject this orientation immediately. This widens the negative story beyond numeric measures: direct KBO-style symbolic comparators already fail at the rule schema before any recursive path descent or modular transformation is introduced. $\square$
Corollary 4.15 (Explicit KO7 KBO impossibility).
No KBO-style order satisfying the standard variable condition can orient the KO7 rec_succ rule, and hence no such KBO-style order can orient the KO7 root relation by decreasing every root step. Machine-checked in Meta/KBO_Impossible.lean as no_kbo_orients_ko7_rec_succ.
Proof sketch. Instantiate Theorem 4.14 with the KO7 duplicating rule schema. The corollary adds no new proof mechanism; it simply records the KBO-facing consequence under a theorem name reviewers can cite directly. $\square$
This is a KO7-facing corollary extracted from the symbolic variable-condition barrier, not a formalization of generic Knuth–Bendix order metatheory over arbitrary signatures.
Note on the vector / pair extensions. The fixed-dimension componentwise barrier and the lexicographic dimension-$2$ barrier (Theorems 4.9 and 4.10) extend the scope of applicability of the barrier beyond scalar measures, but the core obstruction mechanism remains the affine tracked-component argument of Theorem 4.2. The balanced mixed-coordinate theorem (Theorem 4.11) goes one step further: it allows off-diagonal coordinate mixing, but only inside a balanced regime where the aggregate sum of the two coordinates still yields a scalar affine projection. The weighted functional theorem (Theorem 4.12) abstracts both moves: any nonzero fixed scalar projection with an affine interface inherits the same barrier. The scalar-projection meta-theorem (Theorem 4.13) then states that proof shape once and for all. So the matrix story now has four layers, tracked-coordinate, balanced mixed-coordinate, weighted scalar projection, and projection-level unification, while arbitrary mixed matrix interpretations remain outside the present barrier stack.
Corollary 4.16 (Unconditional pumped-subclass forms).
The growth-side hypotheses in the affine, restricted-quadratic, bounded cross-term, generalized bounded-polynomial, bounded multilinear, max-plus, and current projection-based matrix barriers can be internalized into strengthened subclasses. The resulting affine, restricted-quadratic, bounded cross-term, generalized bounded-polynomial, bounded multilinear, max-plus, tracked-primary pair, weighted functional, and balanced mixed-coordinate families therefore admit unconditional exclusions inside their subclass definitions. Machine-checked in Meta/PumpedBarrierClasses.lean; the exact identifiers are listed in Table LABEL:tab:claim-code-barrier.
The bridge from the conditional barrier theorems to these unconditional pumped-subclass corollaries is factored out in Meta/StandardPumpLemmas.lean. That module proves reusable successor-pump and wrap-pump lemmas for the affine, restricted-quadratic, and current tracked-primary dimension-$2$ pair families, and packages the corresponding hypotheses into the strengthened subclasses used by Corollary 4.16. This keeps the theorem statements honest while making the standard growth side conditions available as reusable certified infrastructure.
Corollary 4.17 (Context-closed barrier survival).
Every direct class blocked at the duplicating root step is also blocked for the full context-closed relation StepCtxFull, because every root contraction is a contextual contraction under the identity context. In particular, the additive, transparent-compositional, affine-with-pump, restricted-quadratic, bounded cross-term, bounded multilinear, max-plus, tracked-primary pair, weighted scalar-projection, and balanced mixed-coordinate families all admit explicit context-closed corollaries. Machine-checked in Meta/ContextClosedBarrier.lean as stepCtxFull_orientation_implies_root together with the corresponding class-specific KO7 corollaries.
This corollary does not add new obstruction mechanisms. Its role is semantic: it makes the root-step barrier stack available directly at the same context-closed level used by standard TRS tools and by the external TPDB / CeTA semantics in Section 6.
4.3 KO7 instantiation
Theorem 4.18 (Compositional impossibility for rec_succ).
KO7 instantiates the schema with $\mathrm{base}=\mathrm{void}$, $\mathrm{succ}=\delta$, $\mathrm{wrap}=\mathrm{app}$, $\mathrm{recur}=\mathrm{rec}\Delta$. Both impossibility results hold:
  1. 1.
    Tier 1. For every additive compositional measure $M$ with app-weight $\geq 1$:
    $\neg\bigl(\forall b\,s\,n.\;M(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))<M(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))\bigr).$
  2. 2.
    Tier 2. For every abstract compositional measure $CM$ with app-subterm axioms and $\delta$-transparency at void:
    $\neg\bigl(\forall b\,s\,n.\;CM(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))<CM(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))\bigr).$
Proof sketch. Direct instantiation of Theorem 4.1 with $\mathrm{base}=\mathrm{void}$, $\mathrm{succ}=\delta$, $\mathrm{wrap}=\mathrm{app}$, $\mathrm{recur}=\mathrm{rec}\Delta$. Machine-checked in Meta/CompositionalMeasure_Impossibility.lean. $\square$
The same module also gives the KO7 affine instantiation. It packages a 7-constructor affine evaluation as AffineCompositionalMeasure, lifts it to the schema via toSchemaMeasure, and yields the corresponding KO7 affine barrier with pump corollaries. The KO7 restricted-quadratic and bounded cross-term extensions then appear in Meta/QuadraticBarrier.lean and Meta/QuadraticCrossTermBarrier.lean, each with unbounded, successor-pump, and wrap-pump forms. The multilinear, generalized bounded-polynomial, max-plus, and arctic primary-projection continuations then live in Meta/MultilinearBarrier.lean, Meta/PolynomialBarrierGeneral.lean, Meta/MaxBarrier.lean, and Meta/ArcticBarrier.lean. Meta/PumpedBarrierClasses.lean repackages the affine, restricted-quadratic, bounded cross-term, generalized bounded-polynomial, tracked-primary pair, bounded multilinear, max-plus, weighted functional, and balanced mixed-coordinate KO7 exclusions as unconditional statements for the corresponding strengthened subclasses.
4.4 Essentiality of the transparency hypothesis
Theorem 4.19 (Transparency is essential for Tier 2).
[…]
$\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n)\;<\;\mathrm{rec}\Delta\;b\;s\;(\delta\;n)$
for all $b,s,n$. One concrete witness is
$c_{\mathrm{void}}=1,\qquad c_{\delta}(x)=x+1,\qquad c_{\mathrm{app}}(x,y)=x+y+1,\qquad c_{\mathrm{rec}\Delta}(a,b,c)=a+b(c+1)^{2}+c.$
Once transparency is dropped, the Tier 2 class contains measures that orient the step. Machine-checked in Meta/RecCore.lean.
[…]
Remark 4.20 (Every successful method imports assumptions outside the barrier classes).
The essentiality theorem and the DP escape (Theorem 4.21) instantiate a general pattern: every method that orients the duplicating step does so by violating at least one axiom of the formalized compositional classes. DP projection violates the wrapper-subterm axioms (setting $\pi(\mathrm{app}\;x\;y)=0$). LPO imports the Subterm Property, a structural axiom not derivable from the rewrite rules. The nonlinear witness above violates successor transparency ($c_{\delta}(1)=2\neq 1$) and uses a nonlinear recursor combiner not representable in Tier 1 or the affine class of Theorem 4.2. An arbitrary nonlinear polynomial interpretation over $\mathbb{Z}^{+}$ can likewise orient the full system. For example, assigning $W(\mathrm{rec}\Delta(b,s,n))=(W(n)+1)\cdot(W(s)+W(b)+1)$ with $W(\delta\,t)=W(t)+1$ yields $W(\text{LHS})-W(\text{RHS})=W(b)\geq 1$ on rec_succ. This requires both nonlinearity (cross-term coupling between counter and payload) and non-transparency ($W(\delta\,t)=W(t)+1\neq W(t)$), which place it outside every formalized barrier class. This full-system nonlinear interpretation is machine-checked: W_orients_step in Meta/PolyInterpretation_FullStep.lean proves $W(b)<W(a)$ for every root step $a\to b$, with axiom-violation witnesses W_violates_transparency, W_not_additive, and W_not_affine.
The barrier does not assert that no termination proof exists. It characterizes precisely which structural assumptions any successful proof must import. A method that works on the duplicating step confirms the barrier’s precision by demonstrating which axioms it violates.
Method Additive Transp.-comp. Affine Escape mechanism
$W$ (nonlinear poly.) nonlinear cross-variable coupling
KO7-specialized MPO recursive subterm comparison
DP projection modular SCC decomposition
Mechanized violation witnesses: W_violates_transparency, W_not_additive, W_not_affine in Meta/PolyInterpretation_FullStep.lean.
4.5 The DP escape
Theorem 4.21 (DP projection escapes the impossibility).
Define $\pi:\mathrm{Trace}\to\mathbb{N}$ by:
$\pi(\mathrm{void})=0,\quad\pi(\delta\,t)=\pi(t)+1,\quad\pi(\mathrm{rec}\Delta\,b\,s\,n)=\pi(n),\quad\pi(t)=0\text{ otherwise.}$
Then:
  1. 1.
    $\pi$ orients rec_succ: $\pi(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))<\pi(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))$ for all $b,s,n$.
  2. 2.
    $\pi$ violates the app-subterm axioms: $\pi(\mathrm{app}\;(\delta\,\mathrm{void})\;\mathrm{void})=0\not>1=\pi(\delta\,\mathrm{void})$.
[…]
Proposition 4.22 (The extracted DP pair problem already has a simple linear base order).
There exists a linear polynomial-style base order $\mu:T\to\mathbb{N}$ such that every extracted KO7 dependency pair decreases strictly under $\mu$. Consequently, there is no blanket theorem saying that dependency pairs still fail whenever the chosen base order remains affine or polynomial.
Machine-checked in Meta/DP_BaseOrder_Boundary.lean as counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order, and no_blanket_dp_poly_base_barrier. The point is boundary-setting rather than escape construction: modularity alone is not enough to force failure, but neither does it trivialize the direct barrier story. What matters is the transformed problem and the structure imported by the chosen base order.
4.6 Ablation: removing duplication dissolves the barrier
[…]
$\mathrm{recL}\,b\,s\,(\delta\,n)\;\to\;\mathrm{recL}\,b\,s\,n$
where $s$ is not duplicated. The simpleSize measure (Tier 1, all constructor weights equal to 1) strictly orients both rules:
$\mathrm{simpleSize}(\mathrm{recL}\,b\,s\,(\delta\,n))$ $=1+\mathrm{simpleSize}(b)+\mathrm{simpleSize}(s)$
$\quad+1+\mathrm{simpleSize}(n)$
$>1+\mathrm{simpleSize}(b)+\mathrm{simpleSize}(s)$
$\quad+\mathrm{simpleSize}(n)$
$=\mathrm{simpleSize}(\mathrm{recL}\,b\,s\,n).$
Machine-checked in Meta/LinearRec_Ablation.lean. Removing duplication reduces the problem to a simple additive size measure: the $\delta$-wrapper weight alone drives the decrease, with no need for the multiset or phase-bit components of $\mu^{3}_{c}$. This confirms that step-duplication is the operative source of every barrier theorem in this paper. Symmetric removal theorems for conditions (1) through (3) of §4 are not provided.
Directed wrapper-collapse surrogate.
We also tested a bounded first-order surrogate for the “no axioms inside the object language” condition by extending KO7 with the single directed rule
$\mathrm{app}(t,t)\to t.$
This relaxation does not dissolve the direct barrier. The added collapse rule is harmless in isolation. It is strictly oriented by simpleSize, and the nonlinear full-step witness $W$ still proves root-step termination of the extended relation. However, the additive, transparent-compositional, and affine barrier theorems continue to apply because the duplicating rule rec_succ is unchanged. Machine-checked in Meta/ObjectAxiom_Ablation.lean as simpleSize_orients_app_collapse, wf_StepCollapseRev, no_global_stepCollapse_orientation_additive, no_global_stepCollapse_orientation_compositional_transparent, and no_global_stepCollapse_orientation_affine_of_unbounded. So this particular directed collapse surrogate is not the missing escape mechanism: it changes the system, but it does not neutralize the duplication obstruction isolated above. It is only a first-order syntactic proxy for collapse inside the rewrite language, not a replacement for full rewriting modulo an equational theory.
A separate artifact-only surrogate in Meta/SharingBarrierLift.lean makes the complementary point for condition (2) of §4: once the duplicated payload is represented through explicit sharing rather than copied in the tree, the tree-style direct barrier no longer applies verbatim. The module packages this as a positive counterexample (sharedCounter_orients_step) and a boundary theorem (sharing_breaks_tree_barrier), so the no-sharing hypothesis is not merely presentational.
Bounded SCC composite duplication.
Theorem 4.23 (Bounded delayed-duplication barrier).
Consider a two-node mutually recursive SCC with rules
$\mathrm{recurA}(b,s,\mathrm{succ}(n))$ $\to\mathrm{wrap}(s,\mathrm{recurB}(b,s,n)),$
$\mathrm{recurB}(b,s,\mathrm{succ}(n))$ $\to\mathrm{wrap}(s,\mathrm{recurA}(b,s,n)).$
Each root rule already duplicates the payload once: the same term $s$ appears both as the left wrapper argument and again as the recursive step argument on the right-hand side. The point of the theorem is different: the two-node SCC yields a stable two-step composite profile and an induced minimal context relation on which the additive and affine barrier arguments still go through. Concretely, one root step followed by one step under the right argument of wrap realizes the effective pattern
$\mathrm{recurA}(b,s,\mathrm{succ}(\mathrm{succ}(n)))\;\leadsto\;\mathrm{wrap}(s,\mathrm{wrap}(s,\mathrm{recurA}(b,s,n))).$
Then no additive direct measure on the shared constructors can orient the composite pattern uniformly. Equivalently, no such additive measure globally orients the induced minimal context relation generated by root steps and right-wrapper descent. Under the same kind of unbounded-pump hypothesis used in Theorem 4.2, no affine constructor-local direct measure can orient the same composite pattern uniformly either. The corresponding global affine failure for the induced minimal context relation is mechanized in Meta/MutualDuplication_General.lean; the exact identifiers appear in Table LABEL:tab:claim-code-barrier.
The concrete AltTerm development in Meta/MutualDuplication_Case.lean remains as the worked instance of this bounded SCC theorem. In that instance, the effective target carries three copies of the step payload measure against one copy on the source side, so the additive contradiction closes by the same coefficient comparison as in the schema barrier. The affine companion theorem uses the same two-step composite profile but packages it as a derived duplication schema with doubled successor and nested wrapper, so the existing affine barrier applies under the corresponding unbounded-pump hypothesis. The SCC layer now extends one step further: Meta/MutualDuplication_Transparent.lean proves that the same alternating composite is still blocked for transparent-compositional Tier 2 measures (no_global_orients_ctx_compositional_transparent) and for weighted functional matrix-style families whose projected scalar interface is already blocked (no_global_orients_ctx_matrixFunctional_of_projected_unbounded). A multiplicity-preserving SCC theorem still requires a different mechanism: in first-order tree rewriting, once every rule preserves the relevant payload multiplicities, the current coefficient-growth argument no longer has a duplicated payload to pump against.
Theorem 4.24 (Multiplicity-preserving synchronized SCC barrier).
There exists a two-node first-order SCC together with an explicit payload-counting function such that each individual root rule preserves payload multiplicity exactly. On synchronized inputs carrying the same latent payload on two tagged channels, one full SCC cycle
$\mathrm{recurA}(c,\mathrm{left}(s),\mathrm{right}(s))\;\leadsto\;\mathrm{wrap}(s,\mathrm{wrap}(s,\mathrm{recurA}(c,\mathrm{base},\mathrm{base}))).$
The source has no visible wrapper-carried payload, while the target exposes two visible wrapper copies of the synchronized payload and still preserves the counted multiplicity. Consequently no additive direct measure on this preserving SCC syntax can orient the synchronized composite uniformly. Equivalently, no such additive measure globally orients the induced minimal context relation generated by root steps and right-wrapper descent. Machine-checked in Meta/MutualDuplication_Preserving.lean as step_preserves_payloadCount, synchronized_cycle_realized, synchronized_cycle_exposes_payload, no_additive_orients_synchronized_cycle, and no_global_orients_ctx_additive.
This theorem is stronger than Theorem 4.23 in one precise sense: the local SCC rules no longer increase the tracked payload multiplicity at all. The obstruction comes from synchronized exposure, not from the immediate coefficient-growth profile used in the earlier alternating theorem.
The mechanization also proves a conditional affine extension: under a wrapper-dominance condition $r_{p}+r_{q}<w_{l}+w_{r}\cdot w_{l}$ on the affine coefficients (where $r_{p},r_{q}$ are the recursor’s payload scaling coefficients and $w_{l},w_{r}$ are the wrapper’s left and right scaling coefficients), no affine measure with an unbounded pump can orient the synchronized cycle. The proof pumps the payload evaluation past the source constant, then exploits the coefficient gap to force the target above the source. Machine-checked as no_affine_orients_synchronized_cycle_of_wrapper_dominance in Meta/MutualDuplication_Preserving.lean.
4.7 Proof-theoretic context
Table 3: Five method classes applied to KO7. Every class that succeeds imports structural assumptions outside the formalized barrier (Remark 4.20). The proof-strength column refers to the general metatheory where marked.
Method class Orients KO7? Proof strength Mechanism / imported assumption
Direct/global aggregation (Def. 3.2) No typically arithmetical Aggregates subterms; duplication defeats orientation
Affine/linear constructor-local (Thm. 4.2) No typically arithmetical Coincides with Tier 1 when all scales are 1; independent under unbounded pump
Nonlinear polynomial interpretation Yes (L) PRA Cross-term coupling + non-transparent successor; outside Tier 1, Tier 2, and affine classes (W_orients_step)
DP + subterm criterion Yes elementary ($\ll\varepsilon_{0}$) Projects away duplication; violates wrapper-subterm axioms
[…]
KO7-specialized MPO Yes (L) direct ordinal ranking below $\varphi_{7}(0)$ Fixed 7-constructor signature, fixed precedence; explicit upper bound in Meta/MPO_ProofTheoreticBound.lean (wf_StepRev_mpo_below_veblen7)
[…]
Corollary 4.25 (Non-representability of successful Nat-valued orienters).
Let $\mu:T\to\mathbb{N}$ globally orient a step relation containing the duplicating rule. Then $\mu$ is not representable as
  1. 1.
    an additive schema measure,
  2. 2.
    a Tier 2 compositional measure with successor transparency at $\mathrm{base}$, or
  3. 3.
    an affine / linear constructor-local measure satisfying the unbounded-pump hypothesis of Theorem 4.2.
Machine-checked in Meta/StepDuplicatingSchema.lean.
This is the precise sense in which path-order success lies outside the present Nat-valued measure framework: not because path orders are unsound, but because successful orienters cannot be represented by the formal classes isolated above.
Proposition 4.26 (Nonlinear polynomial full-step escape).
There exists a $\mathbb{N}$-valued interpretation $W:\mathrm{Trace}\to\mathbb{N}$ such that:
  1. 1.
    for every root step $a\to b$ in Step, one has $W(b)<W(a)$, hence the reverse root relation is well-founded;
  2. 2.
    $W$ violates $\delta$-transparency and is neither additive nor affine in the sense of the formalized barrier classes.
Machine-checked in Meta/PolyInterpretation_FullStep.lean as W_orients_step, wf_StepRev_poly, W_violates_transparency, W_not_additive, and W_not_affine.
Proposition 4.27 (KO7-specialized MPO full-step termination).
For the fixed KO7 signature, precedence, and same-head clause formalized in Meta/MPO_FullStep.lean, the reverse root relation induced by Step is well-founded. Machine-checked as wf_MPORev together with the induced reverse-Step theorem wf_StepRev_mpo. This is a fixed-signature KO7 result, not a generic MPO well-foundedness theorem over arbitrary signatures.
Proposition 4.28 (Proof-theoretic upper bound for the KO7-specialized MPO proof).
The fixed-signature KO7 MPO proof is witnessed by an ordinal ranking into the initial segment below $\varphi_{7}(0)$, that is, below Ordinal.veblen70. This gives a concrete proof-theoretic upper bound for the specialized MPO argument without appealing to generic path-order metatheory over arbitrary signatures. Machine-checked in Meta/MPO_ProofTheoreticBound.lean as mpoOrd_lt_mpoVeblenBound and wf_StepRev_mpo_below_veblen7.
Proposition 4.29 (Bad precedence blocks the duplicating rule).
[…]
$\mathrm{rec}\Delta(\mathrm{void},\mathrm{void},\delta(\mathrm{void}))\to\mathrm{app}(\mathrm{void},\mathrm{rec}\Delta(\mathrm{void},\mathrm{void},\mathrm{void})).$
Consequently that bad precedence cannot orient all root steps of Step. Machine-checked in Meta/MPO_Precedence_Barrier.lean as not_mpoBad_rec_succ_instance and no_global_step_orientation_mpo_bad_prec.
This supplies the negative companion to Proposition 4.27: KO7-specialized MPO success is genuine, but it is also precedence-sensitive.
Theorem universe. The following theorem is stated over an explicit direct universe consisting of twelve listed families: additive compositional, transparent-compositional, pumped affine, pumped restricted-quadratic, pumped bounded cross-term quadratic, pumped bounded multilinear, pumped generalized bounded-polynomial, pumped max-plus, tracked-primary componentwise pair, tracked-primary lexicographic pair, KO7-specific max-depth, and KO7-specific pure head-precedence. For the pair-valued families, the theorem reads their first component as the tracked primary scalar. The Lean development first proves the scalar-only preliminary ko7_nat_direct_escape_trichotomy; the theorem stated here is the KO7-specific extension ko7_direct_escape_trichotomy_extended, which adds the tracked-primary pair families. A second theorem, ko7_projection_escape_trichotomy, extends the same classification pattern to successful componentwise orienters that admit a fixed scalar projection whose strict decrease is forced by the ambient order and whose projected scalar lies in the listed Nat-valued universe. This is how the weighted scalar-projection and balanced mixed-coordinate matrix families are absorbed into the trichotomy layer. The generalized degree-bounded polynomial family of Theorem 4.6 is now internalized directly into the Nat-valued universe; projection-based matrix families still enter through the separate projection theorem. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside this theorem universe.
Theorem 4.30 (Escape trichotomy for the explicit direct universe).
Let $\mathcal{O}$ be a successful direct root-step orienter in the explicit theorem universe above, and let $\pi_{\mathcal{O}}:\mathrm{Trace}\to\mathbb{N}$ denote its tracked primary scalar, namely the orienter itself in the scalar cases and the first component in the tracked-pair cases. Then at least one of the following holds:
  1. 1.
    $\pi_{\mathcal{O}}$ fails wrapper-subterm sensitivity for app;
  2. 2.
    $\pi_{\mathcal{O}}$ fails successor transparency at void;
  3. 3.
    $\mathcal{O}$ is not representable by any of the direct families formalized in the artifact: additive compositional, transparent-compositional, affine-with-pump, restricted-quadratic-with-pump, bounded-cross-term-with-pump, bounded-multilinear-with-pump, generalized bounded-polynomial-with-pump, max-plus-with-pump, tracked-primary componentwise pair-with-pump, tracked-primary lexicographic pair-with-pump, max-depth, or pure head-precedence.
Equivalently, every successful direct root-step orienter in the theorem universe escapes by violating wrapper sensitivity on its tracked primary scalar, violating transparency there, or leaving the formalized direct families. Machine-checked in Meta/EscapeTrichotomy.lean as ko7_direct_escape_trichotomy_extended; the scalar-only preliminary theorem is ko7_nat_direct_escape_trichotomy.
Projection-based extension. The same escape pattern also covers the weighted scalar-projection componentwise family and the balanced mixed-coordinate componentwise family. The reason is structural rather than ad hoc: if the ambient componentwise order forces strict decrease of a fixed scalar projection, and that projected scalar already belongs to the explicit Nat-valued direct universe above, then the trichotomy applies to that scalar as well. Machine-checked in Meta/EscapeTrichotomy.lean as ko7_projection_escape_trichotomy, with the concrete representability lemmas matrixFunctional_projection_representable and matrixMix2_sum_projection_representable.
Scope note. The theorem is intentionally restricted to a finite, reviewable direct universe plus a separate projection-based extension principle. This is a deliberate design choice. A theorem over a listed class is verifiable and falsifiable, while an unbounded claim about all imaginable orienters would be either tautological, by defining “outside the formalized classes” to absorb every counterexample, or unprovable with current methods. The bounded cross-term, bounded multilinear, generalized bounded-polynomial, and max-plus scalar families are included directly; the weighted scalar-projection and balanced mixed-coordinate matrix families are included through the projection theorem; and the current matrix-side strengthening now also exposes fixed-row and row-sum projection corollaries in Meta/MatrixProjectionCoverage.lean. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside the present theorem universe, even though the paper separately proves negative or positive results for some of them.
4.8 Scope
Theorem 4.1 covers the four-role schema and the two measure classes of Definition 3.2. Theorem 4.2 adds constructor-local affine/linear measures under an explicit unbounded-growth hypothesis. The nonlinear scalar extensions cover restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus classes; the first two, the generalized bounded-polynomial family, and the multilinear / max-plus families also admit pumped strengthened subclasses in Meta/PumpedBarrierClasses.lean. The vector / pair layer covers fixed-dimension tracked componentwise orderings, dimension-$2$ tracked lexicographic orderings, balanced mixed-coordinate dimension-$2$ componentwise order, and weighted functional componentwise order. The escape trichotomy now classifies the scalar multilinear, generalized bounded-polynomial, and max-plus families directly and absorbs the weighted-functional and balanced mixed-coordinate families through the projection-based extension theorem. The matrix-side coverage is also sharper than before: Meta/MatrixProjectionCoverage.lean makes the fixed-row and row-sum corollaries explicit for the tracked-component and weighted-functional barrier layers. Out of scope are generic DP metatheory, arbitrary mixed matrix interpretations beyond these tracked, balanced, functional, fixed-row, and row-sum cases, arbitrary semantic methods, and duplicating-rule shapes not reduced to a single wrapper. In particular, nonlinear polynomial interpretations with cross-term coupling (such as the witness in Remark 4.20) remain outside the barrier, because they import nonlinearity and non-transparency not present in the formalized direct classes.
Status of full Step
Root-step Step is terminating: mechanized via the nonlinear polynomial interpretation (wf_StepRev_poly) and the KO7-specialized MPO (wf_StepRev_mpo). The full context closure is also strongly normalizing in Lean via wf_StepCtxFullRev_poly, and independently via TTT2/CeTA (§6). Full Step is not locally confluent.
Proposition 4.31 (The unguarded eqW overlap is never locally joinable).
For every trace $a$, the full kernel relation fails local join at $\mathrm{eqW}\;a\;a$: the two one-step reducts $\mathrm{void}$ and $\mathrm{integrate}(\mathrm{merge}\;a\;a)$ are distinct full-step normal forms.
Machine-checked in Meta/EqW_Guard_Barrier.lean as not_localJoinStep_eqW_refl. This is why the SafeStep eqW guards are genuine confluence guards rather than presentation artifacts.
Frontier summary.
The barrier section isolates a concrete frontier. Inside the formalized direct classes, the duplicating step is excluded by the schema, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, max-plus, tracked componentwise fixed-dimension vector, tracked lexicographic pair, balanced mixed-coordinate pair, and weighted scalar-projection componentwise barriers, together with the symbolic variable-condition / KBO corollary. The trichotomy theorem classifies the scalar side through max-plus and multilinear, and extends to the weighted-functional and balanced mixed-coordinate matrix families through projected representability, while the generalized polynomial theorem isolates the remaining bounded direct-polynomial escape condition. Crossing the frontier therefore requires one of three moves exhibited in the artifact: projection-style transformation (DP), direct nonlinear escape outside the formalized classes (the global witness $W$ and the KO7-specialized MPO), or ablation of the offending structure itself (removing duplication dissolves the barrier, while a directed wrapper-collapse surrogate does not). The SCC theorems show that this frontier is not confined to immediate single-rule duplication. Delayed duplication across one alternating cycle already recreates the additive and affine obstruction, and even a multiplicity-preserving SCC can force additive failure once a synchronized cycle exposes the latent payload under visible wrappers.
4.9 Orientation catalog
The theorems at the schema subsume most individual measure witnesses. Items such as additive counters, simple lex pairs, node counts, and linear weights are direct instances of Theorems 4.1 and 4.2. Tree depth and pure precedence are covered by dedicated KO7-specific families in Meta/DepthBarrier.lean and Meta/PrecedenceBarrier.lean. The following catalog entries record results not subsumed by those theorems or by those two KO7-specific families:
  1. 1.
    Polynomial tie (L/T): The multiplicative polynomial $M(\mathrm{rec}\Delta\,b\,s\,n)=M(b)+M(s)\cdot M(n)$ ties exactly on rec_succ (poly_mul_ties_rec_succ). TTT2 POLY returns MAYBE.
  2. 2.
    Multiset / Constellation (L): Single-pass multisets over nesting ranks fail (no_constellation_strict_drop_rec_succ).
  3. 3.
    Semantic Interpretation ($\dagger$): Model-theoretic methods are blocked by the duplicating constructor condition (§4).
  4. 4.
    Path Orders ($\dagger$/L): LPO orients the full TRS (CeTA-certified); this is a framework escape from the $\mathbb{N}$-valued measure classes (Cor. 4.25). Meta/MPO_FullStep.lean proves KO7-specialized MPO orientation (wf_MPORev, wf_StepRev_mpo); this is a fixed-signature comparator, not a generic MPO metatheorem.
The DP framework escapes by projecting to the recursion counter, as formalized in Theorem 4.21, and Proposition 4.22 shows that the extracted pair problem already admits a simple linear polynomial-style base order. Regarding polynomial interpretations more broadly: Theorem 4.2 covers the affine/linear fragment under positive wrapper coefficients and an unbounded pump, and Theorem 4.6 pushes that boundary to bounded degree-bounded polynomial tables with repeated variables under frozen base dominance, but neither theorem extends to arbitrary nonlinear polynomials. Theorem 4.19 gives a concrete nonlinear witness that overcomes duplication via cross-term coupling; any such polynomial necessarily imports structural assumptions outside the formalized barrier classes (Remark 4.20).
4.10 Gallery of explicitly excluded methods
Beyond the schema-level theorems and the catalog items above, Meta/Conjecture_Boundary.lean collects a broader mechanized no-go catalog for the full unrestricted root relation Step. Two stand out.
The Ladder Paradox (exact polynomial tie).
A polynomial measure with multiplicative recursor combiner $M(\mathrm{rec}\Delta(b,s,n))=M(b)+M(s)\cdot M(n)$ and additive app yields an exact equality on the duplicating step, regardless of base weight:
[…]
Machine-checked as poly_mul_ties_rec_succ. This is not merely non-strict failure but exact cancellation: the duplication of $s$ under app is algebraically absorbed by the multiplicative coupling. Only nonlinear cross-variable coupling (as in $W$) breaks this tie.
Dual barrier (conflicting structural demands).
The duplication barrier and the flag barrier target different rules and impose contradictory demands on any single $\mathbb{N}$-valued measure: rec_succ requires insensitivity to duplication of $s$ (additive size fails), while merge_void requires sensitivity to flag structure (the $\delta$-flag must strictly decrease). Machine-checked as dual_barrier_rec_succ_and_merge_void.
Further exclusions.
The remaining no-go results sharpen the negative catalog. Strictly monotone post-processing does not rescue an additive measure, and a single $\delta$-flag cannot orient Step. Nesting depth with merge-neutral aggregation ties on merge_cancel, while standard tree depth strictly increases on deep instances of rec_succ. Node count, pure head precedence, linear KBO weights, and constellation-size measures also fail on the unrestricted root relation. The exact theorem identifiers are listed in Meta/Conjecture_Boundary.lean and indexed in the appendix tables.
Decidability of reachability
In a strongly normalizing and confluent calculus with a computable normalizer and decidable equality, the set $\{t\mid t\Rightarrow^{*}c\}$ for any normal-form target $c$ is decidable: compute $\mathrm{normalizeSafe}(t)$ and compare. This is formalized as reachability_decidable in Meta/Newman_Safe.lean: given $c$ in safe normal form, the predicate $t\mapsto\texttt{SafeStepStar}\;t\;c$ is a DecidablePred, reducing reachability to a single normalization followed by equality comparison. Complexity bounds for related TRS classes are discussed in [2, 6, 15].
5 KO7 certification chain
The guarded SafeStep fragment of KO7 is terminating, root-confluent, and normalizable. This section certifies those properties.
5.1 Strong normalization (SafeStep)
The termination argument uses a triple-lexicographic measure
$\mu^{3}_{c}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\tau(t))$
ordered lexicographically over $(\mathbb{N},\,\mathrm{DM}(\mathrm{Multiset}\;\mathbb{N}),\,\mathbb{N})$:
  • $\delta\text{-flag}(t)=1$ if $t=\mathrm{rec}\Delta\,b\,s\,(\delta\,n)$, else 0.
  • $\kappa^{M}(t)$: multiset of natural numbers, recursively defined; each $\mathrm{rec}\Delta(b,s,n)$ contributes $\mathrm{weight}(n)+1$.
  • $\tau(t)$: weighted node-size with per-constructor base weights; $\tau(\delta\,t)=\tau(t)$.
The components are individually standard: lexicographic arguments originate with Floyd [13]; the multiset ordering is due to Dershowitz and Manna [8]; phase bits appear in DP analyses [1]. The contribution is the instantiation to KO7’s eight rules and the Lean 4 mechanization.
Theorem 5.1 (Per-step decrease).
For every $t\Rightarrow t^{\prime}$ in SafeStep, $\mu^{3}_{c}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}_{c}(t)$.
Proof. By rule head. For rec-succ, $\delta\text{-flag}$ drops ($1\to 0$). For rec-zero, $\kappa^{M}$ drops (DM). For tie branches, $\tau$ resolves.
Corollary 5.2 (Strong normalization).
SafeStep is strongly normalizing.
[…]
$\texttt{ctxFuel}(\texttt{rec}\Delta(b,s,n))\;=\;2^{\,\texttt{ctxFuel}(n)\,+\,\texttt{ctxFuel}(s)\,+\,5}+\texttt{ctxFuel}(b)+1\,$,
with additive or constant clauses for the remaining constructors. The constant +5+5 in the exponent is load-bearing: the central lemma ctxFuel_rec_succ_drop shows that the duplicating step $\texttt{rec}\Delta(b,s,\delta(n))\to\texttt{app}(s,\texttt{rec}\Delta(b,s,n))$ strictly decreases ctxFuel, because $\texttt{ctxFuel}(s)+1<2^{\texttt{ctxFuel}(s)+5}$.
Proposition 5.3 (Confluence of SafeStepCtx).
For the partial context closure SafeStepCtx, the remaining global local-join obligation is discharged, so SafeStepCtx is confluent. The exact Newman characterization is still available:
$\texttt{ConfluentSafeCtx}\quad\Longleftrightarrow\quad\forall a,\;\texttt{LocalJoinCtxAt}\;a.$
Since strong normalization is already proved, this gives both an unconditional contextual confluence theorem and the exact Newman-style equivalence for the partial context closure.
Machine-checked in Meta/SafeStepCtx_Confluence.lean as localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, together with the specialized Newman direction newman_safeCtx.
A second confluence result lands one level up from the guarded fragment. The module Meta/EqGuardedConfluence.lean keeps every root rule of Step except the reflexive eqW-diff overlap and proves the resulting intermediate root relation confluent (confluentEqGuarded). This shows that the full-root confluence failure is concentrated exactly at the unguarded eqW peak rather than in the other root rules.
Theorem 5.4 (Full context-closed strong normalization).
The full contextual closure of the unguarded root relation Step is strongly normalizing.
Machine-checked in Meta/ContextClosed_SN_Full.lean as wf_StepCtxFullRev_poly. The proof lifts the nonlinear polynomial witness $W$ through every constructor position and shows that every contextual contraction still strictly decreases $W$.
Proposition 5.5 (Linear derivation-length bound from the global witness).
If $t\Rightarrow_{\mathrm{ctx}}^{n}u$ is a full contextual reduction of the unguarded system with exactly $n$ steps, then $n+W(u)\leq W(t)$. In particular, every such derivation has length strictly below $W(t)$.
Machine-checked in Meta/SafeStep_Complexity.lean as stepCtxFullPow_length_le_W and stepCtxFullPow_length_lt_W. This is a concrete complexity theorem for the full contextual relation; the ordinal discussion below remains a separate proof-theoretic calibration for the guarded SafeStep argument.
5.2 Certified normalizer
A normalization function is defined by well-founded recursion on the measure-pullback relation. It operates on root SafeStep redexes only. Properties:
(Totality) $\forall t\,\exists n.~\mathrm{SafeStepStar}(t,n)\land\mathrm{NormalFormSafe}(n)$
(Soundness) $\forall t.~\mathrm{SafeStepStar}(t,\mathrm{normalizeSafe}(t))\land\mathrm{NormalFormSafe}(\mathrm{normalizeSafe}(t)).$
Note that normalizeSafe computes normal forms with respect to the guarded relation SafeStep, not the full unguarded Step; a SafeStep-normal form may still admit unguarded reductions. Moreover, SafeStep is deterministic at the root (safeStep_deterministic in Meta/NormalizeSafe_LowerBound.lean), so the certified normalizer follows the unique possible guarded root reduction at each step. Under confluence, reachability reduces to normalization and equality check. For safe normal-form targets, the resulting decision procedure also carries an explicit certified cost envelope: Meta/Reachability_Complexity.lean proves reachabilityDecisionCost_le_complexity_bound, together with the linear lower-bound family reachabilityDecision_has_linear_lower_family.
Proposition 5.6 (Exact normalization cost on a merge spine).
Let $m_{0}:=\mathrm{void}$ and $m_{n+1}:=\mathrm{merge}\;\mathrm{void}\;m_{n}$. The certified normalizer returns $\mathrm{void}$ on this family, and the auxiliary counter normalizeSafeSteps, defined by the same deterministic recursion as normalizeSafe, satisfies $\mathrm{normalizeSafeSteps}(m_{n})=n$.
Machine-checked in Meta/NormalizeSafe_LowerBound.lean via mergeVoidChain_pow_void, normalizeSafe_mergeVoidChain, and normalizeSafeSteps_mergeVoidChain. This gives a concrete linear lower bound for the cost of the certified root normalizer on one explicit family.
Proposition 5.7 (Certified root-normalizer upper envelope).
For every term $t$, the deterministic root-step counter normalizeSafeStepstt is realized by an exact guarded root reduction path and is bounded above by the same explicit size-based tower bound used for SafeStepCtx:
$\texttt{normalizeSafeSteps}(t)\;\leq\;\mathrm{complexity\_bound}(\mathrm{termSize}(t)).$
Machine-checked in Meta/SafeRoot_Complexity.lean; the exact theorem identifiers are listed in the claim-to-code appendix. This yields a fully mechanized algorithmic envelope for the certified root normalizer: an exact linear lower-bound family on merge spines and a computable upper bound inherited from the context-closed guarded measure.
In particular, the guarded root relation has a concrete linear lower-bound family for derivation length: each step from $m_{n+1}$ to $m_{n}$ removes exactly one outer mergevoid layer, so the family grows by one constructor layer per step while requiring exactly $n$ guarded root contractions. The current upper-envelope propositions are still coarse: they route the root normalizer and the partial-context closure through ctxFuel and explicit size-indexed envelopes. The remaining open problem is therefore no longer the absence of any upper theorem, but tightening those explicit envelopes toward a sharper slow-growing extraction from the ordinal calibration below.
5.3 Root-step confluence via Newman
We instantiate Newman’s lemma for SafeStep and discharge the local-join hypothesis $\forall\,a,\;\mathrm{LocalJoinAt}\;a$ by case analysis over all root and context shapes. Generic Newman frameworks exist in Lean 4 [22]; the contribution here is the KO7-specific instantiation.
Theorem 5.8 (Confluence and unique normal forms).
The ARS $(\mathrm{Trace},\,\texttt{SafeStep})$ is confluent. Every term reduces to a unique normal form.
Critical-pair coverage
Table 4: Critical-pair coverage at the root for SafeStep.
Source Lemma
$\mathrm{integrate}(\delta\,t)$ localJoin_int_delta
$\mathrm{merge}\;\mathrm{void}\;t$ localJoin_merge_void_left
$\mathrm{merge}\;t\;\mathrm{void}$ localJoin_merge_void_right
$\mathrm{merge}\;t\;t$ localJoin_merge_tt
$\mathrm{rec}\,\Delta\,b\,s\,\mathrm{void}$ localJoin_rec_zero
$\mathrm{rec}\,\Delta\,b\,s\,(\delta\,n)$ localJoin_rec_succ
$\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=\varnothing$ localJoin_of_unique
The side condition $\kappa^{M}(t)=\varnothing$ on R_merge_cancel is part of the guarded confluence design, not a cosmetic strengthening: when $\kappa^{M}(t)\neq\varnothing$, the collapsing root step is blocked, and the proof switches to the blocked-guard case localJoin_merge_cancel_guard_kappa_ne. The collapsing case itself is handled separately by localJoin_merge_tt. This coupling between the termination measure component $\kappa^{M}$ and the confluence guard is a deliberate design choice: the guarded fragment uses the same multiset infrastructure for both termination (the second component of $\mu^{3}_{c}$) and confluence (the merge-cancel guard). In KO7 the same multiset is exactly what distinguishes harmless collapsing cases from forks that would reintroduce the bad payload, so separating the two would duplicate the same structural case split rather than simplify it.
5.4 Ordinal upper-bound calibration
Theorem 5.9 (Upper bound (mechanized)).
The modules Meta/DM_OrderType.lean and Meta/DM_OrderType_LowerBound.lean together prove:
  1. 1.
    $\forall m:\mathrm{Multiset}\;\mathbb{N},\;\texttt{dmOrdEmbed}(m)<\omega^{\omega}$.
  2. 2.
    $\texttt{DM}\;m_{1}\;m_{2}\Rightarrow\texttt{dmOrdEmbed}(m_{1})<\texttt{dmOrdEmbed}(m_{2})$.
  3. 3.
    $\forall t:\mathrm{Trace},\;\texttt{lex3cToOrd}(\mu_{3c}(t))<\omega^{\omega}\!\cdot\!2<\varepsilon_{0}$.
  4. 4.
    $\texttt{SafeStep}\;a\;b\Rightarrow\texttt{lex3cToOrd}(\mu_{3c}(b))<\texttt{lex3cToOrd}(\mu_{3c}(a))$.
Remark 5.10 (Tight order type; mechanized).
The inner product $(\mathrm{DM},\,\mathbb{N})$ has exact order type $\omega^{\omega}$, and the surrounding two-block carrier $(\{0,1\},\,\mathrm{DM}(\mathrm{Multiset}\;\mathbb{N}),\,\mathbb{N})$ has order type $\omega^{\omega}\!\cdot\!2$. Both $\omega^{\omega}\!\cdot\!2$ and $\omega^{\omega+1}$ (for ambient $\mathbb{N}\times(\mathrm{DM}\times\mathbb{N})$) are below $\varepsilon_{0}$. The theorem dm_order_type_omega_omega bundles the embedding, reflection, bound, and surjectivity into one mechanized statement; lex3c_order_type_bound lifts this to the trace bound $\omega^{\omega}\!\cdot\!2$. The exact $\omega^{\omega}$ calibration concerns the DM component itself; the $\omega^{\omega}\!\cdot\!2$ statement is an upper bound for the triple-lex image used by the SafeStep certificate, not an exact order-type claim for SafeStep reductions.
Theorem 5.11 (Lower bound (mechanized, unconditional)).
The module Meta/DM_OrderType_LowerBound.lean proves:
  1. 1.
[…]
  2. 2.
    Surjectivity below $\omega^{\omega}$ for DM embedding.
  3. 3.
    Order reflection into DM; rank-bridge equality dmOrdEmbed_eq_dmRankOrd.
  4. 4.
    Surjectivity of DM rank below $\omega^{\omega}$.
This yields a mechanized order-type isomorphism: $(\mathrm{Multiset}\;\mathbb{N},\mathrm{DM})\cong\omega^{\omega}$ (cf. [8]). The file now packages the exact statement and the cofinal bridge under the theorem names phaseB_cnf_scaffold_exact_order_type and phaseB_cnf_scaffold_cofinal.
Remark 5.12 (Proof-theoretic context of the bound).
The certified upper bound $\omega^{\omega}\!\cdot\!2<\varepsilon_{0}$ places SafeStep derivation lengths within the scope of multiply recursive functions ($F_{\omega}$ in the fast-growing hierarchy [28]), well below the $\varepsilon_{0}$-calibrated systems arising from Hydra battles or Goodstein sequences. This remark is broader and proof-theoretic in character: unlike Proposition 5.5, it does not constitute a tight derivational-complexity bound. Extracting concrete complexity classes from ordinal embeddings requires an explicit complexity-extraction argument [20]. But it confirms that the proof-theoretic strength of the SafeStep termination argument is modest relative to the full ordinal hierarchy.
Proposition 5.13 (Derivation-length bound for SafeStepCtx).
For any term $t$ of structural size $n$, any SafeStepCtx reduction chain from $t$ has length at most $\mathrm{towerBound}(n)$, where $\mathrm{towerBound}$ is the explicit tower-exponential function
$\mathrm{towerBound}(0)=6,\qquad\mathrm{towerBound}(k{+}1)=2^{2\cdot\mathrm{towerBound}(k)+5}+\mathrm{towerBound}(k)+1.$
Machine-checked in Meta/SafeStep_Complexity_Ordinal.lean as safestep_length_bounded_by_size. The proof factors through the computable numeric witness ctxFuel, which strictly decreases on every context-closed safe step. This bound is numerically much cruder than the ordinal calibration at $\omega^{\omega}\!\cdot\!2$: it gives a concrete tower-style majorant rather than a sharp fast-growing or slow-growing extraction. The technical gap is that ctxFuel absorbs context descent by explicit exponentiation, whereas a genuine Moser–Weiermann extraction would need a direct fundamental-sequence control argument for the calibrated ordinal measure itself. A tighter Moser–Weiermann-style descent argument remains future work.
Proposition 5.14 (Fast-growing upper envelope for SafeStepCtx).
Let $F_{0}(x)=x+1$ and $F_{k+1}(x)=F_{k}^{\,x+1}(x)$ be the finite fast-growing hierarchy on naturals. Define
$G(0)=6,\qquad G(k{+}1)=F_{k+2}\!\bigl(2\,G(k)+6\bigr).$
Then every SafeStepCtx reduction chain from a term $t$ has length at most $G(|t|)$, where $|t|$ is the structural size of $t$.
Machine-checked in Meta/SafeStep_Complexity_FastGrowing.lean as fastGrow, fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope, and safestep_length_bounded_by_fgOmegaEnvelope. This does not yet constitute a full Moser–Weiermann slow-growing extraction from the ordinal embedding, but it replaces the old tower-only presentation with an explicit finite-level fast-growing-family envelope aligned with the $\omega^{\omega}$ calibration.
6 External validation: TTT2 and CeTA
We submitted the full KO7 system (8 rules, no SafeStep guards) to TTT2 [17]. TTT2’s MAYBE results are tool search failures, not proofs of impossibility; the YES result below independently corroborates the Lean theorem 5.4. Here the TPDB/TRS semantics is the usual context-closed rewriting semantics, so the CeTA certificate corresponds to Lean’s StepCtxFull, not merely the root relation Step.
Result.
TTT2 proves termination of the context-closed TRS using the DP framework:
  1. 1.
    Dependency pairs extracted: $\mathrm{rec}\Delta^{\sharp}(b,s,\delta(n))\to\mathrm{rec}\Delta^{\sharp}(b,s,n)$, plus two non-cyclic eqW pairs.
  2. 2.
    SCC decomposition: singleton $\{\mathrm{rec}\Delta^{\sharp}\}$.
  3. 3.
    Subterm criterion with $\pi(\mathrm{rec}\Delta^{\sharp})=3$: third argument decreases from $\delta(n)$ to $n$.
Table 5: TTT2 strategy results. CeTA column: independent certification by CeTA 2.36 [24].
Strategy Method class Result Time CeTA
FAST DP + SCC + subterm criterion YES 0.06s certified
COMP DP + TDG + SCC + subterm + matrix + LPO YES 0.70s certified
LPO Lexicographic path order YES 0.02s certified
POLY Polynomial interpretation MAYBE 0.15s rejected
KBO Knuth-Bendix order MAYBE 0.02s rejected
MAT(2) Matrix interpretation dim 2 MAYBE 0.31s rejected
MAT(3) Matrix interpretation dim 3 MAYBE 0.37s rejected
FBI Forward/backward instantiation MAYBE 0.17s n/a
The three YES results use modular or path-order methods; the five MAYBE results use direct/global methods (polynomial, KBO, matrix, forward/backward instantiation). The barrier theorems (Section 4) cover eight scalar numeric direct classes, additive, transparent-compositional, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, and max-plus, together with one symbolic variable-condition barrier, an explicit KBO corollary, and four vector / pair extensions, tracked fixed-dimension componentwise, tracked lexicographic dimension-$2$, balanced mixed-coordinate dimension-$2$ componentwise, and weighted scalar-projection componentwise. The KBO MAYBE outcome is explained directly by Corollary 4.15. The POLY MAYBE outcome is theorem-covered for the formalized direct polynomial families: affine, restricted quadratic, bounded cross-term, bounded multilinear, generalized bounded polynomial, and max-plus are all blocked, while the successful nonlinear witness $W$ lies outside those direct classes by cross-variable coupling and non-transparency. The MAT(d)(d) MAYBE outcomes now split cleanly. Fixed-row tracked-component and row-sum weighted-projection interpretations are theorem-covered by the existing matrix barriers and the explicit corollaries in Meta/MatrixProjectionCoverage.lean; only arbitrary mixed matrix interpretations remain outside the present theorem stack. The FBI MAYBE entry remains outside the theorem layer here: no general forward/backward-instantiation impossibility result is claimed. CPF certificates and CeTA logs are in Artifacts/ttt2/. The Lean module Meta/TPDB_Export.lean renders the exact KO7 TTT2 input and proves that the generated TPDB text matches the artifact text embedded in that module, which in turn corresponds to Artifacts/ttt2/KO7_full_step.trs. The artifact also includes a runnable verifier, lakeexeverifyTpdbExport, that checks the embedded text against the on-disk .trs file. In addition, Meta/TTT2_CertificateReplay.lean replays the mathematical core of the FAST certificate inside Lean: rather than parsing the external CPF, it re-establishes the relevant mathematical content under Lean definitions, namely the single recursive SCC, the projected argument (tool index $2$, paper index $3$), and the resulting well-founded DP pair proof.
Independent certification.
All three YES strategies received a certified verdict from CeTA 2.36. The four MAYBE strategies were correctly rejected. These external tool runs corroborate the method separation predicted by the barrier theorems. They are not, however, part of the formal impossibility proof: the MAYBE outcomes are search failures of bounded heuristic strategies, not theoretical impossibility results.
7 Formalization structure
The Lean 4 artifact separates the kernel definitions, the barrier at the schema, the KO7-specific results, and the certification chain: Lean 4’s Metatheory library [22] provides abstract rewriting, confluence, and strong-normalization infrastructure, but not a library of TRS termination methods, impossibility barriers, or certification bridges. Relative to the established Isabelle/Coq rewriting ecosystems (IsaFoR/CeTA, CoLoR), this project is not another checker for externally generated positive certificates. It is, to our knowledge, the first large Lean 4 development centered on TRS termination methods, fixed-system impossibility boundaries, concrete negative witnesses, and a bridge back to the standard TRS toolchain.
  • Schema barrier layer: additive / transparent-compositional impossibility, affine and nonlinear scalar barriers, matrix-side projection barriers, the scalar-projection meta-theorem, the symbolic variable-condition barrier, and strengthened pumped-subclass corollaries.
  • Operational boundary tools: computable barrier witnesses, including extended extractors for quadratic, max-plus, and projected matrix-side subclasses, a small synthesis-oracle wrapper around them, and a decidable coefficient-table classifier for the formalized scalar and tracked-primary pair families.
  • Growth and transformed-relation infrastructure: reusable successor- and wrapper-growth bridges, a small reusable DP fragment, and the KO7 extracted DP escape.
[…]
  • Certification chain: per-rule decreases and strong normalization for SafeStep, totality and soundness of normalizeSafe, root-step confluence via Newman, and reachability decidability.
  • Complexity and ordinal layers: guarded-context ctxFuel termination, contextual derivation-length bounds, exact merge-spine normalizer cost, and ordinal upper/lower calibration.
  • Catalog, SCC, and ablation layers: theorem-backed depth and pure-precedence families, verified excluded-method results, the delayed-duplication and multiplicity-preserving SCC theorems, the linear-recursion ablation, and the object-level collapse surrogate.
  • Artifact bridge: Lean-side TPDB export for the full KO7 root TRS, an exact-match theorem for the checked TTT2 artifact text, a narrow internal replay of the FAST DP/subterm certificate core, and a runnable verifier for the on-disk .trs file.
Layer Primary declarations Evidence Scope / side conditions
SafeStep wf_SafeStepRev_c, confluentSafe, normalizeSafe_total, reachability_decidable Internal Lean proof Guarded root relation only
SafeStepCtx wf_SafeStepCtxRev, confluentSafeCtx, safestep_length_bounded_by_size, safestep_length_bounded_by_fgOmegaEnvelope Internal Lean proof Partial context closure of the guarded system
Step wf_StepRev_poly, wf_StepRev_mpo Internal Lean proof Full unguarded root relation; separate from the triple-lex certification layer
StepCtxFull wf_StepCtxFullRev_poly, stepCtxFullPow_length_le_W Internal Lean proof Full context closure, derived from the nonlinear witness $W$
TPDB / FAST bridge ko7_full_step_tpdb_matches_artifact_text, ko7FastReplay_sound Lean bridge + external certificate trail External CeTA verdicts corroborate the StepCtxFull result but are not impossibility proofs
The formalization uses three proof shapes: (i) direct full-step orientation (polynomial $W$, KO7-specialized MPO), stated on the unrestricted Step relation; (ii) guarded-fragment certification (triple-lex measure, confluence, normalizer), on SafeStep; and (iii) transformed-relation arguments (dependency pairs), factored through a small reusable fragment for projection-rank descent and SCC-path reasoning. The reusable schema barrier layer is now also exposed through a dedicated public API module, OperatorKO7/SchemaAPI.lean, so downstream developments can import the generic barrier stack and witness/classifier tooling without pulling in the full KO7 certification chain. The support layer is also computational: Meta/BarrierClass_Classifier.lean classifies coefficient tables into the formalized barrier families, Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean return certified violating triples across the scalar and projected matrix-side subclasses, and Meta/SynthesisOracle.lean packages them into a small oracle interface for direct-orientation attempts. All declarations are sorry-free. The termination, normalization, and barrier developments are computationally effective; the ordinal-calibration layer (Meta/DM_OrderType.lean, Meta/DM_OrderType_LowerBound.lean) uses standard classical infrastructure from Lean 4 / Mathlib (Classical.choice, propext) for noncomputable ordinal reasoning. For reviewability, the repository now ships a bidirectional proof-dependency graph in Docs/KO7_BLUEPRINT.md and Docs/ko7_blueprint.json, mapping paper theorem labels to primary Lean declarations and back.
From a proof perspective, the development is not a Lean transcription of the paper statements. Across more than sixty Lean/build files, its main theorem layer consists of impossibility theorems, computed barrier witnesses, boundary classifiers, and a small internal replay of the FAST certificate core, all inside one Lean library, rather than reconstruction of externally generated positive certificates. Mathlib supplies the well-foundedness background for Dershowitz–Manna multisets and the ambient ordinal infrastructure used in Meta/DM_OrderType.lean, but the precise $\omega^{\omega}$ image and lower-bound argument in Meta/DM_OrderType_LowerBound.lean and the full context-closed lift of the nonlinear witness through Meta/ContextClosed_SN_Full.lean are built in this repository. Those two modules are the clearest mechanization bottlenecks: the former because it must recover a sharp ordinal calibration rather than only well-foundedness, and the latter because the polynomial witness has to be threaded through every constructor context together with an explicit derivation-length control.
The paper does not prove arbitrary-target reachability decidability, an exact order type for the full triple-lex image of guarded reductions, or generic MPO/LPO metatheory over arbitrary signatures. It also does not claim that the computable triple-lex certificate handles unrestricted Step, nor that Meta/KBO_Impossible.lean formalizes generic KBO metatheory; the full root Step arguments are instead the nonlinear witness and KO7-specialized MPO proofs, while the KBO corollary is extracted from the symbolic variable-condition abstraction. Nor does it lift the tree-based barrier stack wholesale to higher-order rewriting; the current artifact instead isolates a sharing-aware surrogate in Meta/SharingBarrierLift.lean showing exactly where the no-sharing assumption is load-bearing. Each of those boundaries is explicit in the corresponding sections below.
For reproducibility, the active artifact is pinned to Lean toolchain leanprover/lean4:v4.22.0-rc4 and Mathlib commit 632465e4b02cb70a5dfa4cfe15468e8a62c2bd85; lake-manifest.json freezes the transitive dependency graph. The repository-level note artifact/REPRODUCIBILITY.md documents the minimal replay path, and scripts/make_referee_bundle.py produces a referee-facing bundle containing the frozen Lean snapshot, generated documentation, and archived Artifacts/ttt2/ validation trail together with replay instructions. Separately, artifact/TPDB_SUBMISSION_PREP.md and scripts/stage_tpdb_submission.py stage, but do not perform, a future TPDB submission.
The active artifact (dual-licensed) is at https://github.com/MosesRahnama/OperatorKO7.
8 Related work
Termination of duplicating systems.
The multiset ordering of Dershowitz and Manna [8] handles rules replacing one term by multiple smaller terms. Recursive path orderings [10] extend this to simplification orders. Duplication and termination are surveyed in [25, 3]. Zantema [30] established a hierarchy between termination method classes. Endrullis, Waldmann, and Zantema [11] proved that specific TRSs cannot be handled by matrix interpretations alone, the closest prior art to a formal impossibility result for a specific method.
Interpretation frameworks.
Yamada’s tuple interpretations [29] give a modern umbrella framework subsuming polynomial, matrix, arctic, and related interpretation methods. This is the right taxonomic backdrop for the present barrier stack. Our twelve blocked direct classes are best read as a certified boundary for direct scalar and tracked-tuple fragments inside that broader interpretation framework, while the successful escape mechanisms in this paper either move to transformed relations (dependency pairs) or to global witnesses that are not members of the blocked direct fragments.
Dependency pairs and modular termination.
The dependency pair method [1] was introduced to overcome limitations of direct simplification orders. Hirokawa and Middeldorp [14] introduced the subterm criterion. Thiemann and Giesl showed size-change termination for TRSs captures only simple termination, while DP goes beyond [26]. IsaFoR/CeTA [24] provides machine-checked formalizations of termination methods in Isabelle; CoLoR [4] and Coccinelle/CiME [7] in Coq. These projects certify positive termination proofs or their soundness conditions. Our artifact occupies a different proof-engineering niche: it internalizes impossibility theorems for direct classes on a fixed terminating system, adds executable witness extraction and classification around those theorems, and only then reconnects to the external TRS toolchain through certified artifacts.
Path orders and proof-theoretic strength.
Ferreira and Zantema [12] showed LPO well-foundedness depends on the subterm property. Buchholz [5] and Dershowitz-Okada [9] analyzed path orders as requiring Kruskal’s Tree Theorem. Rathjen and Weiermann [23] calibrated the resulting general path-order metatheory at $\Pi^{1}_{1}$-CA0. This concerns well-foundedness results for path orders over arbitrary signatures and trees. It does not imply that every fixed-signature specialization must internally reproduce that strength: our Lean MPO result for KO7 is a direct ordinal-ranking proof for the fixed KO7 signature, precedence, and same-head clause. The DP subterm criterion [14] does not require Kruskal-strength principles; its soundness is formalized in IsaFoR at strength below PA.
Confluence and mechanization platforms.
Newman’s lemma [21] has been mechanized in Isabelle/HOL [24] and Lean 4 [22]. Our application is to a root ARS with a discharged local-join hypothesis. The TRS certification ecosystem (IsaFoR/CeTA in Isabelle, CoLoR in Coq, the Ramos library [22] in Lean 4) certifies positive termination proofs but not impossibility theorems for measure classes. IsaFoR also certifies nontermination via loop witnesses [24]; Kim, Saito, Thiemann, and Yamada [16] formalized co-rewrite pairs for non-reachability. Both certify properties of specific systems rather than statements about measure classes. Our contribution is therefore not another certificate checker or soundness layer, but a theorem-level boundary analysis with computed outputs that still connect back to the standard TRS toolchain. The distinction from [19, 27] is noted in §1.
9 Conclusion
We formalized an orientation boundary for first-order rewrite systems with step-duplicating recursors. Twelve direct measure classes are excluded by schema-level impossibility theorems (Theorems 4.1 to 4.13); the escape side is classified by the DP clause, the explicit DP base-order boundary, transparency essentiality, the generalized bounded-polynomial escape condition, the escape trichotomy, and the two SCC synchronization theorems; and the concrete witness calculus KO7 carries a guarded certification chain together with full-step polynomial and MPO witnesses.
The theory is explanatory as well as negative. The barrier stack separates baseline additive/compositional failure, widened scalar families, vector / pair projection layers, and the symbolic variable-condition side; the escape results match that organization closely. It is a boundary theorem for explicit direct families, not a claim that KO7 defeats termination methods wholesale. Dependency pairs, LPO, the nonlinear witness, and the KO7-specialized MPO remain on the successful side of the boundary, and the repository makes those imports explicit at the theorem level.
Open directions include extending the projection-based trichotomy layer to broader mixed matrix and tropical families, tightening the root-normalizer and SafeStepCtx upper envelopes from the current explicit tower and fast-growing-family bounds toward a sharper slow-growing extraction from the $\omega^{\omega}\!\cdot\!2$ calibration, replaying a meaningful fragment of the external certificate chain inside Lean, and lifting the barrier analysis to higher-order or sharing-based rewriting semantics.
Appendix A Addendum
A.1 Lean kernel location
The mechanized kernel is OperatorKO7/Kernel.lean at https://github.com/MosesRahnama/OperatorKO7. Under OperatorKO7/Meta/ lie the safe fragment and its proofs.
A.2 Module map
Table LABEL:tab:module-map lists the theorem-bearing kernel and meta modules in the active artifact. Lightweight test files such as Test/Sanity.lean are intentionally omitted.
Table 6: Lean module map. Key theorem identifiers are shown in parentheses.
Module Contents
Kernel.lean Trace (7 constructors), Step (8 rules), StepStar
Meta/SafeStep_Core weight, kappaM, deltaFlag, guarded SafeStep
Meta/ComputableMeasure tau, Lex3c, per-rule decrease (wf_SafeStepRev_c)
Meta/ComputableMeasure_Verification Sanity checks (deltaFlag_binary, no_infinite_safestep_chain)
Meta/Normalize_Safe Certified normalizer; totality, soundness (normalizeSafe_sound)
Meta/NormalizeSafe_LowerBound Auxiliary step counter for the certified normalizer; exact merge-spine cost family (normalizeSafeSteps_mergeVoidChain)
Meta/SafeStep_Ctx Context closure SafeStepCtx
Meta/ContextClosed_SN ctxFuel measure for the guarded partial context closure (wf_SafeStepCtxRev)
Meta/ContextClosed_SN_Full Full unguarded context closure; polynomial witness lifted through all constructor positions (wf_StepCtxFullRev_poly)
Meta/ContextClosedBarrier Root-to-context barrier survival corollaries for StepCtxFull (stepCtxFull_orientation_implies_root,
no_stepCtxFull_orientation_affine_with_pump)
Meta/SafeStep_Complexity Exact-length contextual derivations; global $W$-based derivation-length bound (stepCtxFullPow_length_le_W)
Meta/SafeStep_Complexity_Ordinal SafeStepCtx derivation-length bound via ctxFuel; size-indexed tower-exponential complexity bound (safestep_length_bounded_by_size)
Meta/SafeStep_Complexity_FastGrowing Explicit finite fast-growing envelope majorizing the same SafeStepCtx length bound (fgOmegaEnvelope, safestep_length_bounded_by_fgOmegaEnvelope)
Meta/SafeRoot_Complexity Exact-path realization and explicit upper envelope for the certified root normalizer (normalizeSafeSteps_realized,
normalizeSafeSteps_le_complexity_bound)
Meta/Reachability_Complexity Explicit decision-cost envelope for safe-normal-form reachability, together with a linear lower-bound family (reachabilityDecisionCost_le_complexity_bound,
reachabilityDecision_has_linear_lower_family)
Meta/SafeStepCtx_Confluence Discharged global contextual local-join obligation and unconditional SafeStepCtx confluence, retaining the exact Newman equivalence (localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll)
Meta/EqGuardedConfluence Intermediate root fragment keeping all root rules except the reflexive eqW-diff overlap; root confluence restored (confluentEqGuarded)
Meta/Confluence_Safe Local-join lemmas (localJoin_all_safe)
Meta/EqW_Guard_Barrier Full-step eqW overlap obstruction; guard-necessity theorem (not_localJoinStep_eqW_refl)
Meta/Newman_Safe Newman instantiation; confluentSafe; unique NFs; reachability_decidable
Meta/Impossibility_Lemmas Failed-measure catalog; historical dead-end notes moved to Notes/Impossibility_DeadEnds.md
Meta/Conjecture_Boundary No-go theorems, GlobalOrients framework
Meta/StepDuplicatingSchema Parametric schema; generic impossibility; affine/linear barrier; projection-rank escape (Thms. 4.14.2;
no_additive_orients_dup_step)
Meta/BarrierWitness Computable certificate extractors: additive_witness, compositional_witness, affine_witness
Meta/BarrierWitness_Extended Extended witness extractors for pumped restricted-quadratic, max-plus, and projected matrix-side families (quadratic_with_pump_witness,
max_with_pump_witness, matrixFunctional_with_projected_affine_pump_witness)
Meta/AffineThresholdSharpness Canonical affine family showing exactness of the generic affine pump bound (affineThresholdMeasure_bound,
affineThresholdMeasure_exact_cutoff)
Meta/SynthesisOracle Small synthesis-oracle wrapper around the barrier certificates, with a local refinement-change invariant (SynthesisOracle, successful_refinement_changes_certificate)
Meta/TTT2_CertificateReplay Narrow Lean-side replay of the FAST certificate core: SCC count, projected argument, and resulting DP-pair well-foundedness (ko7FastReplay_sound, wf_DPPairRev_replayed)
Meta/SharingBarrierLift Sharing-aware surrogate showing exactly where the tree-style direct barrier stops applying (sharedCounter_orients_step, sharing_breaks_tree_barrier)
Meta/TPDB_Export Lean-side TPDB export for the full KO7 root TRS; exporter output equals the embedded artifact text literal (ko7_full_step_tpdb_matches_artifact_text)
Meta/CompositionalMeasure_Impossibility KO7 instantiation; KO7 DP witness; KO7 affine barrier (Thms. 4.184.21;
no_additive_compositional_orients_rec_succ)
Meta/QuadraticBarrier Restricted degree-$2$ barrier without step-counter coupling; KO7 specializations (no_quadratic_counter_orients_dup_step_of_unbounded, no_global_step_orientation_quadratic_of_succ_pump)
Meta/QuadraticCrossTermBarrier Bounded step-counter cross-term quadratic barrier; KO7 specializations under the wrapper-dominance side condition (CrossTermBoundedAtBase,
no_cross_quadratic_orients_dup_step_of_unbounded,
no_global_step_orientation_cross_quadratic_of_succ_pump)
Meta/MultilinearBarrier Finite multilinear direct-polynomial barrier with frozen-coefficient dominance; successor- and wrap-pump corollaries (MultilinearDominatedAtBase, no_multilinear_orients_dup_step_of_unbounded)
Meta/PolynomialBarrierGeneral Generalized bounded direct-polynomial barrier with repeated variables and a necessary frozen base-dominance failure condition (no_polynomial_orients_dup_step_of_unbounded,
polynomial_escape_requires_failure_of_base_dominance)
Meta/MaxBarrier Schema-level max-plus barrier with unbounded, successor-pump, and wrap-pump forms (no_max_orients_dup_step_of_unbounded, no_global_step_orientation_max_of_succ_pump)
Meta/ArcticBarrier Arctic-style primary-projection corollary of the max barrier (no_arctic_primary_orients_dup_step_of_unbounded,
no_global_step_orientation_arctic_primary_of_unbounded)
Meta/TropicalBarrier Broader tropical primary-projection barrier reducing finite tropical coordinates to the same scalar obstruction (no_tropical_primary_orients_dup_step_of_unbounded,
no_global_step_orientation_tropical_primary_of_unbounded)
Meta/PumpedBarrierClasses Strengthened pumped subclasses yielding unconditional affine / quadratic / bounded cross-term / generalized bounded-polynomial / multilinear / max-plus / projection-based matrix corollaries (PolynomialMeasureWithPump,
MultilinearMeasureWithPump, MaxMeasureWithPump,
MatrixFunctionalMeasureWithProjectedAffinePump)
Meta/StandardPumpLemmas Reusable successor-/wrapper-growth lemmas and subclass constructors bridging conditional barriers to pumped subclasses (affine_hasUnboundedRange_of_succ_pump, QuadraticCounterMeasure.withWrapPump, MatrixMeasure2.withPrimarySuccPump)
Meta/EscapeTrichotomy Escape trichotomy for the explicit KO7 direct universe, plus the projection-based extension for weighted functional and balanced mixed-coordinate families (ko7_nat_direct_escape_trichotomy, ko7_direct_escape_trichotomy_extended,
ko7_projection_escape_trichotomy)
Meta/DepthBarrier KO7-specific max-aggregative depth family; dedicated depth barrier theorem (MaxDepthMeasure, no_global_step_orientation_maxDepth)
Meta/PrecedenceBarrier KO7-specific pure head-precedence family; dedicated precedence barrier theorem (HeadPrecedenceFamily, no_global_step_orientation_headPrecedenceFamily)
Meta/ObjectAxiom_Ablation Directed wrapper collapse rule $\mathrm{app}(t,t)\to t$; collapse sanity and preserved direct barriers (wf_StepCollapseRev, no_global_stepCollapse_orientation_additive)
Meta/MutualDuplication_General Bounded two-node SCC theorem; delayed duplication with additive and affine composite-step barriers (alternating_dup2_realized, no_affine_orients_alternating_dup2_composite_of_unbounded)
Meta/MutualDuplication_Case Concrete alternating worked instance of the bounded SCC theorem (alternating_dup2_realized, no_additive_orients_alternating_dup2_step)
Meta/MutualDuplication_Preserving Multiplicity-preserving synchronized SCC barrier; exact payload-count preservation with additive and conditional affine synchronized-exposure obstructions (step_preserves_payloadCount, no_additive_orients_synchronized_cycle, no_affine_orients_synchronized_cycle_of_wrapper_dominance)
Meta/MutualDuplication_Transparent Transparent-compositional and projected matrix-style extension of the alternating SCC barrier (no_global_orients_ctx_compositional_transparent, no_global_orients_ctx_matrixFunctional_of_projected_unbounded)
Meta/MatrixBarrierD Fixed-dimension tracked componentwise vector barrier; KO7 specialization for arbitrary finite dimension (no_matrixD_orients_dup_step_of_componentwise_pump, no_global_step_orientation_matrixD_of_componentwise_pump)
Meta/MatrixBarrier2 Dimension-22 worked instance of the tracked componentwise barrier, plus symmetric second-component lemmas (no_matrix2_orients_dup_step_of_componentwise_pump, no_matrix2_orients_dup_step_of_componentwise_pump_snd)
Meta/MatrixBarrierLex Dimension-22 lexicographic affine pair barrier; tracked primary-component KO7 specialization (no_matrix2_lex_orients_dup_step_of_unbounded_primary, no_global_step_orientation_matrix2_lex_of_componentwise_pump)
Meta/MatrixBarrierMix2 Balanced mixed-coordinate dimension-$2$ pair barrier via aggregate-sum projection (no_matrixMix2_orients_dup_step_of_sum_pump, no_global_step_orientation_matrixMix2_of_sum_pump)
Meta/MatrixBarrierFunctional Weighted scalar-projection componentwise barrier; tracked-coordinate and aggregate-sum projection unification (no_matrixFunctional_orients_dup_step_of_componentwise_pump,
no_global_step_orientation_matrixFunctional_of_componentwise_pump)
Meta/MatrixProjectionCoverage Explicit fixed-row and row-sum corollaries exposing matrix-side projection coverage for tool-facing discussion (no_global_step_orientation_matrix_fixed_row_of_componentwise_pump,
no_global_step_orientation_matrix_row_sum_of_componentwise_pump)
Meta/ScalarProjectionBarrier Generic scalar-projection lift subsuming several projection-based matrix barriers (no_orients_dup_step_of_scalar_projection,
no_matrixMix2_orients_dup_step_of_sum_pump_via_projection)
Meta/SymbolicComparatorBarrier Symbolic variable-condition obstruction for direct KBO-style comparators on the duplicating rule schema (not_orients_dup_rule,
no_symbolic_variable_condition_orients_dup_step)
Meta/KBO_Impossible Explicit KO7 KBO-style impossibility corollary extracted from the symbolic barrier (no_kbo_orients_ko7_rec_succ)
Meta/BarrierClass_Classifier Decidable coefficient-table classifier for the formalized scalar and tracked-primary pair barrier families (ScalarCoeffTable.structuralClasses, PrimaryPairCoeffTable.barrierClass)
Meta/TypedBarrierSurvival Simply-typed first-order recursor fragment; additive survival and affine survival under a typed step pump (no_additive_orients_typed_recSucc,
no_affine_orients_typed_recSucc_of_stepPump)
Meta/DependencyPairs_Fragment Reusable DP fragment: rank-based pair problems and SCC-path descent (DPProjection.wfRev, SCCCycle.target_lt_source)
Meta/DependencyPairs_Works Extracted DP pair; projection rank (wf_DPPairRev)
Meta/DP_BaseOrder_Boundary Boundary theorem showing the extracted DP pair already admits a simple linear polynomial-style base order (counterHeight_orients_dpPair,
no_blanket_dp_poly_base_barrier)
[…]
Meta/DM_OrderType DM-to-ordinal embedding; $\varepsilon_{0}$ bridge (Thm. 5.9; dmOrdEmbed_lt_opow_omega)
[…]
Meta/LinearRec_Ablation Linear recursor; simpleSize orientation (§4.6; wf_linearStep)
Meta/MPO_FullStep KO7-specialized MPO: per-rule orientation; reverse-MPO WF (wf_StepRev_mpo)
Meta/MPO_ProofTheoreticBound Explicit ordinal envelope for the KO7-specialized MPO ranking (mpoOrd_lt_mpoVeblenBound, wf_StepRev_mpo_below_veblen7)
Meta/MPO_Precedence_Barrier Explicit bad-precedence obstruction for the KO7 MPO shape (not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec)
Meta/PolyInterpretation_FullStep Nonlinear polynomial $W$; axiom-violation witnesses (Rem. 4.20;
wf_StepRev_poly)
A.3 Relation/property status matrix
Relation / layer Termination Confluence / normalization Complexity / ordinal Key Lean anchors
Step Root termination: yes Root confluence: no (witness); normalizer: n/a nonlinear polynomial and MPO witnesses; MPO ranking bounded below $\varphi_{7}(0)$ wf_StepRev_poly, wf_StepRev_mpo, wf_StepRev_mpo_below_veblen7, not_localJoinStep_eqW_refl
SafeStep Strong normalization: yes Root confluence + certified normalizer + reachability to safe normal-form targets triple-lex ordinal upper-bound calibration; explicit root-normalizer lower/upper envelope wf_SafeStepRev_c, confluentSafe, normalizeSafe_sound, reachability_decidable, normalizeSafeSteps_le_complexity_bound, safeMeasure_below_epsilon0
SafeStepCtx Partial-context SN: yes Confluence: yes; exact Newman equivalence retained size-indexed tower bound plus explicit fast-growing-family envelope wf_SafeStepCtxRev, localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, safestep_length_bounded_by_size, safestep_length_bounded_by_fgOmegaEnvelope
StepCtxFull Full-context SN: yes Confluence: not claimed; normalizer: n/a explicit $W$-based length bound wf_StepCtxFullRev_poly, stepCtxFullPow_length_le_W
DP transform Extracted DP problem terminates n/a projection-rank descent only dp_projection_orients_rec_succ, dp_projection_violates_sensitivity
A.4 Claim-to-code index
Tables LABEL:tab:claim-code-barrier and LABEL:tab:claim-code-certification index the numbered paper results by Lean identifier, abstraction level, and side conditions. This complements the module map: the tables below are organized by paper statement rather than by file.
Table 7: Barrier, escape, and frontier results indexed by paper label.
Paper label Lean identifier(s) Level Conditions
Thm. 4.1 no_additive_orients_dup_step, no_compositional_orients_dup_step_transparent_succ Schema Tier 1: positive wrap weight. Tier 2: wrap-subterm axioms and transparency at base.
Thm. 4.2 no_affine_orients_dup_step_of_unbounded, no_affine_orients_dup_step_of_succ_pump, no_affine_orients_dup_step_of_wrap_pump Schema Unbounded range, or certified successor / wrap pump.
Thm. 4.3 no_quadratic_counter_orients_dup_step_of_unbounded Schema Unbounded range, or certified successor / wrap pump.
Thm. 4.4 CrossTermBoundedAtBase, no_cross_quadratic_orients_dup_step_of_unbounded Schema Unbounded range plus bounded base coupling, or certified successor / wrap pump.
Thm. 4.5 MultilinearDominatedAtBase, no_multilinear_orients_dup_step_of_unbounded, no_global_step_orientation_multilinear_of_succ_pump Schema / KO7 Finite multilinear monomial table, unbounded range, and frozen pumped-variable coefficient dominated at the base point.
Thm. 4.6 no_polynomial_orients_dup_step_of_unbounded, no_global_step_orientation_polynomial_of_succ_pump,
polynomial_escape_requires_failure_of_base_dominance
Schema / KO7 Finite degree-bounded polynomial table with arbitrary exponents; unbounded or pumped growth; eventual frozen target dominance at the base point.
Thm. 4.7 no_max_orients_dup_step_of_unbounded, no_global_step_orientation_max_of_succ_pump Schema / KO7 Max-plus constructor-local interface with unbounded range, or positive successor / left-wrapper drift.
Cor. 4.8 no_arctic_primary_orients_dup_step_of_unbounded, no_global_step_orientation_arctic_primary_of_unbounded Schema / KO7 Distinguished finite arctic primary projection satisfying the max-plus interface.
Thm. 4.9 no_matrixD_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with one tracked affine coordinate and growth in that coordinate.
Thm. 4.10 no_matrix2_lex_orients_dup_step_of_unbounded_primary Schema / KO7 Lexicographic order by the first component; unbounded or pumped primary coordinate.
Thm. 4.11 no_matrixMix2_orients_dup_step_of_sum_pump Schema / KO7 Balanced mixed $2\times 2$ constructor maps and an unbounded aggregate-sum pump.
Thm. 4.12 no_matrixFunctional_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with a fixed nonzero natural-weighted scalar projection satisfying the affine interface.
Thm. 4.13 no_orients_dup_step_of_scalar_projection,
no_matrix2_orients_dup_step_of_componentwise_pump_via_projection,
no_matrixMix2_orients_dup_step_of_sum_pump_via_projection
Schema meta-layer A codomain order whose strict decrease forces strict decrease of an already-blocked scalar projection.
Thm. 4.14 not_orients_dup_rule, no_symbolic_variable_condition_orients_dup_step Symbolic schema Standard variable condition on the strict symbolic comparator.
Cor. 4.15 no_kbo_orients_ko7_rec_succ KO7 symbolic corollary KBO-style order satisfying the standard variable condition.
Cor. 4.16 no_affine_with_pump_orients_dup_step,
no_quadratic_with_pump_orients_dup_step,
no_cross_quadratic_with_pump_orients_dup_step,
no_matrix2_with_primary_pump_componentwise_orients_dup_step,
no_matrix2_with_primary_pump_lex_orients_dup_step
Strengthened subclasses Unconditional inside the subclass definitions.
Cor. 4.17 stepCtxFull_orientation_implies_root, no_stepCtxFull_orientation_affine_with_pump,
no_stepCtxFull_orientation_matrixFunctional_with_projected_affine_pump
Context-closed corollary layer Identity-context embedding of root steps into StepCtxFull.
Thm. 4.18 no_additive_compositional_orients_rec_succ, no_compositional_orients_rec_succ_transparent_delta KO7 Tier 1: positive app weight. Tier 2: app-subterm axioms and delta-transparency at void.
[…]
Thm. 4.21 dp_projection_orients_rec_succ, dp_projection_violates_sensitivity KO7 / DP None.
Prop. 4.22 counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order,
no_blanket_dp_poly_base_barrier
KO7 / DP Extracted DP pair problem; linear polynomial-style counter-height base order.
Thm. 4.23 alternating_dup2_realized,
no_additive_orients_alternating_dup2_composite,
no_affine_orients_alternating_dup2_composite_of_unbounded
Alternating SCC schema Additive branch: none. Affine branch: unbounded pump.
Thm. 4.24 step_preserves_payloadCount,
synchronized_cycle_realized,
synchronized_cycle_exposes_payload,
no_additive_orients_synchronized_cycle,
no_affine_orients_synchronized_cycle_of_wrapper_dominance
Synchronized SCC Additive: synchronized tagged payload inputs. Affine: wrapper-dominance condition + unbounded pump.
Cor. 4.25 no_global_orients_additive, no_global_orients_compositional_transparent_succ, no_global_orients_affine_of_unbounded Schema / global Affine branch inherits the unbounded-pump condition.
Prop. 4.26 W_orients_step, wf_StepRev_poly KO7 root Step None.
Prop. 4.27 wf_StepRev_mpo KO7 root Step Fixed signature, fixed precedence, fixed same-head clause.
Prop. 4.28 mpoOrd_lt_mpoVeblenBound, wf_StepRev_mpo_below_veblen7 KO7 root Step Fixed signature; ordinal envelope below Ordinal.veblen70.
Prop. 4.29 not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec KO7 root Step Bad-precedence hypothesis.
Thm. 4.30 ko7_direct_escape_trichotomy_extended (preliminary: ko7_nat_direct_escape_trichotomy; projection extension: ko7_projection_escape_trichotomy) KO7 direct universe Successful orienter must violate wrapper sensitivity, violate transparency, or leave the twelve explicit representable families.
Prop. 4.31 not_localJoinStep_eqW_refl KO7 root Step None.
Table 8: Certification and ordinal-calibration results indexed by paper label.
Paper label Lean identifier(s) Level Conditions
Thm. 5.1 measure_decreases_safe_c SafeStep Guarded root step.
Cor. 5.2 wf_SafeStepRev_c SafeStep None.
Prop. 5.3 localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, newman_safeCtx SafeStepCtx Strong normalization already available; the contextual local-join obligation is discharged, yielding unconditional confluence.
Thm. 5.4 wf_StepCtxFullRev_poly StepCtxFull None.
Prop. 5.5 stepCtxFullPow_length_le_W, stepCtxFullPow_length_lt_W StepCtxFull Exact-length contextual reduction hypothesis.
Prop. 5.13 safestep_length_bounded_by_size SafeStepCtx Size-indexed tower-exponential bound via ctxFuel.
Prop. 5.14 fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope,
safestep_length_bounded_by_fgOmegaEnvelope
SafeStepCtx Explicit finite fast-growing-family envelope majorizing the same contextual length bound.
Prop. 5.6 mergeVoidChain_pow_void, normalizeSafe_mergeVoidChain, normalizeSafeSteps_mergeVoidChain SafeStep normalizer Merge-spine family.
Prop. 5.7 normalizeSafeSteps_realized, normalizeSafeSteps_le_ctxFuel,
normalizeSafeSteps_le_complexity_bound, normalizeSafeSteps_has_linear_lower_family
SafeStep normalizer Exact guarded root path; upper bound factored through ctxFuel / complexity_bound.
Thm. 5.8 localJoin_all_safe, confluentSafe SafeStep None.
Thm. 5.9 dmOrdEmbed_lt_opow_omega, safeMeasure_strictMono_explicit, safeMeasure_below_epsilon0 Ordinal calibration Guarded step for the strict-decrease clause.
Thm. 5.11 phaseB_cnf_scaffold_exact_order_type, phaseB_cnf_scaffold_cofinal,
dmOrdEmbed_eq_dmRankOrd, lex3c_order_type_bound
DM / ordinal scaffold None.
A.5 Local-join report (confluence)
Table 9: Local-join witnesses for the safe relation.
Type Shape Join lemma name Notes
Root merge void void localJoin_merge_void_void both branches to void
Root merge t t localJoin_merge_tt joins at $t$
Root integrate (delta t) localJoin_int_delta unique target void
[…]
Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
Root eqW a a, $\kappa^{M}(a)\neq 0$ localJoin_eqW_refl_guard_ne vacuous join
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root merge
Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
Ctx eqW a a, $\kappa^{M}(a)\neq 0$ localJoin_ctx_eqW_refl_guard_ne vacuous join
Ctx eqW a a (norm + guards) localJoin_ctx_eqW_refl_if_normalizes_to_delta via normalizeSafe