### 08 Apr 2026 v8

# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

Abstract
We formalize the orientation boundary for first-order step-duplicating recursors. At the schema level, no measure in twelve formalized direct classes in the barrier package can strictly orient the duplicating step uniformly. These classes include additive and transparent-compositional measures, several scalar growth families, and tracked vector / pair families under stated pump or growth hypotheses. Two meta-results sharpen that stack: a scalar-projection theorem for the matrix-side barriers, and a symbolic variable-condition barrier yielding a KBO-style impossibility corollary for the concrete witness system studied later. To our knowledge, this is the first mechanized object-level barrier theorem, proved without reductions or undecidability arguments, showing that a fixed terminating system lies outside several named direct termination-measure classes. All internal proofs are formalized in Lean 4.
The paper also isolates three escape routes inside the explicit direct universe and extends that analysis on the matrix side by scalar-projection representability. It proves a transparency-essentiality theorem with a concrete witness, a dependency-pair escape characterization, a generalized polynomial boundary theorem isolating failure of frozen base dominance, computable barrier-witness extractors with explicit construction budgets, a coefficient-table classifier, a small synthesis-oracle layer, ablations, and two SCC synchronization families that continue from bounded witness systems to finite cyclic and raw-graph forms. Within this explicit direct universe, every successful orienter succeeds only by importing structural assumptions outside the formalized barrier classes.
The concrete witness calculus KO7 carries a two-layer formal development. For the guarded fragment we internally certify strong normalization, confluence, a certified normalizer, and decidable reachability to safe normal-form targets. We also prove a single-exponential SafeStepCtx derivation-length theorem with a matching explicit lower family, an exact DM-component ordinal calibration at $\omega^{\omega}$, a triple-lex certificate image below $\omega^{\omega}\cdot 2$, and explicit realized lower families in both blocks of that triple-lex image. For the full unguarded system we separately prove root termination by a nonlinear polynomial witness and by a KO7-specialized MPO, lift the nonlinear witness to context-closed strong normalization with an explicit derivation-length bound, and connect the Lean development to TTT2/CeTA through a checked TPDB export, a runnable verifier bridge, and a narrow internal replay of the FAST certificate core.
The Lean 4 artifact is available at https://github.com/MosesRahnama/OperatorKO7.
1 Introduction
Consider a first-order rewrite rule of the form
$\mathrm{recur}(b,\,s,\,\mathrm{succ}(n))\;\to\;\mathrm{wrap}(s,\,\mathrm{recur}(b,\,s,\,n))$
where the step argument $s$ appears once on the left and twice on the right. Any direct aggregative termination measure that aggregates subterm contributions and respects the wrapper constructor must account for both copies of $s$ on the right-hand side. Once the measured value of $s$ exceeds the fixed successor-side contribution available in the chosen class, the extra copy prevents a uniform strict decrease. Modular methods (dependency pairs with subterm criterion) escape by projecting to the recursion counter and discarding the duplicated mass.
The central question is: which classes of termination measures can orient this rule, and which cannot? We call the dividing line between these classes the orientation boundary. On one side sit measure classes that provably fail; on the other, methods that succeed by importing structural assumptions outside those classes.
This paper formalizes both sides of this boundary. We prove impossibility theorems at the schema for twelve axiomatized direct measure classes (the barrier package, enumerated below), classify the escape mechanisms, and validate the boundary on a concrete system, KO7, carrying a full certification chain. These twelve classes are not ad hoc: they are the direct scalar and tracked-tuple fragments of the broader tuple-interpretation taxonomy discussed later in Section 4. A separate twelve-family universe drives the escape trichotomy in Section 4; the two are reconciled there.
Prior work
The observation that additive and KBO-style direct measures cannot orient duplicating rules is classical: KBO cannot orient any duplicating rule because of the variable condition [4], and the dependency pair framework was introduced to handle systems where simplification orderings fail [2]. Middeldorp and Zantema [19] established the hierarchy between simple termination and full termination. Our contribution is to axiomatize a broader transparency-constrained compositional class together with a separate affine/linear constructor-local class, and then prove barrier theorems, an escape characterization, a transparency-essentiality theorem, and a mechanized ablation for those classes. To our knowledge, no prior work formalizes impossibility theorems for named termination-measure classes on a fixed terminating system in the strong object-level sense used here: no member of the specified class can uniformly orient the duplicating step of that system. Mitterwallner, Middeldorp, and Thiemann [20, 28] formalized in Isabelle/HOL the undecidability of orientation by linear polynomial interpretations and by KBO with subterm coefficients. That is a result about the decision problem across all inputs. Their work does not produce standalone impossibility theorems for named systems; ours does. The broader formalization ecosystem (IsaFoR/CeTA, CoLoR) certifies positive termination proofs but not impossibility theorems for measure classes.
The closest neighboring lines of work address different questions. The LICS/AFP undecidability line asks whether a method class is algorithmically decidable across all inputs; certificate ecosystems such as IsaFoR/CeTA and CoLoR ask whether a successful external termination argument can be reconstructed and checked. Our claim is narrower: for one fixed terminating system, we prove object-level impossibility theorems for explicit direct measure families, explain the escape routes, and package constructive witnesses and classifiers around that boundary. The difference is architectural as well as mathematical. The artifact is not another generic certificate checker. It develops a negative boundary theory for one fixed terminating system and then reconnects that Lean development to standard TRS tooling through a checked TPDB export and verifier bridge.
Correspondingly, the paper does not claim that KO7 defeats termination methods in general. The no-go theorems concern the twelve direct measure families of the barrier package and an adjacent direct symbolic barrier. Modular methods and fixed-signature path-order arguments remain on the successful side of the boundary, and the artifact proves exactly that through the DP escape, the nonlinear full-step witness $W$, and the KO7-specialized MPO. KO7 itself is deliberately minimal rather than ad hoc: it is a witness calculus designed to exhibit, in one fixed first-order system, the four structural features that make the boundary stack load-bearing, namely duplication, visible wrappers, a transparent-successor interface on the direct side, and enough auxiliary structure to expose both confluence and certification behavior.
Contributions
All results below are machine-checked in Lean 4 unless otherwise noted.
A. The orientation barrier (primary contribution). The main theorem package is a twelve-class direct boundary (the barrier package) for the duplicating step, with the latter ten classes blocked under their stated pump or growth hypotheses and the matrix side unified by a scalar-projection meta-theorem. The count is $2+6+4=12$: two unconditional classes (additive and transparency-constrained compositional), six scalar growth classes (affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus), and four tracked vector/pair families (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). Two structural results sit above that stack: a scalar-projection meta-theorem (Theorem 4.14) unifying the projection-based matrix arguments, and a symbolic variable-condition barrier (Theorem 4.15) together with an explicit KO7 KBO corollary (Corollary 4.16) for direct KBO-style comparators. The generalized bounded-polynomial theorem also now carries a WPO-facing direct polynomial-branch corollary (Corollary 4.7) for orders whose strict comparison is certified by such a bounded polynomial algebra. Meta/PumpedBarrierClasses.lean supplies unconditional strengthened-subclass corollaries for the affine, restricted-quadratic, bounded cross-term quadratic, generalized bounded-polynomial, tracked-primary pair, bounded multilinear, max-plus, weighted-functional projected matrix, and balanced mixed-coordinate sum-pump families. The affine contradiction cutoff is also shown exact for a canonical family in Meta/AffineThresholdSharpness.lean. Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean provide concrete violating triples for the additive and transparent-compositional classes, a packaged affine contradiction once a pump term is supplied, and additional certified extractors for the restricted-quadratic, max-plus, and projected matrix-side pumped subclasses; Meta/BarrierWitness_Budgets.lean packages the corresponding canonical construction budgets, showing that the extracted payload is always base, succIter $k$, or wrapIter $k$ for an explicit computable budget $k$.
B. Escape characterization and precision. A DP escape clause (Theorem 4.22), a transparency-essentiality theorem with a concrete witness (Theorem 4.20), an explicit DP base-order boundary note (Proposition 4.23), a generalized degree-bounded polynomial barrier with a necessary base-dominance failure condition (Theorem 4.6), and an escape trichotomy over an explicit direct universe, extended on the matrix side by projected representability (Theorem 4.31), turn the no-go side into an explicit escape theory. Two ablations and two SCC synchronization theorems sharpen the boundary: removing step-duplication dissolves the barrier; adding a wrapper-collapse rule does not; delayed and multiplicity-preserving SCC duplication both inherit the obstruction (§4.6).
C. Certified implementation and artifact engineering. For the guarded SafeStep fragment of KO7, the artifact certifies strong normalization, root-step confluence, a certified normalizer, decidable reachability to safe normal-form targets, partial context-closure strong normalization and confluence, a single-exponential contextual derivation-length theorem with an explicit lower family, an exact DM-component ordinal calibration at $\omega^{\omega}$, and a triple-lex upper-bound calibration below $\omega^{\omega}\!\cdot\!2$ (Theorems 5.95.11). This pins the guarded contextual complexity story at a single-exponential scale up to constant factors in the exponent and isolates the exact $\omega^{\omega}$ order type of the DM component as a standalone calibration result. For the full unguarded Step relation it formalizes two independent root-step termination proofs (nonlinear polynomial and KO7-specialized MPO), a Lean proof of full context-closed strong normalization by lifting the nonlinear witness through contexts, and a concrete derivation-length bound for contextual reductions. The support layer is executable: Meta/BarrierClass_Classifier.lean, Meta/BarrierWitness.lean, and Meta/SynthesisOracle.lean expose a classifier / witness / oracle pipeline for direct-orientation attempts, while Meta/TPDB_Export.lean and lakeexeverifyTpdbExport provide a checked bridge to the external TTT2/CeTA artifact. TTT2 independently reproduces the context-closed termination result; CeTA certifies the external proof.
Organization
[…]
2 Background
Standard abstract reduction and term rewriting notions apply [4, 26]. A term is in normal form if no rule applies. Strong normalization (SN) means that no infinite reduction exists. Confluence means that for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma: SN + local confluence $\Rightarrow$ confluence [22], yielding unique normal forms.
[…]
3.1 The step-duplicating schema
Definition 3.1 (Step-duplicating schema).
A step-duplicating schema consists of a term carrier $T$ with designated operations
$\mathrm{base}\in T,\qquad\mathrm{succ}:T\to T,\qquad\mathrm{wrap}:T\times T\to T,\qquad\mathrm{recur}:T^{3}\to T,$
and a step relation containing the duplicating rule
$\mathrm{recur}(b,s,\mathrm{succ}(n))\to\mathrm{wrap}(s,\mathrm{recur}(b,s,n)).$
Definition 3.2 (Direct compositional measures).
Two compositional axiom systems, both mapping terms to $\mathbb{N}$:
  1. 1.
    Tier 1 (additive). A fixed weight per constructor; the measure of a compound term is the constructor’s weight plus the sum of subterm measures. Hypothesis: wrap contributes at least 1.
  2. 2.
    Tier 2 (abstract, with wrapper-subterm axioms). Each constructor has an arbitrary combining function on subterm measures, subject to:
    $c_{\mathrm{wrap}}(x,y)>x\qquad\text{and}\qquad c_{\mathrm{wrap}}(x,y)>y.$
    No growth property is assumed for $\mathrm{succ}$. The Tier 2 impossibility additionally requires $\mathrm{succ}$-transparency at $\mathrm{base}$: $c_{\mathrm{succ}}(c_{\mathrm{base}})=c_{\mathrm{base}}$.
A measure orients a rule instance if the measured value strictly decreases across that instance.
Separately, Theorem 4.2 isolates an affine/linear constructor-local fragment:
$M(\mathrm{succ}(t))=\alpha_{s}+\beta_{s}\,M(t),\qquad M(\mathrm{wrap}(x,y))=\alpha_{w}+\beta_{w}\,M(x)+\gamma_{w}\,M(y),$
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n),$
with no cross terms. Its barrier uses an explicit unbounded-pump hypothesis rather than only the local axioms of Definition 3.2.
The schema and barrier theorems are formulated for a single self-recursive duplicating rule. Mutually recursive step-duplicating systems, where two or more functions duplicate each other’s arguments across distinct rules, are not analyzed here and may require separate barrier arguments.
Remark 3.3 (Classic duplicating TRS as schema instance).
The schema also captures standard parameterized duplicating rules from first-order rewriting. For example, the rule
$f(x,\,s(y))\;\to\;g(x,\,f(x,y))$
is an exact instance: take $\mathrm{base}=b$ as a dummy parameter, let $\mathrm{succ}=s$ act on the second argument, set $\mathrm{wrap}(a,r)=g(a,r)$, and define $\mathrm{recur}(b,a,n)=f(a,n)$. Then the schema step
$\mathrm{recur}(b,a,\mathrm{succ}\,n)\;\to\;\mathrm{wrap}(a,\mathrm{recur}(b,a,n))$
becomes precisely the displayed rule. The Tier 1 and Tier 2 barrier theorems therefore apply directly to this familiar duplicating shape. Thus the formalized barriers are not artifacts of KO7 but of the underlying self-recursive duplication pattern.
Meta/TextbookDupInstance.lean packages this case explicitly: the textbook rule $f(x,s(y))\to g(x,f(x,y))$ is turned into a separate schema-level case study, with direct additive, transparent-compositional, and affine barrier corollaries and specialized witness-extractor aliases. The reusable boundary layer therefore applies to a second named duplicating TRS, not only to KO7.
3.2 KO7
KO7 is a first-order ground term calculus with 7 constructors and 8 rules.
Kernel signature
$t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
Table 1: KO7’s 8 kernel rules. R_rec_succ (*) duplicates the step argument $s$. In full Step, R_eq_diff has no side condition and applies for all $a,b$.
Rule Head Arity Shape Dup?
R_merge_void_left merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R_merge_void_right merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R_merge_cancel merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R_int_delta integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R_eq_refl eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
In full Step, R_eq_refl and R_eq_diff overlap at $\mathrm{eqW}\,a\,a$. The SafeStep fragment adds per-rule guards: $\delta$-flag guards on the merge-void and rec-zero rules serve termination; the side conditions $\delta\text{-flag}(t)=0$ and $\kappa^{M}(t)=\varnothing$ on R_merge_cancel control the collapsing merge branch; and $\kappa^{M}(a)=\varnothing$ on R_eq_refl together with $a\neq b$ on R_eq_diff serve confluence.
Relations
Table 2: Certified status of the primary KO7 reduction relations.
Relation Scope SN Confluence Normalizer
Step Full, root Yes (mech.) No (witness) n/a
SafeStep Guarded, root Yes (mech.) Yes, root ARS Yes (mech.)
SafeStepCtx Partial ctx. Yes (mech.) Yes; Newman characterization retained n/a
The primary motivation for introducing SafeStep is confluence rather than root termination. The chosen guards are also coordinated with the computable termination certificate: deltaFlag and $\kappa^{M}$ appear both in the guarded rules and in the triple-lexicographic decrease argument. Later sections prove that the full root relation Step is terminating, but it already fails local confluence at the concrete term $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$:
$\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}\to\mathrm{void}\qquad\text{and}\qquad\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}\to\mathrm{integrate}(\mathrm{merge}\;\mathrm{void}\;\mathrm{void}).$
Both targets are full-Step root normal forms, so they do not join. The guarded relation SafeStep is introduced to block this overlap and recover unique normal forms for the root ARS.
The constructor app has no reduction rules. When rec_succ fires, the duplicated $s$ inside $\mathrm{app}(s,\mathrm{rec}\Delta(b,s,n))$ cannot feed back as a recursion counter: no rule has app at the root of its LHS. The duplicated payload does not itself become a new root recursive control position.
[…]
The impossibility and DP escape depend only on four constructors and two rules:
$\begin{array}[]{rcl}t&::=&\mathrm{void}\mid\delta\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\Delta\,t\,t\,t\\[4.0pt] \mathrm{rec}\Delta\,b\,s\,\mathrm{void}&\to&b\\ \mathrm{rec}\Delta\,b\,s\,(\delta\,n)&\to&\mathrm{app}\,s\,(\mathrm{rec}\Delta\,b\,s\,n)\end{array}$
[…]
4 The orientation barrier
Section preview.
This section isolates a concrete frontier between direct measure classes that provably fail and methods that succeed by importing structure outside those classes. Inside the formalized direct universe, the duplicating step is excluded by the schema, affine, restricted-quadratic, bounded cross-term, bounded multilinear, generalized bounded-polynomial, max-plus, tracked componentwise / lexicographic / balanced mixed-coordinate / weighted scalar-projection vector and pair barriers, together with the symbolic variable-condition (KBO) corollary. Crossing the frontier requires one of three moves exhibited in the artifact: projection-style transformation (dependency pairs), direct nonlinear escape outside the formalized classes (the global witness $W$ and the KO7-specialized MPO), or ablation of the offending structure itself. A more detailed recap appears at the end of the section as the Frontier summary.
The orientation barrier arises under four structural conditions, all satisfied by KO7:
  1. 1.
    No binding operators. First-order rewriting; no $\lambda$-abstraction, no free variables. Semantic methods (reducibility candidates, logical relations) rely on a typed or higher-order structure to supply the well-founded induction; KO7’s untyped first-order setting does not provide this structure.
  2. 2.
    Tree semantics (no sharing). All state is in the term tree; no external tape, graph-reduction sharing [4], or shared pointers. The step argument must be syntactically duplicated. This condition is part of the theorem scope: graph rewriting, shared environments, or explicit let-style sharing can change the duplication accounting and therefore fall outside the present barrier argument.
  3. 3.
    No axioms inside the object language. No equational theories or imported well-ordering principles are built into the rewrite language itself. (The rule $\mathrm{merge}(t,t)\to t$ is a directed reduction rule, not an algebraic identity.) The wrapper-collapse ablation in Section 4.6 should therefore be read only as a syntactic surrogate for object-level collapse, not as rewriting modulo an ambient equational theory.
  4. 4.
    Unrestricted step-duplicating recursor. A rule of the form $\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n))$ with $s$ unrestricted. Our ablation (§4.6) shows that removing duplication dissolves the barrier.
Simple typing by itself is not an escape from the barrier. The artifact includes a simply-typed first-order recursor fragment with separate result, step, and counter sorts, and proves that the additive barrier still survives there. It also proves that the affine barrier survives under an explicit typed step-pump hypothesis, namely, that the step sort still contains arbitrarily large closed terms. The point is structural: typing can block the old untyped pump patterns, but it does not dissolve the obstruction once a typed unbounded step family remains available. The Lean proofs are no_additive_orients_typed_recSucc and no_affine_orients_typed_recSucc_of_stepPump in Meta/TypedBarrierSurvival.lean. The same sort-indexed syntax is also repackaged in Meta/ManySortedBarrierSurvival.lean as a specialized many-sorted first-order presentation, yielding no_additive_orients_manySorted_recSucc and no_affine_orients_manySorted_recSucc_of_stepPump. This is a theorem-backed many-sorted extension of the KO7 recursor fragment, not a full mechanization of the general Aoto–Yamada translation theorem [1].
Under these conditions, the barrier stack has a clear internal structure rather than a flat catalog. Theorems 4.1 and 4.2 establish the baseline: additive, transparency-constrained compositional, and affine constructor-local measures already fail. Theorems 4.3 to 4.8 then widen the scalar side in stages, asking how much extra nonlinear growth or alternative aggregation can be added before the freeze-and-pump argument breaks; each theorem identifies the surviving obstruction. Theorems 4.10 to 4.14 form the vector / pair block, ending with a single scalar-projection meta-theorem that states the common matrix-side proof shape once. Theorem 4.15 and Corollary 4.16 close the direct symbolic side. This organization matters because the escape results later mirror it: successful methods work only by importing structure outside one of these blocks. Table 3 summarizes the resulting boundary.
This twelve-class stack should also be read against the modern interpretation-method taxonomy rather than as an ad hoc list. Tuple interpretations provide a unifying framework subsuming polynomial, matrix, arctic, and related interpretation families [30]. Most of our scalar and tracked vector / pair classes are direct syntactic fragments of that broader framework: the additive, affine, quadratic, multilinear, generalized bounded-polynomial, and max-plus barriers are scalar or one-coordinate fragments, while the fixed-dimension, lexicographic, balanced mixed-coordinate, and weighted scalar-projection theorems are tracked-tuple fragments with restricted comparison interfaces. The extra pump and dominance hypotheses in our statements are exactly the explicit growth conditions under which those direct fragments can be tested against the step-duplicating schema. The escape side then has a clean interpretation: DP projection and the full-step witnesses do not refute tuple-interpretation methods wholesale; they succeed by leaving this direct boundary through projection or a stronger global proof shape.
Table 3: Orientation boundary map for KO7’s step-duplicating recursor.
Bucket KO7 status
Provably blocked classes Twelve direct families are ruled out by theorem: additive and transparent-compositional measures; affine and nonlinear scalar extensions (restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus); and tracked vector/pair extensions (fixed-dimension componentwise, dimension-$2$ lexicographic, dimension-$2$ balanced mixed-coordinate, and weighted scalar-projection componentwise). The additive and transparent-compositional classes are excluded unconditionally; the remaining ten are excluded under their stated pump or growth hypotheses (Theorems 4.1, 4.2, 4.3, 4.4, 4.5, 4.6, 4.8, 4.10, 4.11, 4.12, 4.13).
[…]
Transformed / projection escape Dependency-pair projection succeeds by discarding duplicated payload and thereby violating wrapper-subterm sensitivity (Theorem 4.22).
Direct out-of-class global orienters The full root relation is oriented internally by the nonlinear polynomial witness and the KO7-specialized MPO, both outside the formalized direct barrier classes (Propositions 4.27, 4.28).
Table 4: Status of the main barrier side conditions.
Class block Status Explanation
Additive; transparent-compositional Unconditional No extra growth or dominance side condition is needed.
Affine; restricted quadratic; max-plus; tracked-primary vector / pair; weighted scalar-projection; balanced mixed-coordinate Standard pump hypotheses The required growth hypotheses are easy to discharge from a successor pump or wrap/base pump and are internalized as strengthened subclasses where appropriate in Meta/PumpedBarrierClasses.lean.
Bounded cross-term quadratic; bounded multilinear; generalized bounded polynomial Load-bearing dominance hypotheses These classes carry genuine frozen-coefficient or base-dominance constraints that delimit the theorem scope rather than following automatically from the schema alone.
This summary is only a status map for the stated hypotheses, not an extra theorem. Its purpose is to separate routine pump side conditions from the genuinely restrictive dominance assumptions that remain part of the mathematical boundary.
4.1 Barrier at the schema
Theorem 4.1 (Barrier at the schema).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema (Definition 3.1). Then:
  1. 1.
    No Tier 1 measure with positive wrap-weight strictly orients the duplicating step uniformly.
  2. 2.
    No Tier 2 measure with the wrap-subterm axioms and successor transparency at $\mathrm{base}$ strictly orients the duplicating step uniformly.
  3. 3.
    Neither class globally orients any relation containing the duplicating step.
The Lean proof is in Meta/StepDuplicatingSchema.lean. Computable certificate extractors in Meta/BarrierWitness.lean (additive_witness, compositional_witness, affine_witness) return a concrete violating triple $(b,s,n)$ with a bundled proof of orientation failure: directly from the measure data in the additive and transparent-compositional cases, and from the measure together with a supplied pump term above the required bound in the affine case. The companion module Meta/BarrierWitness_Budgets.lean packages the corresponding construction-budget statements: for the current extractor layer, the generated step payload is always the base term, a successor chain succIterk, or a wrapper chain wrapIterk, with an explicit budget $k$ read off from the coefficient data.
Argument. For Tier 1, set $b=n=\mathrm{base}$ and $s=\mathrm{wrapIter}(w_{\mathrm{succ}})$ (iterated wrapper). By additivity, $M(s)\geq w_{\mathrm{succ}}$. The rewrite right-hand side $\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base}))$ contributes $w_{\mathrm{wrap}}+M(s)$ beyond the shared recursive subterm, while the left-hand side $\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base}))$ contributes only $w_{\mathrm{succ}}$. Since $w_{\mathrm{wrap}}\geq 1$ and $M(s)\geq w_{\mathrm{succ}}$, the right-hand side is at least as large, contradicting strict orientation. For Tier 2, set $b=s=n=\mathrm{base}$. By transparency, $c_{\mathrm{succ}}(c_{\mathrm{base}})=c_{\mathrm{base}}$, so $M(\mathrm{recur}(\mathrm{base},\mathrm{base},\mathrm{succ}(\mathrm{base})))=c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}})$. The right-hand side $c_{\mathrm{wrap}}(c_{\mathrm{base}},c_{\mathrm{recur}}(c_{\mathrm{base}},c_{\mathrm{base}},c_{\mathrm{base}}))$ exceeds the left-hand side by the wrapper-subterm axiom $c_{\mathrm{wrap}}(x,y)>y$. We retain the symmetric pair of wrapper-subterm axioms in the theorem interface for standardity and for downstream uniformity across the compositional API, although only the second inequality drives the contradiction. $\square$
4.2 Affine / linear constructor-local measures
Theorem 4.2 (Affine / linear barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy
$M(\mathrm{succ}(t))=\alpha_{s}+\beta_{s}\,M(t),$
$M(\mathrm{wrap}(x,y))=\alpha_{w}+\beta_{w}\,M(x)+\gamma_{w}\,M(y),$
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n),$
with $\beta_{w},\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, i.e.
$\forall k\in\mathbb{N}\;\exists t\in T.\;M(t)\geq k,$
then $M$ does not strictly orient the duplicating step uniformly. In particular, the same conclusion follows if either
  1. 1.
    $\alpha_{s}\geq 1$ and $\beta_{s}\geq 1$ (positive successor pump), or
  2. 2.
    $\alpha_{w}+\gamma_{w}\,M(\mathrm{base})\geq 1$ (positive wrap/base pump).
Formalized in Meta/StepDuplicatingSchema.lean.
Both alternatives entail unbounded attainable values: the successor clause pumps directly, while the wrap/base clause pumps through the left-nested recurrence
$x_{k+1}=\alpha_{w}+\beta_{w}x_{k}+\gamma_{w}M(\mathrm{base}),$
that is, through wrapIter with the left argument iterated and the right argument frozen at $\mathrm{base}$. With $\beta_{w}\geq 1$ and positive constant term $\alpha_{w}+\gamma_{w}M(\mathrm{base})\geq 1$, this yields unbounded growth.
Fix $b=n=\mathrm{base}$ and pump $s$. Write
$C_{\mathrm{src}}:=\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\delta_{r}\,M(\mathrm{succ}(\mathrm{base})),\qquad C_{\mathrm{tgt}}:=\alpha_{w}+\gamma_{w}\!\bigl(\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\delta_{r}\,M(\mathrm{base})\bigr),$
so the two sides of the duplicating rule become
$M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))=C_{\mathrm{src}}+\gamma_{r}\,M(s),$
$M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))=C_{\mathrm{tgt}}+(\beta_{w}+\gamma_{w}\gamma_{r})\,M(s).$
Hence the pumped-variable coefficient on the target exceeds the source coefficient by
$(\beta_{w}+\gamma_{w}\gamma_{r})-\gamma_{r}=\beta_{w}+(\gamma_{w}-1)\gamma_{r}\geq 1.$
Once $M(s)$ is larger than the fixed constant gap between $C_{\mathrm{src}}$ and $C_{\mathrm{tgt}}$, the target exceeds the source, contradicting strict orientation. The later scalar barrier theorems reuse this same coefficient comparison after freezing their additional nonlinear terms at the base point. $\square$
Sharpness note. The contradiction bound used in Theorem 4.2 is not merely a proof artifact. In Meta/AffineThresholdSharpness.lean we define a canonical affine family on $\mathbb{N}$ with
$M(\mathrm{base})=0,\qquad M(\mathrm{succ}(t))=1,\qquad M(\mathrm{wrap}(x,y))=x+y,\qquad M(\mathrm{recur}(b,s,n))=\delta_{r}\,n,$
for which the generic affine pump bound is exactly $\delta_{r}$, and the distinguished root-step instance satisfies
$M(\mathrm{wrap}(k,\mathrm{recur}(\mathrm{base},k,\mathrm{base})))<M(\mathrm{recur}(\mathrm{base},k,\mathrm{succ}(\mathrm{base})))\quad\Longleftrightarrow\quad k<\delta_{r}.$
Thus the generic affine contradiction bound is exact for this family. Intuitively, the canonical family makes the duplicated wrapper contribution grow exactly like the pumped payload $k$, while the source still only sees the fixed counter coefficient $\delta_{r}$; the contradiction therefore flips precisely at the cutoff $k=\delta_{r}$. Formalized as affineThresholdMeasure_exact_cutoff.
Scope note. Theorems 4.1 and 4.2 cover the four-role schema and three axiomatized measure classes. They do not cover nonlinear polynomial interpretations with unrestricted cross-term coupling, generic DP frameworks, or duplicating rule shapes not reduced to a single wrapper. The following extensions address restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, max-plus, and tracked or projected vector / pair classes within the same schema; the full scope discussion is in §4.8.
Theorem 4.3 (Restricted quadratic counter barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy the same affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$ as in Theorem 4.2, and let the recursor take the restricted degree-$2$ form
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n)+\varepsilon_{r}\,M(n)^{2},$
with $\beta_{w},\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump, exactly as in Theorem 4.2. See Meta/QuadraticBarrier.lean (no_quadratic_counter_orients_dup_step_of_unbounded and the successor- and wrap-pump corollaries).
The argument follows the same freeze-and-pump pattern. Fix $b=n=\mathrm{base}$ and pump the step argument $s$. The rewrite source is
$M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))={}$ $\underbrace{\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\delta_{r}\,M(\mathrm{succ}(\mathrm{base}))}_{\text{constant in }M(s)}$
$+\underbrace{\varepsilon_{r}\,M(\mathrm{succ}(\mathrm{base}))^{2}}_{\text{constant in }M(s)}+\gamma_{r}\,M(s).$
The rewrite target is
$M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))={}$ $\alpha_{w}+\beta_{w}\,M(s)$
$+\gamma_{w}(\alpha_{r}+\beta_{r}\,M(\mathrm{base})+\gamma_{r}\,M(s)$
$\qquad+\delta_{r}\,M(\mathrm{base})+\varepsilon_{r}\,M(\mathrm{base})^{2}).$
[…]
Theorem 4.4 (Bounded cross-term quadratic barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy the same affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$ as in Theorem 4.2, and let the recursor take the degree-$2$ form
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}\,M(b)+\gamma_{r}\,M(s)+\delta_{r}\,M(n)+\varepsilon_{r}\,M(n)^{2}+\zeta_{r}\,M(s)M(n).$
Assume $M$ has unbounded attainable values, and that the cross-term coupling satisfies the bounded-regime condition
$\gamma_{r}+\zeta_{r}\,M(\mathrm{succ}(\mathrm{base}))+1\;\leq\;\beta_{w}+\gamma_{w}\bigl(\gamma_{r}+\zeta_{r}\,M(\mathrm{base})\bigr),$
formalized in Meta/QuadraticCrossTermBarrier.lean as CrossTermBoundedAtBase. Then $M$ does not strictly orient the duplicating step uniformly. In particular, the same conclusion follows from either a positive successor pump or a positive wrap/base pump together with this bounded-coupling condition. Proved in Meta/QuadraticCrossTermBarrier.lean (no_cross_quadratic_orients_dup_step_of_unbounded and KO7 specializations).
Non-vacuous example. The bounded-coupling condition is not an ad hoc exclusion of the known nonlinear witness. For example, if $M(\mathrm{base})=0$, $M(\mathrm{succ}(t))=1+M(t)$, $M(\mathrm{wrap}(x,y))=1+2M(x)+M(y)$, and
$M(\mathrm{recur}(b,s,n))=M(b)+M(n)+M(n)^{2}+M(s)M(n),$
then $\zeta_{r}=1$ gives genuine step-counter coupling, yet CrossTermBoundedAtBase still holds because the wrapper gain on the duplicated step argument dominates the one-step cross-term increment at the base point.
By contrast, the successful nonlinear witness $W$ lies outside this bounded-coupling fragment: its escape uses broader cross-variable coupling together with successor non-transparency, so the present base-point dominance condition is not the mechanism that certifies it.
Again fix $b=n=\mathrm{base}$ and pump the step argument $s$. The source coefficient in the pump variable becomes $\gamma_{r}+\zeta_{r}\,M(\mathrm{succ}(\mathrm{base}))$, while the target coefficient becomes $\beta_{w}+\gamma_{w}(\gamma_{r}+\zeta_{r}\,M(\mathrm{base}))$. The bounded-coupling hypothesis says the target coefficient still dominates the source coefficient by at least one. Once the pumped step argument exceeds the frozen constant part, the target overtakes the source and strict decrease is impossible. So the barrier reaches one step beyond the pure counter-square class: some step-counter coupling is still excluded, provided the wrapper gain remains dominant at the base point. $\square$
In the rest of this subsection, the bounded cross-term, bounded multilinear, and generalized bounded-polynomial barriers should be read as exclusion theorems for the frozen-base-dominated subclass, not for the entire syntactic family. The dominance predicate is the load-bearing hypothesis in each statement and is summarized again in Table 4.
For the multilinear theorem, if $B_{0}=M(\mathrm{base})$ and $S_{0}=M(\mathrm{succ}(\mathrm{base}))$, the frozen dominance condition is
$\mathrm{sourceCoeff}+1\leq\mathrm{targetCoeff},$
where $\mathrm{sourceCoeff}$ is the total coefficient of the pumped variable $M(s)$ in the frozen source polynomial $M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))$, and $\mathrm{targetCoeff}$ is the coefficient of $M(s)$ in the frozen target polynomial $M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))$. In Lean this is the explicit predicate MultilinearDominatedAtBase; once $b$ and $n$ are frozen, every multilinear monomial becomes affine in the pumped variable.
Theorem 4.5 (Bounded multilinear barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy the same affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$ as in Theorem 4.2, and let the recursor take the form
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\beta_{r}M(b)+\gamma_{r}M(s)+\delta_{r}M(n)+\sum_{m\in\mathcal{T}}m(M(b),M(s),M(n)),$
where $\mathcal{T}$ is a finite table of multilinear monomials, each variable appearing at most once per monomial. Assume $M$ has unbounded attainable values and that the frozen dominance condition $\mathrm{sourceCoeff}+1\leq\mathrm{targetCoeff}$ holds. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a positive successor pump or a positive wrap/base pump under the same dominance condition. The Lean development is Meta/MultilinearBarrier.lean (no_multilinear_orients_dup_step_of_unbounded, successor-pump, wrap-pump, and KO7 corollaries).
Proof idea. Freeze $b=\mathrm{base}$ and pump the step argument $s$. Each multilinear monomial becomes affine in $M(s)$ once the other arguments are frozen, so the entire finite monomial table collapses to a constant part plus a frozen coefficient times $M(s)$ on both the source and target sides. The explicit dominance condition says the wrapper-side coefficient still exceeds the frozen source coefficient by at least one. Once $M(s)$ exceeds the frozen constant part, the target overtakes the source and strict decrease becomes impossible. A large finite multilinear family therefore still fails as long as its pumped-variable contribution remains frozen and dominated in this sense. $\square$
Theorem 4.6 (Generalized degree-bounded polynomial barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy constructor-local affine equations for $\mathrm{succ}$ and $\mathrm{wrap}$, and let the recursor be given by a finite polynomial table in the three tracked scalar arguments $M(b)$, $M(s)$, and $M(n)$, with arbitrary natural exponents allowed in each monomial. Assume $M$ has unbounded attainable values and that, after freezing $b=\mathrm{base}$ and comparing the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$, the frozen target-side polynomial eventually dominates the frozen source-side polynomial as the pumped step value grows. Then $M$ does not strictly orient the duplicating step uniformly. The same conclusion follows from either a certified successor pump or a certified wrap pump under the same frozen dominance condition. Equivalently, any successful orienter inside this bounded direct-polynomial family must violate the frozen base-dominance condition. Formalized as no_polynomial_orients_dup_step_of_unbounded, no_global_step_orientation_polynomial_of_succ_pump, and polynomial_escape_requires_failure_of_base_dominance in Meta/PolynomialBarrierGeneral.lean.
Explicitly, the dominance condition is
$\exists K\,\forall X\geq K,\;P_{\mathrm{src}}(X)\leq P_{\mathrm{tgt}}(X),$
where $P_{\mathrm{src}}(X)$ is the one-variable polynomial obtained from $M(\mathrm{recur}(\mathrm{base},s,\mathrm{succ}(\mathrm{base})))$ by freezing $M(b)=M(\mathrm{base})$ and replacing $M(s)$ by $X$, and $P_{\mathrm{tgt}}(X)$ is the corresponding frozen target polynomial for $M(\mathrm{wrap}(s,\mathrm{recur}(\mathrm{base},s,\mathrm{base})))$. In Lean this predicate is EventuallyDominatedAtBase.
Freeze the non-pumped arguments at the base point and compare the source counter $\mathrm{succ}(\mathrm{base})$ with the target counter $\mathrm{base}$. Each monomial then becomes a one-variable polynomial in the pumped step value, possibly with repeated occurrences of that variable. The explicit dominance hypothesis says that from some cutoff onward the frozen target-side polynomial overtakes the frozen source-side polynomial. Choose the pumped step value beyond that cutoff. The target is then at least the source, so strict decrease fails. The argument extends beyond the multilinear regime because repeated variables are allowed; however, the resulting one-variable growth must remain dominated at the frozen base point. $\square$
Accordingly, Theorem 4.6 should be read as the artifact’s parametric direct bounded-polynomial barrier for the frozen-base-dominated subclass: any direct bounded-degree polynomial algebra with the standard succ/wrap interface and the same frozen base-dominance condition falls under it, regardless of the specific finite recursor monomial table.
Corollary 4.7 (WPO-style bounded polynomial branch barrier).
Let $\succ$ be any direct order on a step-duplicating schema such that every strict comparison $x\succ y$ forces strict descent in a bounded degree-bounded polynomial algebra of Theorem 4.6. Under the same unbounded or certified pump hypotheses and the same frozen base-dominance condition, $\succ$ cannot orient the duplicating step uniformly. In particular, the KO7 root relation admits no such direct WPO-style polynomial-branch orienter. See Meta/WPO_PolynomialBarrier.lean (no_wpoPolynomialDirect_orients_dup_step_of_unbounded, no_global_step_orientation_wpoPolynomialDirect_of_succ_pump, wpoPolynomialDirect_escape_requires_failure_of_base_dominance).
No new polynomial argument is needed. If strict comparison in the candidate order already implies strict descent of the bounded polynomial algebra, then any successful orientation would induce a successful instance of Theorem 4.6. The polynomial barrier therefore transfers immediately through that soundness implication. The scope is intentionally narrow: the corollary captures the direct bounded polynomial-algebra branch used inside WPO-style tool implementations, not generic recursive path descent or the full WPO metatheory. $\square$
Theorem 4.8 (Schema-level max-plus barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $M:T\to\mathbb{N}$ satisfy constructor-local max-plus equations of the form
$M(\mathrm{succ}(t))=\alpha_{s}+M(t),$
$M(\mathrm{wrap}(x,y))=\alpha_{w}+\max(\beta_{w}+M(x),\gamma_{w}+M(y)),$
$M(\mathrm{recur}(b,s,n))=\alpha_{r}+\max(\beta_{r}+M(b),\max(\gamma_{r}+M(s),\delta_{r}+M(n))).$
Assume moreover that $\gamma_{w}\geq 1$. If $M$ has unbounded attainable values, then $M$ does not strictly orient the duplicating step uniformly. The same conclusion also follows from either a positive successor drift or a positive left-wrapper drift, but these are separate pump-side hypotheses used only to generate unbounded inputs. The barrier proof itself uses the positive offset on the wrapped recursive branch. The Lean proof is no_max_orients_dup_step_of_unbounded in Meta/MaxBarrier.lean, with successor-pump, wrap-pump, and KO7 corollaries in the same module.
Sketch of the max argument. Pump the step argument until the step branch $\gamma_{r}+M(s)$ dominates both the base branch and the counter branch inside the recursor. At that point the source becomes $\alpha_{r}+\gamma_{r}+M(s)$, while the target wraps the same visible branch again through a max-plus wrapper. Because the wrapper keeps a visible copy of the pumped branch, and because the right wrapper branch carries the positive offset $\gamma_{w}\geq 1$ (absorbing the at-most-one-step successor offset $\alpha_{s}$), the target is at least the source plus one extra unit on that branch, so strict decrease fails. The obstruction therefore persists even when aggregation is by $\max$ rather than by addition. $\square$
Corollary 4.9 (Arctic primary-projection barrier).
Any direct arctic-style interpretation with a distinguished finite primary projection satisfying the max-plus interface of Theorem 4.8 is blocked by the same argument. In particular, the KO7 root relation admits no such arctic primary-projection orienter under the corresponding unbounded or certified pump hypotheses. Proved as no_arctic_primary_orients_dup_step_of_unbounded in Meta/ArcticBarrier.lean; KO7 corollary: no_global_step_orientation_arctic_primary_of_unbounded.
Forget the arctic carrier and read only the distinguished finite primary coordinate. By hypothesis that projection already satisfies the max-plus interface, so Theorem 4.8 applies directly. The corollary is a tool-facing instantiation of the scalar max barrier rather than a separate max-algebra proof. $\square$
The same projection pattern extends further to a broader tropical primary-projection corollary, formalized in Meta/TropicalBarrier.lean. We treat that module as an artifact-level continuation of the max-plus / arctic line rather than as a thirteenth direct family in the main theorem count.
Theorem 4.10 (Fixed-dimension componentwise affine barrier).
Fix any finite dimension $d\geq 1$. Let $M:T\to\mathbb{N}^{d}$ be a vector-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume one tracked coordinate of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that tracked coordinate. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, the same conclusion follows if the tracked coordinate admits either a positive successor pump or a positive wrap/base pump. Formalized in Meta/MatrixBarrierD.lean (no_matrixD_orients_dup_step_of_componentwise_pump); the earlier dimension-$2$ development is retained in Meta/MatrixBarrier2.lean as the concrete low-dimensional worked instance.
Strict componentwise orientation implies strict decrease in every coordinate, hence in the tracked one. Project the vector-valued measure to that tracked coordinate and apply Theorem 4.2; the contradiction is immediate. The result covers arbitrary fixed dimension under strict componentwise order, but still only through one tracked affine coordinate, not through arbitrary mixed matrix interactions. $\square$
Theorem 4.11 (Dimension-$2$ lexicographic affine barrier).
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered lexicographically by the first component. Assume the first component of $M$ satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that first component. Then the duplicating step cannot be strictly oriented lexicographically by $M$. In particular, the same conclusion follows if the first component admits either a positive successor pump or a positive wrap/base pump. See Meta/MatrixBarrierLex.lean (no_matrix2_lex_orients_dup_step_of_unbounded_primary and KO7 specializations).
Lexicographic decrease implies that the first component of the target is at most the first component of the source. Pump the tracked first component one unit beyond the scalar affine bound. The first component of the duplicating-step target then exceeds the first component of the source, so lexicographic decrease is impossible regardless of what the second component does. $\square$
Theorem 4.12 (Balanced mixed-coordinate dimension-$2$ barrier).
Let $M:T\to\mathbb{N}^{2}$ be a pair-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume each constructor acts by a full mixed $2\times 2$ linear map on the two coordinates, and that the two column sums of each such map agree. Then the aggregate coordinate sum $M_{1}+M_{2}$ is an affine scalar measure. If that aggregate sum has an unbounded pump and the aggregate wrapper coefficients are positive, then the duplicating step cannot be strictly oriented componentwise by $M$. The Lean proof is no_matrixMix2_orients_dup_step_of_sum_pump in Meta/MatrixBarrierMix2.lean.
Aggregate-sum reduction. Balanced column sums make the aggregate quantity $M_{1}(t)+M_{2}(t)$ behave like a scalar affine interpretation, even though each coordinate may mix both inputs through off-diagonal coefficients. Strict componentwise decrease forces strict decrease of this aggregate sum, so Theorem 4.2 applies to the aggregate projection. Unlike the earlier matrix barriers, the proof here is not a single-coordinate projection theorem; it still stops short of arbitrary mixed matrix interpretations however. $\square$
Theorem 4.13 (Weighted functional componentwise affine barrier).
Fix any finite dimension $d\geq 1$. Let $M:T\to\mathbb{N}^{d}$ be a vector-valued constructor-local measure on a step-duplicating schema, ordered by strict componentwise decrease. Assume there is a fixed nonzero natural weight vector $\lambda\in\mathbb{N}^{d}$ such that the scalar projection
$\lambda\cdot M(t):=\sum_{i}\lambda_{i}M_{i}(t)$
satisfies the affine equations of Theorem 4.2 with positive wrapper coefficients and an unbounded pump in that weighted projection. Then the duplicating step cannot be strictly oriented componentwise by $M$. In particular, singleton weights recover the tracked-coordinate argument, while the all-ones vector recovers aggregate-sum arguments. Proved in Meta/MatrixBarrierFunctional.lean (no_matrixFunctional_orients_dup_step_of_componentwise_pump), with successor-pump and wrap-pump corollaries.
Strict componentwise decrease forces strict decrease of every nonzero natural-weighted scalar projection. Apply this to $\lambda\cdot M$. By hypothesis, that projection already satisfies the scalar affine interface of Theorem 4.2. The contradiction therefore reduces to the scalar affine barrier after one projection step, unifying the tracked-coordinate and balanced aggregate-sum proof shapes under one principle: once a fixed scalar projection carries the affine interface, strict componentwise orientation is impossible. $\square$
Theorem 4.14 (Scalar-projection meta-barrier).
Let $(T,\mathrm{base},\mathrm{succ},\mathrm{wrap},\mathrm{recur})$ be a step-duplicating schema. Let $\mu:T\to A$ be any direct orienter into a codomain $A$, let $R$ be the chosen strict order on $A$, and let $\pi:A\to\mathbb{N}$ be a tracked scalar projection. If every strict $R$-decrease forces strict decrease of $\pi$, and if strict decrease of $\pi\circ\mu$ on the duplicating step is already ruled out by a scalar barrier theorem, then $R$ itself cannot orient the duplicating step uniformly. Formalized as no_orients_dup_step_of_scalar_projection in Meta/ScalarProjectionBarrier.lean, with concrete rederivations for tracked componentwise pairs, weighted functional matrices, and balanced mixed-coordinate pairs.
Suppose for contradiction that $R$ oriented the duplicating step uniformly. Every such $R$-decrease forces strict decrease of $\pi$, so the projected scalar family $\pi\circ\mu$ would also orient the duplicating step uniformly, contradicting the already available scalar barrier theorem. The result captures in one statement the common proof shape shared by the tracked-component, weighted-functional, and balanced mixed-coordinate matrix barriers. $\square$
Theorem 4.15 (Symbolic variable-condition barrier).
Let $\succ$ be any strict symbolic comparator on the abstract rule schema
$\mathrm{recur}(b,s,\mathrm{succ}(n))\to\mathrm{wrap}(s,\mathrm{recur}(b,s,n))$
such that $\succ$ satisfies the standard variable condition: if $x\succ y$, then every variable occurs in $y$ at most as often as in $x$. Then $\succ$ cannot orient the duplicating step from left to right. See Meta/SymbolicComparatorBarrier.lean (not_orients_dup_rule).
The variable $s$ occurs once on the source side and twice on the target side. Any order satisfying the variable condition must reject this orientation immediately, widening the negative story beyond numeric measures: direct KBO-style symbolic comparators already fail at the rule schema before any recursive path descent or modular transformation is introduced. $\square$
Corollary 4.16 (Explicit KO7 KBO impossibility).
No KBO-style order on concrete KO7 traces, once restricted along the instantiated duplicating schema and assumed to satisfy the standard variable condition there, can orient the KO7 rec_succ rule. Consequently, no such KBO-style order can orient the KO7 root relation by decreasing every root step. The formalized concrete bridge theorem is no_kbo_orients_ko7_rec_succ_trace in Meta/KBO_Impossible.lean; the underlying abstract schema-level obstruction remains no_kbo_orients_ko7_rec_succ.
Instantiate Theorem 4.15 with the KO7 duplicating rule schema. No new proof mechanism is needed; the corollary records the KBO-facing consequence under a theorem name reviewers can cite directly. $\square$
The corollary is extracted from the symbolic variable-condition barrier and does not formalize generic Knuth–Bendix order metatheory over arbitrary signatures.
Architectural note on Meta/KBO_Impossible.lean. The cited Lean module is intentionally a thin paper-facing renaming layer. The substantive obstruction lives in Meta/SymbolicComparatorBarrier.lean (not_orients_dup_rule) and is reused unchanged. The KBO_Impossible module contributes only a type alias KBOStyleOrder:=VariableConditionOrder, two one-line forwarding theorems under KBO-facing names, and the new concrete-trace bridge no_kbo_orients_ko7_rec_succ_trace that lifts the schema-level statement to literal Trace terms via the instantiate map. Readers searching the artifact for the actual proof of the variable-condition obstruction should look at Meta/SymbolicComparatorBarrier.lean, not at this file.
Note on the vector / pair extensions. The fixed-dimension componentwise barrier and the lexicographic dimension-$2$ barrier (Theorems 4.10 and 4.11) push the barrier beyond scalar measures, but the obstruction is still the affine tracked-component argument of Theorem 4.2. The balanced mixed-coordinate theorem (Theorem 4.12) allows off-diagonal coordinate mixing, but only in a balanced regime where the aggregate sum of the two coordinates still yields a scalar affine projection. The weighted functional theorem (Theorem 4.13) abstracts both moves: any nonzero fixed scalar projection with an affine interface inherits the same barrier. The scalar-projection meta-theorem (Theorem 4.14) states that proof shape once and for all. The matrix story therefore has four layers: tracked-coordinate, balanced mixed-coordinate, weighted scalar projection, and projection-level unification. Arbitrary mixed matrix interpretations remain outside the present barrier stack.
Corollary 4.17 (Unconditional pumped-subclass forms).
The growth-side hypotheses in the affine, restricted-quadratic, bounded cross-term, generalized bounded-polynomial, bounded multilinear, max-plus, and current projection-based matrix barriers can be internalized into strengthened subclasses. The resulting affine, restricted-quadratic, bounded cross-term, generalized bounded-polynomial, bounded multilinear, max-plus, tracked-primary pair, weighted functional, and balanced mixed-coordinate families therefore admit unconditional exclusions inside their subclass definitions. Machine-checked in Meta/PumpedBarrierClasses.lean; the exact identifiers are listed in Table LABEL:tab:claim-code-barrier.
The bridge from the conditional barrier theorems to these unconditional pumped-subclass corollaries is factored out in Meta/StandardPumpLemmas.lean. That module proves reusable successor-pump and wrap-pump lemmas for the affine, restricted-quadratic, and current tracked-primary dimension-$2$ pair families, and packages the corresponding hypotheses into the strengthened subclasses used by Corollary 4.17. This keeps the theorem statements honest while making the standard growth side conditions available as reusable certified infrastructure.
Corollary 4.18 (Context-closed barrier survival).
Every direct class blocked at the duplicating root step is also blocked for the full context-closed relation StepCtxFull, because every root contraction is a contextual contraction under the identity context. In particular, the additive, transparent-compositional, affine-with-pump, restricted-quadratic, bounded cross-term, bounded multilinear, max-plus, tracked-primary pair, weighted scalar-projection, and balanced mixed-coordinate families all admit explicit context-closed corollaries. Proved in Meta/ContextClosedBarrier.lean as stepCtxFull_orientation_implies_root together with the corresponding class-specific KO7 corollaries.
This corollary adds no new obstruction mechanism. It simply states the root-step barrier stack at the same context-closed semantic level used by standard TRS tools and by the external TPDB / CeTA semantics in Section 6.
4.3 KO7 instantiation
Theorem 4.19 (Compositional impossibility for rec_succ).
KO7 instantiates the schema with $\mathrm{base}=\mathrm{void}$, $\mathrm{succ}=\delta$, $\mathrm{wrap}=\mathrm{app}$, $\mathrm{recur}=\mathrm{rec}\Delta$. Both impossibility results hold:
  1. 1.
    Tier 1. For every additive compositional measure $M$ with app-weight $\geq 1$:
    $\neg\bigl(\forall b\,s\,n.\;M(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))<M(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))\bigr).$
  2. 2.
    Tier 2. For every abstract compositional measure $CM$ with app-subterm axioms and $\delta$-transparency at void:
    $\neg\bigl(\forall b\,s\,n.\;CM(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))<CM(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))\bigr).$
Direct instantiation of Theorem 4.1 with $\mathrm{base}=\mathrm{void}$, $\mathrm{succ}=\delta$, $\mathrm{wrap}=\mathrm{app}$, $\mathrm{recur}=\mathrm{rec}\Delta$; see Meta/CompositionalMeasure_Impossibility.lean. $\square$
The same module also gives the KO7 affine instantiation. It packages a 7-constructor affine evaluation as AffineCompositionalMeasure, lifts it to the schema via toSchemaMeasure, and yields the corresponding KO7 affine barrier with pump corollaries. The KO7 restricted-quadratic and bounded cross-term extensions then appear in Meta/QuadraticBarrier.lean and Meta/QuadraticCrossTermBarrier.lean, each with unbounded, successor-pump, and wrap-pump forms. The multilinear, generalized bounded-polynomial, max-plus, and arctic primary-projection continuations then live in Meta/MultilinearBarrier.lean, Meta/PolynomialBarrierGeneral.lean, Meta/MaxBarrier.lean, and Meta/ArcticBarrier.lean. Meta/PumpedBarrierClasses.lean repackages the affine, restricted-quadratic, bounded cross-term, generalized bounded-polynomial, tracked-primary pair, bounded multilinear, max-plus, weighted functional, and balanced mixed-coordinate KO7 exclusions as unconditional statements for the corresponding strengthened subclasses.
4.4 Essentiality of the transparency hypothesis
Theorem 4.20 (Transparency is essential for Tier 2).
[…]
$CM.\mathrm{eval}(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))\;<\;CM.\mathrm{eval}(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))$
for all $b,s,n$. One concrete witness is
$c_{\mathrm{void}}=1,\qquad c_{\delta}(x)=x+1,\qquad c_{\mathrm{app}}(x,y)=x+y+1,\qquad c_{\mathrm{rec}\Delta}(a,b,c)=a+b(c+1)^{2}+c.$
Once transparency is dropped, the Tier 2 class contains measures that orient the step (see Meta/RecCore.lean).
[…]
Remark 4.21 (Every successful method imports assumptions outside the barrier classes).
The essentiality theorem and the DP escape (Theorem 4.22) instantiate a general pattern: every method that orients the duplicating step does so by violating at least one axiom of the formalized compositional classes. DP projection violates the wrapper-subterm axioms (setting $\pi(\mathrm{app}\;x\;y)=0$). LPO imports the Subterm Property, a structural axiom not derivable from the rewrite rules. The nonlinear witness above violates successor transparency ($c_{\delta}(1)=2\neq 1$) and uses a nonlinear recursor combiner not representable in Tier 1 or the affine class of Theorem 4.2. An arbitrary nonlinear polynomial interpretation over $\mathbb{Z}^{+}$ can likewise orient the full system. For example, assigning $W(\mathrm{rec}\Delta(b,s,n))=(W(n)+1)\cdot(W(s)+W(b)+1)$ with $W(\delta\,t)=W(t)+1$ yields $W(\text{LHS})-W(\text{RHS})=W(b)\geq 1$ on rec_succ. This requires both nonlinearity (cross-term coupling between counter and payload) and non-transparency ($W(\delta\,t)=W(t)+1\neq W(t)$), which place it outside every formalized barrier class. This full-system nonlinear interpretation is machine-checked: W_orients_step in Meta/PolyInterpretation_FullStep.lean proves $W(b)<W(a)$ for every root step $a\to b$, with axiom-violation witnesses W_violates_transparency, W_not_additive, and W_not_affine.
The barrier does not assert that no termination proof exists. It characterizes precisely which structural assumptions any successful proof must import. A method that works on the duplicating step confirms the barrier’s precision by demonstrating which axioms it violates.
Table 5: Escape methods and the structural assumptions they import.
Escape method Broken / imported structural assumption Resulting escape shape
DP projection Wrapper sensitivity on the duplicated payload is discarded by projecting to the counter position. Modular transformed-relation escape.
Nonlinear $W$ witness Successor transparency and direct affine / additive structure are violated by cross-variable coupling. Global nonlinear full-step orientation.
KO7-specialized MPO The direct Nat-valued measure interface is left in favor of fixed-signature path-order structure, recursive subterm comparison, and a specialized same-head clause. Fixed-signature path-order escape.
Mechanized violation witnesses: W_violates_transparency, W_not_additive, W_not_affine in Meta/PolyInterpretation_FullStep.lean.
4.5 The DP escape
Theorem 4.22 (DP projection escapes the impossibility).
Define $\pi:\mathrm{Trace}\to\mathbb{N}$ by:
$\pi(\mathrm{void})=0,\quad\pi(\delta\,t)=\pi(t)+1,\quad\pi(\mathrm{rec}\Delta\,b\,s\,n)=\pi(n),\quad\pi(t)=0\text{ otherwise.}$
Then:
  1. 1.
    $\pi$ orients rec_succ: $\pi(\mathrm{app}\;s\;(\mathrm{rec}\Delta\;b\;s\;n))<\pi(\mathrm{rec}\Delta\;b\;s\;(\delta\;n))$ for all $b,s,n$.
  2. 2.
    $\pi$ violates the app-subterm axioms: $\pi(\mathrm{app}\;(\delta\,\mathrm{void})\;\mathrm{void})=0\not>1=\pi(\delta\,\mathrm{void})$.
[…]
The dependency-pair support layer has also been widened. Meta/DependencyPairs_FiniteGraph.lean reuses the finite-SCC search surface on any finite decidable dependency-pair relation and repackages a discovered SCC directly as the standard SCCCycle. Meta/DependencyPairs_CallGraph.lean and Meta/DependencyPairs_ExtractedCallGraph.lean remove the manual relation and finite-type packaging layers. Meta/DependencyPairs_TPDBExtraction.lean extracts the array-backed call graph from a finite TPDB-style rule list by computing defined heads and right-hand-side call heads, while Meta/DependencyPairs_FirstOrderExtraction.lean lifts the same step to a generic finite first-order TRS layer over arbitrary symbol and variable types.
The remaining frontend layers strip away packaging one stage at a time. Meta/DependencyPairs_FirstOrderFrontend.lean accepts an arbitrary finite rule-record array together with lhs / rhs extractors and reexports the same call-graph plus SCC-search / contradiction surface. Meta/DependencyPairs_FirstOrderEngine.lean packages that interface as a single internal engine object carrying Rule, rules, lhs, and rhs; it also bundles the typeclass-level first-order view and the raw-term bridge for engines that keep their own syntax. Meta/DependencyPairs_HeadView.lean works with the smaller head-view interface when only dependency-pair extraction matters. Meta/DependencyPairs_FiniteCarrierView.lean collects the finite rule-carrier, finite raw rule-carrier, finite head-carrier, and finite extracted-data carrier bridges in one module. Finally, Meta/DependencyPairs_KernelFirstOrder.lean packages the full KO7 root TRS directly over an internal symbol type and feeds that stack through the raw-term, head-view, and finite-carrier layers.
Proposition 4.23 (The extracted DP pair problem already has a simple linear base order).
There exists a linear polynomial-style base order $\mu:T\to\mathbb{N}$ such that every extracted KO7 dependency pair decreases strictly under $\mu$. Consequently, there is no blanket theorem saying that dependency pairs still fail whenever the chosen base order remains affine or polynomial.
The Lean proof is in Meta/DP_BaseOrder_Boundary.lean (counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order, no_blanket_dp_poly_base_barrier). The point is boundary-setting rather than escape construction: modularity alone is not enough to force failure, but neither does it trivialize the direct barrier story. What matters is the transformed problem and the structure imported by the chosen base order.
4.6 Ablation: removing duplication dissolves the barrier
[…]
$\mathrm{recL}\,b\,s\,(\delta\,n)\;\to\;\mathrm{recL}\,b\,s\,n$
where $s$ is not duplicated. The simpleSize measure (Tier 1, all constructor weights equal to 1) strictly orients both rules:
$\mathrm{simpleSize}(\mathrm{recL}\,b\,s\,(\delta\,n))$ $=1+\mathrm{simpleSize}(b)+\mathrm{simpleSize}(s)$
$\quad+1+\mathrm{simpleSize}(n)$
$>1+\mathrm{simpleSize}(b)+\mathrm{simpleSize}(s)$
$\quad+\mathrm{simpleSize}(n)$
$=\mathrm{simpleSize}(\mathrm{recL}\,b\,s\,n).$
Formalized in Meta/LinearRec_Ablation.lean. Removing duplication reduces the problem to a simple additive size measure: the $\delta$-wrapper weight alone drives the decrease, with no need for the multiset or phase-bit components of $\mu^{3}_{c}$. This confirms that step-duplication is the operative source of the core immediate-step barrier theorems in this paper. Symmetric removal theorems for conditions (1) through (3) of §4 are not provided.
Directed wrapper-collapse surrogate.
We also tested a bounded first-order surrogate for the “no axioms inside the object language” condition by extending KO7 with the single directed rule
$\mathrm{app}(t,t)\to t.$
This relaxation does not dissolve the direct barrier. The added collapse rule is harmless in isolation. It is strictly oriented by simpleSize, and the nonlinear full-step witness $W$ still proves root-step termination of the extended relation. However, the additive, transparent-compositional, and affine barrier theorems continue to apply because the duplicating rule rec_succ is unchanged. The Lean development is in Meta/ObjectAxiom_Ablation.lean; the relevant identifiers are simpleSize_orients_app_collapse, wf_StepCollapseRev, no_global_stepCollapse_orientation_additive, no_global_stepCollapse_orientation_compositional_transparent, and no_global_stepCollapse_orientation_affine_of_unbounded. So this particular directed collapse surrogate is not the missing escape mechanism: it changes the system, but it does not neutralize the duplication obstruction isolated above. It is only a first-order syntactic proxy for collapse inside the rewrite language, not a replacement for full rewriting modulo an equational theory.
A separate artifact-only surrogate in Meta/SharingBarrierLift.lean makes the complementary point for condition (2) of §4: once the duplicated payload is represented through explicit sharing rather than copied in the tree, the tree-style direct barrier no longer applies verbatim. The module packages this as a positive counterexample (sharedCounter_orients_step) and a boundary theorem (sharing_breaks_tree_barrier), so the no-sharing hypothesis is not merely presentational. Correspondingly, any higher-order lift that remains tree-like and forbids such sharing should still be expected to inherit the same duplication pressure; proving that conditional survival statement is outside the present first-order artifact.
Alternating-cycle composite-step barrier.
Theorem 4.24 (Alternating-cycle composite-step barrier).
Consider a two-node mutually recursive SCC with rules
$\mathrm{recurA}(b,s,\mathrm{succ}(n))$ $\to\mathrm{wrap}(s,\mathrm{recurB}(b,s,n)),$
$\mathrm{recurB}(b,s,\mathrm{succ}(n))$ $\to\mathrm{wrap}(s,\mathrm{recurA}(b,s,n)).$
This theorem is not about non-duplicating local rules: each root rule already duplicates the payload once, since the same term $s$ appears both as the left wrapper argument and again as the recursive step argument on the right-hand side. Its content is that the two-node SCC yields a stable two-step composite profile and an induced minimal context relation on which the additive and affine barrier arguments still go through. Concretely, one root step followed by one step under the right argument of wrap passes through the intermediate trace $\mathrm{wrap}(s,\mathrm{recurB}(b,s,\mathrm{succ}(n)))$ and realizes the effective pattern
$\mathrm{recurA}(b,s,\mathrm{succ}(\mathrm{succ}(n)))\;\leadsto\;\mathrm{wrap}(s,\mathrm{wrap}(s,\mathrm{recurA}(b,s,n))).$
Then no additive direct measure on the shared constructors can orient the composite pattern uniformly. Equivalently, no such additive measure globally orients the induced minimal context relation generated by root steps and right-wrapper descent. Under the same kind of unbounded-pump hypothesis used in Theorem 4.2, no affine constructor-local direct measure can orient the same composite pattern uniformly either. The corresponding global affine failure for the induced minimal context relation is mechanized in Meta/MutualDuplication_General.lean; the exact identifiers appear in Table LABEL:tab:claim-code-barrier.
The concrete AltTerm development in Meta/MutualDuplication_Case.lean remains the worked instance of this bounded SCC theorem. In that instance, the effective target carries three copies of the step payload measure against one copy on the source side, so the additive contradiction closes by the same coefficient comparison as in the schema barrier. The affine companion theorem uses the same two-step composite profile but repackages it as a derived duplication schema with doubled successor and nested wrapper, so the existing affine barrier applies under the corresponding unbounded-pump hypothesis. The delayed-duplication branch then extends to finite cyclic ‘k+1‘-node SCCs in Meta/MutualDuplication_KNode.lean. That file generalizes the alternating two-node construction to local rules $\mathrm{recur}_{i}(b,s,\mathrm{succ}(n))\to\mathrm{wrap}(s,\mathrm{recur}_{i+1}(b,s,n))$ modulo the cycle length, proves exact one-cycle realization (cycle_realized), and rederives both additive and affine minimal-context barriers for the finite cyclic family (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_unbounded).
That delayed-duplication finite-cycle branch is now factored one layer further through an abstract one-cycle interface in Meta/MutualDuplication_CycleFlow.lean. The abstract source state is $\mathrm{recur}(b,s,\mathrm{succ}^{m}(n))$, the abstract target state is $\mathrm{wrap}^{m}(s,\mathrm{recur}(b,s,n))$, and a certified cycle witness supplies the exact contextual path between them. On top of that witness, the file packages four generic contextual obstructions: additive (no_global_orients_ctx_additive), affine under the derived unbounded-pump hypothesis (no_global_orients_ctx_affine_of_unbounded), transparent-compositional with successor transparency at the base point (no_global_orients_ctx_compositional_transparent), and a scalar-projection lift reducing richer codomain orders to an affine contradiction on the derived one-cycle schema (no_global_orients_ctx_of_scalar_projection_affine_of_unbounded).
The companion bridge file Meta/MutualDuplication_KNode_Abstract.lean packages the finite cyclic ‘k+1‘-node delayed SCC as an instance of that abstract interface and rederives the additive, affine, transparent-compositional, and projected matrix-style corollaries as no_global_orients_ctx_additive_via_cycleFlow, no_global_orients_ctx_affine_of_unbounded_via_cycleFlow, no_global_orients_ctx_compositional_transparent_via_cycleFlow, and no_global_orients_ctx_matrixFunctional_of_projected_unbounded_via_cycleFlow. So the finite cyclic syntax is now the concrete witness, not the theorem shape itself. The earlier Meta/MutualDuplication_Transparent.lean file remains the original two-node transparent/projection presentation of the same delayed-duplication mechanism.
The delayed branch also has a raw-graph formulation in Meta/MutualDuplication_GraphCycle.lean. That module fixes an arbitrary node type and edge relation, equips the graph-indexed recursor family with a certified closed cycle witness (CyclePath, toCycleWitness), and proves that any such raw graph instance inherits the same additive, affine, transparent-compositional, and scalar-projection contextual barriers as the finite cyclic presentation (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_unbounded, no_global_orients_ctx_compositional_transparent, no_global_orients_ctx_of_scalar_projection_affine_of_unbounded).
The utility files Meta/GraphPathExtraction.lean and Meta/FiniteGraphReachability.lean remove most of the remaining witness plumbing. Any nonempty closed transitive-closure proof or round-trip SCC witness is converted automatically into the required raw-graph cycle via ofClosedTransGen and ofRoundTrip, and for finite decidable graphs the wrappers no_global_orients_ctx_additive_of_finiteRoundTrip and its affine / transparent / scalar-projection companions work directly from bounded reachability facts rather than explicit path proofs. Meta/FiniteGraphSCC.lean then adds a finite search surface on node pairs: a candidate-pair finset sccPairs, an optional SCC-pair search function, and an equivalence between successful search and HasNontrivialSCC. From there the wrappers no_global_orients_ctx_additive_of_hasNontrivialSCC and its affine / transparent / scalar-projection companions recover the concrete witness pair and both reachability directions internally.
The relation-level wrapper file Meta/MutualDuplication_RelationalGraph.lean removes one more manual layer. Starting only from a node type, edge relation, shared constructor interface, and local edge-realization theorem, it builds the delayed raw-graph system automatically and reexports the same round-trip / finite-round-trip / HasNontrivialSCC wrappers together with direct variants that consume a successful finite-SCC search result. Two concrete frontends sit on top of that layer. Meta/MutualDuplication_CallGraph.lean starts from extracted call-graph data (‘nodeKey‘, ‘succKeys‘) on a finite node type and builds the same delayed SCC interface automatically, while Meta/MutualDuplication_ExtractedCallGraph.lean removes the finite-type packaging too: raw extracted node data stored in an array yields the node type ‘Fin nodes.size‘, the call graph, the finite SCC search surface, and then the delayed barrier wrappers. The delayed SCC obstruction is therefore packaged at ten levels: explicit finite cycle, abstract one-cycle witness, raw graph with a certified cycle object, raw graph with automatic cycle extraction from a closed-path witness, raw graph with finite-round-trip reachability wrappers, raw graph with finite-SCC witness packaging, raw graph with finite search-result wrappers, relation-level construction of the raw graph, call-graph construction from extracted successor-key data, and array-backed construction from raw extracted node data.
A multiplicity-preserving SCC theorem still requires a different mechanism: in first-order tree rewriting, once every rule preserves the relevant payload multiplicities, the current coefficient-growth argument no longer has a duplicated payload to pump against. Delayed branch: per-rule duplication, composite-step barrier. Preserving branch: per-rule multiplicity preserved, synchronized-exposure barrier.
Theorem 4.25 (Multiplicity-preserving synchronized SCC barrier).
There exists a two-node first-order SCC together with an explicit payload-counting function such that each individual root rule preserves payload multiplicity exactly. On synchronized inputs carrying the same latent payload on two tagged channels, one full SCC cycle realizes the contextual reduction
$\mathrm{recurA}(c,\mathrm{left}(s),\mathrm{right}(s))\;\leadsto\;\mathrm{wrap}(s,\mathrm{wrap}(s,\mathrm{recurA}(c,\mathrm{base},\mathrm{base}))).$
The source has no visible wrapper-carried payload, while the target exposes two visible wrapper copies of the synchronized payload and still preserves the counted multiplicity. Consequently no additive direct measure on this preserving SCC syntax can orient the synchronized composite uniformly. Equivalently, no such additive measure globally orients the induced minimal context relation generated by root steps and right-wrapper descent. Formalized in Meta/MutualDuplication_Preserving.lean as step_preserves_payloadCount, synchronized_cycle_realized, synchronized_cycle_exposes_payload, no_additive_orients_synchronized_cycle, and no_global_orients_ctx_additive.
This theorem is stronger than Theorem 4.24 in one precise sense: the local SCC rules no longer increase the tracked payload multiplicity at all. The obstruction comes from synchronized exposure, not from the immediate coefficient-growth profile used in the earlier alternating theorem.
The mechanization also proves a conditional affine extension: under a wrapper-dominance condition $r_{p}+r_{q}<w_{l}+w_{r}\cdot w_{l}$ on the affine coefficients (where $r_{p},r_{q}$ are the recursor’s payload scaling coefficients and $w_{l},w_{r}$ are the wrapper’s left and right scaling coefficients), no affine measure with an unbounded pump can orient the synchronized cycle. The proof pumps the payload evaluation past the source constant, then exploits the coefficient gap to force the target above the source. Proved as no_affine_orients_synchronized_cycle_of_wrapper_dominance in Meta/MutualDuplication_Preserving.lean.
A finite cyclic synchronized-packet extension appears in Meta/MutualDuplication_Preserving_KNode.lean. There the latent payload is carried by an explicit packet of indexed channels of length ‘k+1‘; one root step exposes the current packet head under wrap and forwards the remaining packet to the next cycle node. The file proves exact payload preservation for every local step (step_preserves_payloadCount), exact one-cycle realization (synchronized_cycle_realized), additive failure of the induced minimal context relation (no_global_orients_ctx_additive), and a conditional affine failure theorem under a generalized wrapper-dominance condition (no_global_orients_ctx_affine_of_wrapper_dominance). This is a narrowed explicit payload-flow model rather than a fully abstract SCC metatheorem, but it shows that the synchronized-exposure obstruction survives beyond the original two-node presentation.
That explicit packet model is then factored through an abstract payload-flow interface in Meta/MutualDuplication_PayloadFlow.lean. The companion bridge file Meta/MutualDuplication_Preserving_Abstract.lean packages the finite synchronized-packet SCC as an instance of that interface and rederives the additive and affine global barrier theorems from the abstract cycle lemmas no_global_orients_ctx_additive and no_global_orients_ctx_affine_of_wrapper_dominance. It also rederives the packet-transparent and projected packet-transparent corollaries via no_global_orients_ctx_packetTransparent_via_payloadFlow and no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent_via_payloadFlow. The packet syntax is therefore a concrete witness presentation, not the theorem shape itself: the barrier mechanism has been isolated as a generic one-cycle payload-flow metatheorem.
A narrow transparent-compositional downstream corollary is also available in Meta/MutualDuplication_Preserving_Transparent.lean. That module treats the synchronized packet constructors as evaluation-transparent bookkeeping and proves both the cycle-level contradiction no_compositional_orients_synchronized_cycle_packetTransparent and the induced minimal-context corollary no_global_orients_ctx_packetTransparent. It also adds a generic scalar-projection lift, no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent: any componentwise matrix-style orienter whose chosen weighted scalar projection collapses to one of these packet-transparent measures is blocked as well. The preserving finite-cycle branch thus admits both a direct transparent obstruction and a projection-based downstream corollary, without introducing a separate preserving-specific matrix syntax layer.
The preserving branch also has a raw-graph formulation in Meta/MutualDuplication_PacketGraph.lean. That module fixes an arbitrary node type and edge relation, equips the graph-indexed packet system with a certified closed cycle witness (CyclePath, toCycleWitness), and proves additive, affine, transparent, and scalar-projection contextual barriers directly for that raw graph presentation (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_wrapper_dominance, no_global_orients_ctx_transparent, no_global_orients_ctx_of_scalar_projection_transparent).
Here too the path-extraction and finite-reachability utilities remove most of the cycle bookkeeping. A nonempty closed transitive-closure proof or round-trip SCC witness builds the required cycle object automatically via ofClosedTransGen and ofRoundTrip; for finite decidable graphs the *_of_finiteRoundTrip wrappers apply directly from bounded reachability facts. The finite-SCC layer in Meta/FiniteGraphSCC.lean goes one step further: it exposes a finite pair-search surface built around sccPairs, proves that successful search is equivalent to HasNontrivialSCC, and lets the wrappers no_global_orients_ctx_additive_of_hasNontrivialSCC and its affine / transparent / scalar-projection companions recover the concrete pair and both reachability directions internally.
The same relation-level wrapper file, Meta/MutualDuplication_RelationalGraph.lean, builds the preserving raw-graph system from the smaller relation-level presentation and reexports the same round-trip / finite-round-trip / HasNontrivialSCC wrappers together with direct variants that consume a successful finite-SCC search result. On top of that, Meta/MutualDuplication_CallGraph.lean and Meta/MutualDuplication_ExtractedCallGraph.lean provide the same preserving barrier surface starting respectively from extracted call-graph data on a finite node type and from raw extracted node data stored in an array. The obstruction is therefore no longer tied to finite packet syntax, hand-written cycle indices or explicit transitive-closure witnesses, manual raw-graph packaging, hand-supplied SCC pairs or separate SCC-existence certificates once finite search succeeds, or a hand-built finite node type / call-graph layer for extracted dependency data.
4.7 Proof-theoretic context
Table 6: Five method classes applied to KO7. Every class that succeeds imports structural assumptions outside the formalized barrier (Remark 4.21). The proof-strength column refers to the general metatheory where marked.
Method class Orients KO7? Proof strength Mechanism / imported assumption
Direct/global aggregation (Def. 3.2) No typically arithmetical Aggregates subterms; duplication defeats orientation
Affine/linear constructor-local (Thm. 4.2) No typically arithmetical Coincides with Tier 1 when all scales are 1; independent under unbounded pump
Nonlinear polynomial interpretation Yes (L) PRA Cross-term coupling + non-transparent successor; outside Tier 1, Tier 2, and affine classes (W_orients_step)
DP + subterm criterion Yes elementary ($\ll\varepsilon_{0}$) Projects away duplication; violates wrapper-subterm axioms
[…]
KO7-specialized MPO Yes (L) direct ordinal ranking below $\varphi_{7}(0)$ Fixed 7-constructor signature, fixed precedence; explicit upper bound in Meta/MPO_ProofTheoreticBound.lean (wf_StepRev_mpo_below_veblen7)
[…]
Corollary 4.26 (Non-representability of successful Nat-valued orienters).
Let $\mu:T\to\mathbb{N}$ globally orient a step relation containing the duplicating rule. Then $\mu$ is not representable as
  1. 1.
    an additive schema measure,
  2. 2.
    a Tier 2 compositional measure with successor transparency at $\mathrm{base}$, or
  3. 3.
    an affine / linear constructor-local measure satisfying the unbounded-pump hypothesis of Theorem 4.2.
See Meta/StepDuplicatingSchema.lean.
Here and below, the schema-level API uses the prefix no_global_orients_*, while the KO7-instantiation API uses no_global_step_orientation_*. This is the precise sense in which path-order success lies outside the present Nat-valued measure framework: not because path orders are unsound, but because successful orienters cannot be represented by the formal classes isolated above.
Proposition 4.27 (Nonlinear polynomial full-step escape).
There exists a $\mathbb{N}$-valued interpretation $W:\mathrm{Trace}\to\mathbb{N}$ such that:
  1. 1.
    for every root step $a\to b$ in Step, one has $W(b)<W(a)$, hence the reverse root relation is well-founded;
  2. 2.
    $W$ violates $\delta$-transparency and is neither additive nor affine in the sense of the formalized barrier classes.
The Lean proof is in Meta/PolyInterpretation_FullStep.lean (W_orients_step, wf_StepRev_poly, W_violates_transparency, W_not_additive, W_not_affine).
In that interpretation the eqW clause is $W(\mathrm{eqW}\;a\;b)=W(a)+W(b)+3$. The constant +3+3 is tight: R_eq_diff needs $W(\mathrm{eqW}\;a\;b)>W(\mathrm{integrate}(\mathrm{merge}\;a\;b))=W(a)+W(b)+2$, while R_eq_refl needs $W(\mathrm{eqW}\;\mathrm{void}\;\mathrm{void})>W(\mathrm{void})=1$. So $3$ is the smallest uniform choice that makes both rules strictly decreasing.
Proposition 4.28 (KO7-specialized MPO full-step termination).
[…]
Proposition 4.29 (Proof-theoretic upper bound for the KO7-specialized MPO proof).
The fixed-signature KO7 MPO proof is witnessed by an ordinal ranking into the initial segment below $\varphi_{7}(0)$, that is, below Ordinal.veblen70. Here $\varphi_{\alpha}$ denotes the Veblen hierarchy, so $\varphi_{7}(0)$ lies below the Feferman–Schütte ordinal $\Gamma_{0}$. This gives a concrete proof-theoretic upper bound for the specialized MPO argument without appealing to generic path-order metatheory over arbitrary signatures. Proved as mpoOrd_lt_mpoVeblenBound and wf_StepRev_mpo_below_veblen7 in Meta/MPO_ProofTheoreticBound.lean.
Proposition 4.30 (Bad precedence blocks the duplicating rule).
[…]
$\mathrm{rec}\Delta(\mathrm{void},\mathrm{void},\delta(\mathrm{void}))\to\mathrm{app}(\mathrm{void},\mathrm{rec}\Delta(\mathrm{void},\mathrm{void},\mathrm{void})).$
Consequently that bad precedence cannot orient all root steps of Step. See Meta/MPO_Precedence_Barrier.lean (not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec).
This supplies the negative companion to Proposition 4.28: KO7-specialized MPO success is genuine, but it is also precedence-sensitive.
Theorem universe. The following theorem is stated over an explicit direct universe consisting of twelve listed families, namely the twelve-family trichotomy universe distinct from the twelve-family barrier package discussed above: additive compositional, transparent-compositional, pumped affine, pumped restricted-quadratic, pumped bounded cross-term quadratic, pumped bounded multilinear, pumped generalized bounded-polynomial, pumped max-plus, tracked-primary componentwise pair, tracked-primary lexicographic pair, KO7-specific max-depth, and KO7-specific pure head-precedence. For the pair-valued families, the theorem reads their first component as the tracked primary scalar. The Lean development first proves the scalar-only preliminary ko7_nat_direct_escape_trichotomy; the theorem stated here is the KO7-specific extension ko7_direct_escape_trichotomy_extended, which adds the tracked-primary pair families. A second theorem, ko7_projection_escape_trichotomy, extends the same classification pattern to successful componentwise orienters that admit a fixed scalar projection whose strict decrease is forced by the ambient order and whose projected scalar lies in the listed Nat-valued universe. This is how the weighted scalar-projection and balanced mixed-coordinate matrix families are absorbed into the trichotomy layer. The generalized degree-bounded polynomial family of Theorem 4.6 is now internalized directly into the Nat-valued universe; projection-based matrix families still enter through the separate projection theorem. So the twelve-family barrier package and the twelve-family trichotomy universe are not identical: the former lists the projection-based matrix families directly, while the latter replaces them by KO7-specific max-depth and pure head-precedence and then reintroduces the projection families through the separate representability theorem. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside this theorem universe.
Table 7: Barrier-package families versus the trichotomy theorem universe.
Shared by both universes Barrier package only Trichotomy universe only
Additive, transparent-compositional, pumped affine, pumped restricted-quadratic, pumped bounded cross-term quadratic, pumped bounded multilinear, pumped generalized bounded-polynomial, pumped max-plus, and the tracked-primary componentwise / lexicographic pair families Projection-based matrix families, reintroduced later through the separate projection theorem KO7-specific max-depth and KO7-specific pure head-precedence
Theorem 4.31 (Escape trichotomy for the explicit direct universe).
Let $\mathcal{O}$ be a successful direct root-step orienter in the explicit theorem universe above, and let $\pi_{\mathcal{O}}:\mathrm{Trace}\to\mathbb{N}$ denote its tracked primary scalar, namely the orienter itself in the scalar cases and the first component in the tracked-pair cases. Then at least one of the following holds:
  1. 1.
    $\pi_{\mathcal{O}}$ fails wrapper-subterm sensitivity for app;
  2. 2.
    $\pi_{\mathcal{O}}$ fails successor transparency at void;
  3. 3.
    $\mathcal{O}$ is not representable by any of the direct families formalized in the artifact: additive compositional, transparent-compositional, affine-with-pump, restricted-quadratic-with-pump, bounded-cross-term-with-pump, bounded-multilinear-with-pump, generalized bounded-polynomial-with-pump, max-plus-with-pump, tracked-primary componentwise pair-with-pump, tracked-primary lexicographic pair-with-pump, max-depth, or pure head-precedence.
The third disjunct is the explicit twelve-family universe enumeration just listed. Equivalently, every successful direct root-step orienter in the theorem universe escapes by violating wrapper sensitivity on its tracked primary scalar, violating transparency there, or leaving the formalized direct families. Formalized in Meta/EscapeTrichotomy.lean (ko7_direct_escape_trichotomy_extended; scalar-only preliminary: ko7_nat_direct_escape_trichotomy).
Projection-based extension. The same escape pattern also covers the weighted scalar-projection componentwise family and the balanced mixed-coordinate componentwise family. The reason is structural rather than ad hoc: if the ambient componentwise order forces strict decrease of a fixed scalar projection, and that projected scalar already belongs to the explicit Nat-valued direct universe above, then the trichotomy applies to that scalar as well. The Lean proof is ko7_projection_escape_trichotomy in Meta/EscapeTrichotomy.lean, with the concrete representability lemmas matrixFunctional_projection_representable and matrixMix2_sum_projection_representable.
Scope note. The theorem is intentionally restricted to a finite, reviewable direct universe plus a separate projection-based extension principle. This is a deliberate design choice. A theorem over a listed class is verifiable and falsifiable, while an unbounded claim about all imaginable orienters would be either tautological, by defining “outside the formalized classes” to absorb every counterexample, or unprovable with current methods. This is therefore a classification theorem for the listed universe, not a dichotomy between success and failure: the first two disjuncts are testable structural predicates, while the third localizes the remaining open space outside the formalized direct families. The bounded cross-term, bounded multilinear, generalized bounded-polynomial, and max-plus scalar families are included directly; the weighted scalar-projection and balanced mixed-coordinate matrix families are included through the projection theorem; and the current matrix-side strengthening now also exposes fixed-row and row-sum projection corollaries in Meta/MatrixProjectionCoverage.lean. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside the present theorem universe, even though the paper separately proves negative or positive results for some of them.
4.8 Scope
Theorem 4.1 covers the four-role schema and the two measure classes of Definition 3.2. Theorem 4.2 adds constructor-local affine/linear measures under an explicit unbounded-growth hypothesis. The nonlinear scalar extensions cover restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized degree-bounded polynomial, and max-plus classes; the first two, the generalized bounded-polynomial family, and the multilinear / max-plus families also admit pumped strengthened subclasses in Meta/PumpedBarrierClasses.lean. The vector / pair layer covers fixed-dimension tracked componentwise orderings, dimension-$2$ tracked lexicographic orderings, balanced mixed-coordinate dimension-$2$ componentwise order, and weighted functional componentwise order. The escape trichotomy now classifies the scalar multilinear, generalized bounded-polynomial, and max-plus families directly and absorbs the weighted-functional and balanced mixed-coordinate families through the projection-based extension theorem. The matrix-side coverage is also sharper than before: Meta/MatrixProjectionCoverage.lean makes the fixed-row and row-sum corollaries explicit for the tracked-component and weighted-functional barrier layers. Out of scope are generic DP metatheory, arbitrary mixed matrix interpretations beyond these tracked, balanced, functional, fixed-row, and row-sum cases, arbitrary semantic methods, and duplicating-rule shapes not reduced to a single wrapper. In particular, nonlinear polynomial interpretations with cross-term coupling (such as the witness in Remark 4.21) remain outside the barrier, because they import nonlinearity and non-transparency not present in the formalized direct classes.
Status of full Step. Root-step Step is terminating: mechanized via the nonlinear polynomial interpretation (wf_StepRev_poly) and the KO7-specialized MPO (wf_StepRev_mpo). The full context closure is also strongly normalizing in Lean via wf_StepCtxFullRev_poly, and independently via TTT2/CeTA (§6). Full Step is not locally confluent.
Proposition 4.32 (The unguarded eqW overlap is never locally joinable).
For every trace $a$, the full kernel relation fails local join at $\mathrm{eqW}\;a\;a$: the two one-step reducts $\mathrm{void}$ and $\mathrm{integrate}(\mathrm{merge}\;a\;a)$ are distinct full-step root normal forms.
Proved as not_localJoinStep_eqW_refl in Meta/EqW_Guard_Barrier.lean. The second reduct is in normal form for a simple structural reason: R_int_delta only fires on $\mathrm{integrate}(\delta\,t)$, while $\mathrm{merge}\;a\;a$ is never $\delta$-wrapped, so no rule can act on $\mathrm{integrate}(\mathrm{merge}\;a\;a)$. This is why the SafeStep eqW guards are genuine confluence guards rather than presentation artifacts.
Frontier summary.
The barrier section isolates a concrete frontier. Inside the formalized direct classes, the duplicating step is excluded by the schema, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, max-plus, tracked componentwise fixed-dimension vector, tracked lexicographic pair, balanced mixed-coordinate pair, and weighted scalar-projection componentwise barriers, together with the symbolic variable-condition / KBO corollary. The trichotomy theorem classifies the scalar side through max-plus and multilinear, and extends to the weighted-functional and balanced mixed-coordinate matrix families through projected representability, while the generalized polynomial theorem isolates the remaining bounded direct-polynomial escape condition. Crossing the frontier therefore requires one of three moves exhibited in the artifact: projection-style transformation (DP), direct nonlinear escape outside the formalized classes (the global witness $W$ and the KO7-specialized MPO), or ablation of the offending structure itself (removing duplication dissolves the barrier, while a directed wrapper-collapse surrogate does not). The SCC theorems show that this frontier is not confined to immediate single-rule duplication. Delayed duplication across finite cyclic SCCs, and more generally across graph-indexed SCC families once a nonempty closed path, round-trip SCC witness, or bounded finite-round-trip reachability package is supplied, recreates the additive and affine obstruction. Even multiplicity-preserving synchronized cycles still force additive failure, with a conditional affine companion and transparent / scalar-projection downstream corollaries, once the latent packet is exposed under visible wrappers.
4.9 Orientation catalog
The theorems at the schema subsume most individual measure witnesses. Items such as additive counters, simple lex pairs, node counts, and linear weights are direct instances of Theorems 4.1 and 4.2. Tree depth and pure precedence are covered by dedicated KO7-specific families in Meta/DepthBarrier.lean and Meta/PrecedenceBarrier.lean. The following catalog entries record results not subsumed by those theorems or by those two KO7-specific families:
  1. 1.
    Polynomial tie (L/T): The multiplicative polynomial $M(\mathrm{rec}\Delta\,b\,s\,n)=M(b)+M(s)\cdot M(n)$ ties exactly on rec_succ (poly_mul_ties_rec_succ). TTT2 POLY returns MAYBE.
  2. 2.
    Multiset / Constellation (L): Single-pass multisets over nesting ranks fail (no_constellation_strict_drop_rec_succ).
  3. 3.
    Semantic Interpretation ($\dagger$): Model-theoretic methods are blocked by the duplicating constructor condition (§4).
  4. 4.
    Path Orders ($\dagger$/L): LPO orients the full TRS (CeTA-certified); this is a framework escape from the $\mathbb{N}$-valued measure classes (Cor. 4.26). Meta/MPO_FullStep.lean proves KO7-specialized MPO orientation (wf_MPORev, wf_StepRev_mpo); this is a fixed-signature comparator, not a generic MPO metatheorem.
Here (L) means Lean-checked inside the artifact, (T) records the corresponding external TTT2 outcome when one is reported, and $\dagger$ marks theory-level or tool-level remarks rather than a new standalone Lean family theorem.
The DP framework escapes by projecting to the recursion counter, as formalized in Theorem 4.22, and Proposition 4.23 shows that the extracted pair problem already admits a simple linear polynomial-style base order. Regarding polynomial interpretations more broadly: Theorem 4.2 covers the affine/linear fragment under positive wrapper coefficients and an unbounded pump, and Theorem 4.6 pushes that boundary to bounded degree-bounded polynomial tables with repeated variables under frozen base dominance, but neither theorem extends to arbitrary nonlinear polynomials. Theorem 4.20 gives a concrete nonlinear witness that overcomes duplication via cross-term coupling; any such polynomial necessarily imports structural assumptions outside the formalized barrier classes (Remark 4.21).
4.10 Gallery of explicitly excluded methods
Beyond the schema-level theorems and the catalog items above, Meta/Conjecture_Boundary.lean collects a broader mechanized no-go catalog for the full unrestricted root relation Step. Two stand out.
Exact polynomial tie (poly_mul_ties_rec_succ).
A polynomial measure with multiplicative recursor combiner $M(\mathrm{rec}\Delta(b,s,n))=M(b)+M(s)\cdot M(n)$ and additive app yields an exact equality on the duplicating step, regardless of base weight:
[…]
Formalized as poly_mul_ties_rec_succ. This is not merely non-strict failure but exact cancellation: the duplication of $s$ under app is algebraically absorbed by the multiplicative coupling. Only nonlinear cross-variable coupling (as in $W$) breaks this tie.
Dual barrier (conflicting structural demands).
The duplication barrier and the flag barrier target different rules and impose contradictory demands on any single $\mathbb{N}$-valued measure: rec_succ requires insensitivity to duplication of $s$ (additive size fails), while merge_void requires sensitivity to flag structure (the $\delta$-flag must strictly decrease). See dual_barrier_rec_succ_and_merge_void.
Further exclusions.
The remaining no-go results sharpen the negative catalog. Strictly monotone post-processing does not rescue an additive measure, and a single $\delta$-flag cannot orient Step. Nesting depth with merge-neutral aggregation ties on merge_cancel, while standard tree depth strictly increases on deep instances of rec_succ. Node count, pure head precedence, linear KBO weights, and constellation-size measures also fail on the unrestricted root relation. The exact theorem identifiers are listed in Meta/Conjecture_Boundary.lean and indexed in the appendix tables.
Decidability of reachability
In a strongly normalizing and confluent calculus with a computable normalizer and decidable equality, the set $\{t\mid t\Rightarrow^{*}c\}$ for any normal-form target $c$ is decidable: compute $\mathrm{normalizeSafe}(t)$ and compare. This is formalized as the Decidable instance reachability_decidable in Meta/Newman_Safe.lean: given $c$ in safe normal form, the predicate $t\mapsto\texttt{SafeStepStar}\;t\;c$ is a DecidablePred, reducing reachability to a single normalization followed by equality comparison. Complexity bounds for related TRS classes are discussed in [3, 7, 16].
5 KO7 certification chain
The guarded SafeStep fragment of KO7 is terminating, root-confluent, and normalizable. This section certifies those properties.
5.1 Strong normalization (SafeStep)
The termination argument uses a triple-lexicographic measure
$\mu^{3}_{c}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\tau(t))$
ordered lexicographically over $(\mathbb{N},\,\mathrm{DM}(\mathrm{Multiset}\;\mathbb{N}),\,\mathbb{N})$:
  • $\delta\text{-flag}(t)=1$ if $t=\mathrm{rec}\Delta\,b\,s\,(\delta\,n)$, else 0.
  • $\kappa^{M}(t)$: multiset of natural numbers, recursively defined; each $\mathrm{rec}\Delta(b,s,n)$ contributes $\mathrm{weight}(n)+1$.
  • $\tau(t)$: weighted node-size with per-constructor base weights; $\tau(\delta\,t)=\tau(t)$.
The components are individually standard: lexicographic arguments originate with Floyd [14]; the multiset ordering is due to Dershowitz and Manna [9]; phase bits appear in DP analyses [2]. The contribution is the instantiation to KO7’s eight rules and the Lean 4 mechanization.
Theorem 5.1 (Per-step decrease).
For every $t\Rightarrow t^{\prime}$ in SafeStep, $\mu^{3}_{c}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}_{c}(t)$.
Proof. By exhaustive case analysis over the eight SafeStep constructors, discharged in Lean as measure_decreases_safe_c in Meta/ComputableMeasure.lean. For rec_succ, $\delta$-flag drops ($1\to 0$). For rec_zero and the guarded merge-cancel branch, $\kappa^{M}$ drops in the Dershowitz–Manna order. In the remaining guarded merge / eqW / integrate cases, the first two coordinates tie and the weighted size $\tau$ strictly decreases.
Corollary 5.2 (Strong normalization).
SafeStep is strongly normalizing.
[…]
$\texttt{ctxFuel}(\texttt{void})$ $=0,$
$\texttt{ctxFuel}(\delta\,t)$ $=2^{\,\texttt{ctxFuel}(t)+1},$
$\texttt{ctxFuel}(\texttt{integrate}\,t)$ $=\texttt{ctxFuel}(t)+1,$
$\texttt{ctxFuel}(\texttt{merge}(a,b))$ $=\texttt{ctxFuel}(a)+\texttt{ctxFuel}(b)+2,$
$\texttt{ctxFuel}(\texttt{app}(a,b))$ $=\texttt{ctxFuel}(a)+\texttt{ctxFuel}(b)+1,$
$\texttt{ctxFuel}(\texttt{rec}\Delta(b,s,n))$ $=2^{\,\texttt{ctxFuel}(n)+\texttt{ctxFuel}(s)+5}+\texttt{ctxFuel}(b)+1,$
$\texttt{ctxFuel}(\texttt{eqW}(a,b))$ $=\texttt{ctxFuel}(a)+\texttt{ctxFuel}(b)+4.$
[…]
Proposition 5.3 (Confluence of SafeStepCtx).
For the partial context closure SafeStepCtx, the remaining global local-join obligation is discharged, so SafeStepCtx is confluent. The Newman characterization is retained as a corollary:
$\texttt{ConfluentSafeCtx}\quad\Longleftrightarrow\quad\forall a,\;\texttt{LocalJoinCtxAt}\;a.$
Since strong normalization is already proved, this gives both an unconditional contextual confluence theorem and the corresponding SN-plus-local-join characterization for the partial context closure.
See Meta/SafeStepCtx_Confluence.lean (localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, and the specialized Newman direction newman_safeCtx).
A second confluence result lands one level up from the guarded fragment. The module Meta/EqGuardedConfluence.lean keeps every root rule of Step except the reflexive eqW-diff overlap and proves the resulting intermediate root relation confluent (confluentEqGuarded). This shows that the full-root confluence failure is concentrated exactly at the unguarded eqW peak rather than in the other root rules.
Theorem 5.4 (Full context-closed strong normalization).
The full contextual closure of the unguarded root relation Step is strongly normalizing.
The Lean proof (wf_StepCtxFullRev_poly in Meta/ContextClosed_SN_Full.lean) lifts the nonlinear polynomial witness $W$ through every constructor position and shows that every contextual contraction still strictly decreases $W$.
Proposition 5.5 (Linear derivation-length bound from the global witness).
If $t\Rightarrow_{\mathrm{ctx}}^{n}u$ is a full contextual reduction of the unguarded system with exactly $n$ steps, then $n+W(u)\leq W(t)$. In particular, every such derivation has length strictly below $W(t)$.
Formalized in Meta/SafeStep_Complexity.lean (stepCtxFullPow_length_le_W, stepCtxFullPow_length_lt_W). The result is a concrete complexity theorem for the full contextual relation; the ordinal discussion below remains a separate proof-theoretic calibration for the guarded SafeStep argument.
5.2 Certified normalizer
A normalization function is defined by well-founded recursion on the measure-pullback relation. It operates on root SafeStep redexes only. Properties:
(Totality) $\forall t\,\exists n.~\mathrm{SafeStepStar}(t,n)\land\mathrm{NormalFormSafe}(n)$
(Soundness) $\forall t.~\mathrm{SafeStepStar}(t,\mathrm{normalizeSafe}(t))\land\mathrm{NormalFormSafe}(\mathrm{normalizeSafe}(t)).$
Note that normalizeSafe computes normal forms with respect to the guarded relation SafeStep, not the full unguarded Step; a SafeStep-normal form may still admit unguarded reductions. Moreover, SafeStep is deterministic at the root (safeStep_deterministic in Meta/NormalizeSafe_LowerBound.lean), so the certified normalizer follows the unique possible guarded root reduction at each step. Under confluence, reachability reduces to normalization and equality check. For safe normal-form targets, the resulting decision procedure also carries an explicit certified cost envelope: Meta/Reachability_Complexity.lean proves both the linear size upper bound reachabilityDecisionCost_le_linear_termSize and the coarser inherited tower bound reachabilityDecisionCost_le_complexity_bound, together with the matching linear families reachabilityDecision_has_linear_lower_family and reachabilityDecision_has_linear_size_family.
Proposition 5.6 (Exact normalization cost on a merge spine).
Let $m_{0}:=\mathrm{void}$ and $m_{n+1}:=\mathrm{merge}\;\mathrm{void}\;m_{n}$. The certified normalizer returns $\mathrm{void}$ on this family, and the auxiliary counter normalizeSafeSteps, defined by the same deterministic recursion as normalizeSafe, satisfies $\mathrm{normalizeSafeSteps}(m_{n})=n$.
See Meta/NormalizeSafe_LowerBound.lean (mergeVoidChain_pow_void, normalizeSafe_mergeVoidChain, normalizeSafeSteps_mergeVoidChain). This gives a concrete linear lower bound for the cost of the certified root normalizer on one explicit family.
Proposition 5.7 (Certified root-normalizer upper envelope).
For every term $t$, the deterministic root-step counter normalizeSafeStepstt is realized by an exact guarded root reduction path and satisfies the linear size bound
$\texttt{normalizeSafeSteps}(t)+2\;\leq\;2\,\mathrm{termSize}(t).$
Hence it is also bounded above by the older explicit size-based tower bound inherited from the coarse SafeStepCtx analysis.
The Lean proofs are in Meta/SafeRoot_Complexity.lean; the exact theorem identifiers are listed in the claim-to-code appendix. This yields a fully mechanized algorithmic envelope for the certified root normalizer: an exact linear lower-bound family on merge spines, a linear upper bound in structural term size, and the coarser inherited ctxFuel / cichon corollaries.
In particular, the guarded root relation has a concrete linear lower-bound family for derivation length: each step from $m_{n+1}$ to $m_{n}$ removes exactly one outer mergevoid layer, so the family grows by one constructor layer per step while requiring exactly $n$ guarded root contractions, and $\mathrm{termSize}(m_{n})=2n+1$. Thus the certified root normalizer has linear complexity in the structural size of its input. The coarse ctxFuel and finite fast-growing envelopes remain useful explicit corollaries, and the artifact also exposes the root and guarded-context upper bounds through Cichon-style theorems in the same lex3Note family used by the ordinal calibration.
5.3 Root-step confluence via Newman
We instantiate Newman’s lemma for SafeStep and discharge the local-join hypothesis $\forall\,a,\;\mathrm{LocalJoinAt}\;a$ by case analysis over all root and context shapes. Generic Newman frameworks exist in Lean 4 [23]; the contribution here is the KO7-specific instantiation.
Theorem 5.8 (Confluence and unique normal forms).
The ARS $(\mathrm{Trace},\,\texttt{SafeStep})$ is confluent. Every term reduces to a unique normal form.
Critical-pair coverage
Table 8: Critical-pair coverage at the root for SafeStep.
Source Lemma
$\mathrm{integrate}(\delta\,t)$ localJoin_int_delta
$\mathrm{merge}\;\mathrm{void}\;t$ localJoin_merge_void_left
$\mathrm{merge}\;t\;\mathrm{void}$ localJoin_merge_void_right
$\mathrm{merge}\;t\;t$ localJoin_merge_tt
$\mathrm{rec}\,\Delta\,b\,s\,\mathrm{void}$ localJoin_rec_zero
$\mathrm{rec}\,\Delta\,b\,s\,(\delta\,n)$ localJoin_rec_succ
$\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=\varnothing$ localJoin_of_unique
The side condition $\kappa^{M}(t)=\varnothing$ on R_merge_cancel is part of the guarded confluence design, not a cosmetic strengthening: when $\kappa^{M}(t)\neq\varnothing$, the collapsing root step is blocked, and the proof switches to the blocked-guard case localJoin_merge_cancel_guard_kappa_ne. The collapsing case itself is handled separately by localJoin_merge_tt. This coupling between the termination measure component $\kappa^{M}$ and the confluence guard is a deliberate design choice: the guarded fragment uses the same multiset infrastructure for both termination (the second component of $\mu^{3}_{c}$) and confluence (the merge-cancel guard). In KO7 the same multiset is exactly what distinguishes harmless collapsing cases from forks that would reintroduce the bad payload, so separating the two would duplicate the same structural case split rather than simplify it. Concretely, if $\mathrm{merge}\;t\;t\to t$ were allowed while $t$ still contained an active $\mathrm{rec}\Delta$ subterm, the two copies could unfold differently before any later merge, producing distinct descendants instead of a trivially joinable collapse.
5.4 Ordinal upper-bound calibration
Theorem 5.9 (Upper bound (mechanized)).
The modules Meta/DM_OrderType.lean and Meta/DM_OrderType_LowerBound.lean together prove:
  1. 1.
    $\forall m:\mathrm{Multiset}\;\mathbb{N},\;\texttt{dmOrdEmbed}(m)<\omega^{\omega}$.
  2. 2.
    $\texttt{DM}\;m_{1}\;m_{2}\Rightarrow\texttt{dmOrdEmbed}(m_{1})<\texttt{dmOrdEmbed}(m_{2})$.
  3. 3.
    $\forall t:\mathrm{Trace},\;\texttt{lex3cToOrd}(\mu_{3c}(t))<\omega^{\omega}\!\cdot\!2<\varepsilon_{0}$.
  4. 4.
    $\texttt{SafeStep}\;a\;b\Rightarrow\texttt{lex3cToOrd}(\mu_{3c}(b))<\texttt{lex3cToOrd}(\mu_{3c}(a))$.
Remark 5.10 (Tight order type; mechanized).
The inner product $(\mathrm{DM},\,\mathbb{N})$ has exact order type $\omega^{\omega}$, and the surrounding two-block carrier $(\{0,1\},\,\mathrm{DM}(\mathrm{Multiset}\;\mathbb{N}),\,\mathbb{N})$ has order type $\omega^{\omega}\!\cdot\!2$. Both $\omega^{\omega}\!\cdot\!2$ and $\omega^{\omega+1}$ (for ambient $\mathbb{N}\times(\mathrm{DM}\times\mathbb{N})$) are below $\varepsilon_{0}$. The theorem dm_order_type_omega_omega bundles the embedding, reflection, bound, and surjectivity into one mechanized statement; lex3c_order_type_bound lifts this to the trace bound $\omega^{\omega}\!\cdot\!2$. The exact $\omega^{\omega}$ calibration concerns the DM component itself; the $\omega^{\omega}\!\cdot\!2$ statement is an upper bound for the triple-lex image used by the SafeStep certificate, not an exact order-type claim for SafeStep reductions.
Theorem 5.11 (Lower bound (mechanized, unconditional)).
The module Meta/DM_OrderType_LowerBound.lean proves:
  1. 1.
[…]
  2. 2.
    Surjectivity below $\omega^{\omega}$ for DM embedding.
  3. 3.
    Order reflection into DM; rank-bridge equality dmOrdEmbed_eq_dmRankOrd.
  4. 4.
    Surjectivity of DM rank below $\omega^{\omega}$.
This yields a mechanized order-type isomorphism: $(\mathrm{Multiset}\;\mathbb{N},\mathrm{DM})\cong\omega^{\omega}$ (cf. [9]). The exact statement and the cofinal bridge appear in the theorem names phaseB_cnf_scaffold_exact_order_type and phaseB_cnf_scaffold_cofinal.
Proposition 5.12 (Realized lower families for the full $\mu_{3c}$ image).
For every $n\in\mathbb{N}$, there exist traces $t_{0},t_{1}$ such that
$\omega^{n+2}\;\leq\;\texttt{lex3cToOrd}(\mu_{3c}(t_{0}))\;<\;\omega^{\omega}$
and
$\omega^{\omega}+\omega^{n+2}\;\leq\;\texttt{lex3cToOrd}(\mu_{3c}(t_{1}))\;<\;\omega^{\omega}\cdot 2.$
Formalized in Meta/Mu3c_Image_LowerBound.lean as payloadTower_mu3c_lower, payloadTower_mu3c_lt_block_cap, flaggedTower_mu3c_lower, flaggedTower_mu3c_lt_two_block_cap, mu3c_lower_family_block0, and mu3c_lower_family_block1. This does not upgrade the full $\mu_{3c}$ story to an exact order-type theorem, but it does show that the realized trace image is nontrivial in both $\delta$-blocks and reaches arbitrarily high ordinals below the established $\omega^{\omega}\!\cdot\!2$ cap.
Remark 5.13 (Proof-theoretic context of the bound).
The certified upper bound $\omega^{\omega}\!\cdot\!2<\varepsilon_{0}$ places SafeStep derivation lengths within the scope of multiply recursive functions ($F_{\omega}$ in the fast-growing hierarchy [29]), well below the $\varepsilon_{0}$-calibrated systems arising from Hydra battles or Goodstein sequences. This remark is broader and proof-theoretic in character: unlike Proposition 5.5, it does not constitute a tight derivational-complexity bound. Extracting concrete complexity classes from ordinal embeddings requires an explicit complexity-extraction argument [21]. But it confirms that the proof-theoretic strength of the SafeStep termination argument is modest relative to the full ordinal hierarchy.
Proposition 5.14 (Single-exponential derivation-length bound for SafeStepCtx).
For any term $t$ of structural size $n$, any SafeStepCtx reduction chain $t\to^{m}u$ satisfies
$m+1\;\leq\;2^{2n}.$
[…]
Proposition 5.15 (Explicit exponential lower family for SafeStepCtx).
Let
$t_{0}:=\mathrm{merge}\;\mathrm{void}\;\mathrm{void},\qquad t_{n+1}:=\mathrm{rec}\Delta\;\mathrm{void}\;t_{n}\;(\delta(\delta\,\mathrm{void})).$
Then there exist terms $u_{n}$ and exact contextual lengths $m_{n}$ such that
$\texttt{SafeStepCtxPow}\;m_{n}\;t_{n}\;u_{n},\qquad 2^{n}\leq m_{n},$
$\mathrm{termSize}(t_{n})=5n+3,\qquad\mathrm{termSize}(u_{n})=2^{n+2}-3.$
Hence the guarded context-closed derivational complexity is genuinely single-exponential in structural size, up to constant factors in the exponent, and the same family exhibits explicit exponential normal-form size growth from linear-size guarded inputs.
The Lean proof is in Meta/SafeStepCtx_Complexity_LowerBound.lean; the relevant identifiers are ctxLowerFamily, ctxLowerNF, ctxLowerLen, termSize_ctxLowerFamily, termSize_ctxLowerNF, ctxLowerFamily_pow_nf, two_pow_le_ctxLowerLen, safeStepCtx_has_singleExponential_lower_family, safeStepCtx_has_singleExponential_size_family, and safeStepCtx_has_singleExponential_output_family. Together with Proposition 5.14, this pins the guarded contextual derivation-length story between matching single exponentials: the old ctxFuel/towerBound theorem remains available as a coarse explicit envelope, but the actual certified complexity class is now exponential rather than merely bounded above by a tower. The same family also provides a direct duplication-growth witness relevant to bounded-growth heuristics such as match-bound-style approaches, without claiming any generic match-bound metatheory.
Proposition 5.16 (Fast-growing upper envelope for SafeStepCtx).
Let $F_{0}(x)=x+1$ and $F_{k+1}(x)=F_{k}^{\,x+1}(x)$ be the finite fast-growing hierarchy on naturals. Define
$G(0)=6,\qquad G(k{+}1)=F_{k+2}\!\bigl(2\,G(k)+6\bigr).$
Then every SafeStepCtx reduction chain from a term $t$ has length at most $G(|t|)$, where $|t|$ is the structural size of $t$.
See Meta/SafeStep_Complexity_FastGrowing.lean (fastGrow, fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope, and safestep_length_bounded_by_fgOmegaEnvelope). This remains a coarse explicit corollary to Proposition 5.14. Proposition 5.15 shows that the actual guarded-context complexity is already exponential, so the fast-growing envelope should be read only as a retained explicit majorant. The paper-facing guarded-context hierarchy theorems are the two Cichon-style propositions below.
Proposition 5.17 (Cichon-style upper bound for SafeStepCtx).
Define the contextual notation
$\mathrm{mwCtxNote}(t)\;=\;\mathrm{lex3Note}\bigl(\mathrm{deltaFlag}(t),(\kappa^{M}(t),\texttt{ctxFuel}(t))\bigr)$
and the associated hierarchy bound
$\mathrm{mwCtxBound}(t)\;=\;\mathrm{cichon}\bigl(\mathrm{mwCtxNote}(t),0\bigr).$
Then for every SafeStepCtx reduction chain $t\to^{n}u$,
$n\;\leq\;\mathrm{mwCtxBound}(t).$
Moreover $\mathrm{mwCtxNote}(t)<\omega^{\omega}\!\cdot\!2$ for every $t$.
Proved in Meta/SafeStep_Complexity_MW_Ctx.lean (mwCtxNote, mwCtxBound, mwCtxNote_lt_opow_omega_mul_two, ctxFuel_le_mwCtxBound, safestepctx_length_le_mwCtxBound, ctxOrdinalObstruction_rootNote_increases, and not_all_safeStepCtx_rootNote_drop). This is a conservative Moser–Weiermann lift for the context-closed relation: it packages the already-certified contextual potential ctxFuel into the finite tail of the same lex3Note family used by the root calibration, yielding a genuine cichon theorem while keeping the bound in the calibrated $\omega^{\omega}\!\cdot\!2$ block. The same file also proves that the exported root-side calibrated note can increase on a concrete SafeStepCtx step, so Proposition 5.17 should be read as a theorem-backed conservative lift rather than as a direct contextual fundamental-sequence descent theorem.
Proposition 5.18 (Direct exact contextual extraction on a separate note package).
Define the exact contextual note
$\mathrm{ctxExactNote}(t)\;=\;\mathrm{lex3Note}\bigl(0,(0,\texttt{ctxFuel}(t))\bigr)$
and the associated hierarchy bound
$\mathrm{ctxExactBound}(t)\;=\;\mathrm{cichon}\bigl(\mathrm{ctxExactNote}(t),0\bigr).$
Then for every SafeStepCtx reduction chain $t\to^{n}u$,
$n\;\leq\;\mathrm{ctxExactBound}(t).$
Moreover $\mathrm{ctxExactNote}(t)<\omega$ for every $t$, and every contextual chain induces an exact fundamental-sequence descent on this new note family with total length $\texttt{ctxFuel}(t)-\texttt{ctxFuel}(u)$.
The Lean development is in Meta/SafeStep_Complexity_MW_CtxExact.lean (ctxExactNote, ctxExactBound, ctxExactNote_lt_omega, safeStepCtx_exact_step, safeStepCtxPow_exact_drop, and safestepctx_length_le_ctxExactBound). This gives the direct contextual exact-control theorem on a separate contextual note/control package. It does not reuse the root-calibrated note; that impossibility is recorded in Proposition 5.17. Instead, it turns the already-certified contextual potential into a standalone exact fundamental-sequence package.
6 External validation: TTT2 and CeTA
We submitted the full KO7 system (8 rules, no SafeStep guards) to TTT2 [18]. TTT2’s MAYBE results are tool search failures, not proofs of impossibility; the three YES results below independently corroborate the Lean theorem 5.4. Here the TPDB/TRS semantics is the usual context-closed rewriting semantics, so the CeTA certificate corresponds to Lean’s StepCtxFull, not merely the root relation Step.
Result.
TTT2 proves termination of the context-closed TRS using the DP framework:
  1. 1.
    Dependency pairs extracted: $\mathrm{rec}\Delta^{\sharp}(b,s,\delta(n))\to\mathrm{rec}\Delta^{\sharp}(b,s,n)$, plus two non-cyclic eqW pairs.
  2. 2.
    SCC decomposition: singleton $\{\mathrm{rec}\Delta^{\sharp}\}$.
  3. 3.
    Subterm criterion with $\pi(\mathrm{rec}\Delta^{\sharp})=3$: third argument decreases from $\delta(n)$ to $n$.
Table 9: TTT2 strategy results. CeTA column: independent certification by CeTA 2.36 [25].
Strategy Method class Result Time CeTA
FAST DP + SCC + subterm criterion YES 0.06s certified
COMP DP + TDG + SCC + subterm + matrix + LPO YES 0.70s certified
LPO Lexicographic path order YES 0.02s certified
POLY Polynomial interpretation MAYBE 0.15s rejected
KBO Knuth-Bendix order MAYBE 0.02s rejected
MAT(2) Matrix interpretation dim 2 MAYBE 0.31s rejected
MAT(3) Matrix interpretation dim 3 MAYBE 0.37s rejected
FBI Forward/backward instantiation MAYBE 0.17s n/a
The three YES results use modular or path-order methods; the five MAYBE results use direct/global methods (polynomial, KBO, matrix, forward/backward instantiation). The barrier theorems (Section 4) cover eight scalar numeric direct classes, additive, transparent-compositional, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, and max-plus, together with one symbolic variable-condition barrier, an explicit KBO corollary, and four vector / pair extensions, tracked fixed-dimension componentwise, tracked lexicographic dimension-$2$, balanced mixed-coordinate dimension-$2$ componentwise, and weighted scalar-projection componentwise. The KBO MAYBE outcome is explained directly by Corollary 4.16. The POLY MAYBE outcome is theorem-covered for the formalized direct polynomial families: affine, restricted quadratic, bounded cross-term, bounded multilinear, generalized bounded polynomial, and max-plus are all blocked, while the successful nonlinear witness $W$ lies outside those direct classes by cross-variable coupling and non-transparency. The MAT(d)(d) MAYBE outcomes now split cleanly. Fixed-row tracked-component and row-sum weighted-projection interpretations are theorem-covered by the existing matrix barriers and the explicit corollaries in Meta/MatrixProjectionCoverage.lean; only arbitrary mixed matrix interpretations remain outside the present theorem stack. The FBI MAYBE entry remains outside the theorem layer here: no general forward/backward-instantiation impossibility result is claimed. CPF certificates and CeTA logs are in Artifacts/ttt2/. The Lean module Meta/TPDB_Export.lean renders the exact KO7 TTT2 input and proves that the generated TPDB text matches the artifact text embedded in that module, which in turn corresponds to Artifacts/ttt2/KO7_full_step.trs. The artifact also includes a runnable verifier, lakeexeverifyTpdbExport, that checks the embedded text against the on-disk Artifacts/ttt2/KO7_full_step.trs file. In addition, Meta/TTT2_CertificateReplay.lean is a narrow Lean-side replay of the mathematical core of the FAST certificate: rather than parsing the external CPF, it re-establishes under Lean definitions the single recursive SCC, the projected argument (tool index $2$, paper index $3$), and the resulting well-founded DP pair proof.
Independent certification.
All three YES strategies received a certified verdict from CeTA 2.36. The four MAYBE strategies were rejected in the narrow certification sense that no checkable termination certificate was produced for CeTA to replay; this is not a proof of impossibility. These external tool runs corroborate the method separation predicted by the barrier theorems. They are not, however, part of the formal impossibility proof: the MAYBE outcomes are search failures of bounded heuristic strategies, not theoretical impossibility results.
7 Formalization structure
The Lean 4 artifact separates the kernel definitions, the barrier at the schema, the KO7-specific results, and the certification chain: Lean 4’s Metatheory library [23] provides abstract rewriting, confluence, and strong-normalization infrastructure, but not a library of TRS termination methods, impossibility barriers, or certification bridges. Relative to the established Isabelle/Coq rewriting ecosystems (IsaFoR/CeTA, CoLoR), this project is not another checker for externally generated positive certificates. It is, to our knowledge, the first large Lean 4 development centered on TRS termination methods, fixed-system impossibility boundaries, concrete negative witnesses, and a bridge back to the standard TRS toolchain.
  • Schema barrier layer: additive / transparent-compositional impossibility, affine and nonlinear scalar barriers, matrix-side projection barriers, the scalar-projection meta-theorem, the symbolic variable-condition barrier, and strengthened pumped-subclass corollaries.
  • Operational boundary tools: computable barrier witnesses, including extended extractors for quadratic, max-plus, and projected matrix-side subclasses, theorem-backed canonical witness budgets for those extractors, a small synthesis-oracle wrapper around them, and a decidable coefficient-table classifier for the formalized scalar and tracked-primary pair families.
  • Growth and transformed-relation infrastructure: reusable successor- and wrapper-growth bridges, a small reusable DP fragment, and the KO7 extracted DP escape.
[…]
  • Certification chain: per-rule decreases and strong normalization for SafeStep, totality and soundness of normalizeSafe, root-step confluence via Newman, and reachability decidability.
  • Complexity and ordinal layers: guarded-context ctxFuel termination, a new single-exponential contextual derivation-length theorem with matching explicit lower family, exact merge-spine normalizer cost, and ordinal upper/lower calibration.
  • Catalog, SCC, and ablation layers: theorem-backed depth and pure-precedence families, verified excluded-method results, the delayed-duplication SCC theorem together with its finite cyclic, raw-graph, automatic closed-path / round-trip extraction, and bounded finite-round-trip reachability extensions, the multiplicity-preserving SCC theorem together with its finite cyclic, raw-graph, automatic closed-path / round-trip extraction, and bounded finite-round-trip reachability extensions, the linear-recursion ablation, and the object-level collapse surrogate.
  • Artifact bridge: Lean-side TPDB export for the full KO7 root TRS, an exact-match theorem for the checked TTT2 artifact text, a narrow internal replay of the FAST DP/subterm certificate core, and a runnable verifier for the on-disk Artifacts/ttt2/KO7_full_step.trs file.
Table 10: Formalization layers, primary declarations, and evidence types in the active artifact.
Layer Primary declarations Evidence Scope / side conditions
SafeStep wf_SafeStepRev_c, confluentSafe, normalizeSafe_total, reachability_decidable Internal Lean proof Guarded root relation only
SafeStepCtx wf_SafeStepCtxRev, confluentSafeCtx, safeStepCtx_length_le_contextualExpBound, safestep_length_bounded_by_fgOmegaEnvelope Internal Lean proof Partial context closure of the guarded system; single-exponential structural bound with matching explicit lower family plus coarser inherited envelopes
Step wf_StepRev_poly, wf_StepRev_mpo Internal Lean proof Full unguarded root relation; separate from the triple-lex certification layer
StepCtxFull wf_StepCtxFullRev_poly, stepCtxFullPow_length_le_W Internal Lean proof Full context closure, derived from the nonlinear witness $W$
TPDB / FAST bridge ko7_full_step_tpdb_matches_artifact_text, ko7FastReplay_sound Lean bridge + external certificate trail External CeTA verdicts corroborate the StepCtxFull result but are not impossibility proofs
The formalization uses three proof shapes: (i) direct full-step orientation (polynomial $W$, KO7-specialized MPO), stated on the unrestricted Step relation; (ii) guarded-fragment certification (triple-lex measure, confluence, normalizer), on SafeStep; and (iii) transformed-relation arguments (dependency pairs), factored through a small reusable fragment for projection-rank descent and SCC-path reasoning. The reusable schema barrier layer is also exposed through a dedicated public API module, OperatorKO7/SchemaAPI.lean, so downstream developments can import the generic barrier stack and witness/classifier tooling without pulling in the full KO7 certification chain.
The support layer is computational as well. Meta/BarrierClass_Classifier.lean classifies coefficient tables into the formalized barrier families. Meta/BarrierWitness.lean and Meta/BarrierWitness_Extended.lean return certified violating triples across the scalar and projected matrix-side subclasses. Meta/BarrierWitness_Budgets.lean packages the corresponding canonical construction budgets, and Meta/SynthesisOracle.lean wraps the witness layer in a small oracle interface for direct-orientation attempts. All declarations are sorry-free.
The termination, normalization, and barrier developments are computationally effective. The ordinal and hierarchy-calibration layer (Meta/DM_OrderType.lean, Meta/DM_OrderType_LowerBound.lean, Meta/SafeStep_Complexity_MW_Root.lean, Meta/SafeStep_Complexity_MW_Ctx.lean, and Meta/MPO_ProofTheoreticBound.lean) uses standard classical infrastructure from Lean 4 / Mathlib (Classical.choice, propext) for noncomputable ordinal reasoning. For reviewability, the repository ships a bidirectional proof-dependency graph in Docs/KO7_BLUEPRINT.md and Docs/ko7_blueprint.json, mapping paper theorem labels to primary Lean declarations and back.
From a proof perspective, the development is not a Lean transcription of the paper statements. Across more than sixty Lean/build files, the main theorem layer consists of impossibility theorems, computed barrier witnesses, boundary classifiers, and a small internal replay of the FAST certificate core, all inside one Lean library rather than as reconstructions of externally generated positive certificates. Mathlib supplies the well-foundedness background for Dershowitz–Manna multisets and the ambient ordinal infrastructure used in Meta/DM_OrderType.lean, but the precise $\omega^{\omega}$ image and lower-bound argument in Meta/DM_OrderType_LowerBound.lean and the full context-closed lift of the nonlinear witness through Meta/ContextClosed_SN_Full.lean are built in this repository.
The repository also includes an upstream-prep staging surface for the general DM package in Meta/DM_UpstreamSurface.lean. This is a preparation artifact for a future Mathlib contribution, not a claimed upstream merge. Those two modules are the clearest mechanization bottlenecks: the former because it must recover a sharp ordinal calibration rather than only well-foundedness, and the latter because the polynomial witness has to be threaded through every constructor context together with an explicit derivation-length control.
The paper does not prove arbitrary-target reachability decidability, an exact order type for the full triple-lex image of guarded reductions, or generic MPO/LPO metatheory over arbitrary signatures. It also does not claim that the computable triple-lex certificate handles unrestricted Step, nor that Meta/KBO_Impossible.lean formalizes generic KBO metatheory; the full root Step arguments are instead the nonlinear witness and KO7-specialized MPO proofs, while the KBO corollary is extracted from the symbolic variable-condition abstraction. Likewise, Meta/WPO_PolynomialBarrier.lean is not generic WPO metatheory: it only lifts the bounded direct-polynomial barrier to direct orders whose strict comparison is certified by that same algebraic branch. Nor does it lift the tree-based barrier stack wholesale to higher-order rewriting; the current artifact instead isolates a sharing-aware surrogate in Meta/SharingBarrierLift.lean showing exactly where the no-sharing assumption is load-bearing. Each of those boundaries is explicit in the corresponding sections below.
For reproducibility, the active artifact is pinned to the following Lean / Mathlib environment:
Lean toolchain: leanprover/lean4:v4.22.0-rc4
Mathlib commit: 632465e4b02cb70a5dfa4cfe15468e8a62c2bd85
The file lake-manifest.json freezes the transitive dependency graph. Public artifact-facing notes and staged outputs are collected under the single repository root Artifacts/. The repository-level note Artifacts/REPRODUCIBILITY.md documents the minimal replay path. Artifacts/MICRO_BENCHMARKS.md records representative local build/check times together with the archived TTT2 run times. The script scripts/make_referee_bundle.py writes a referee-facing bundle in the public Artifacts/ tree, containing the frozen Lean snapshot, generated documentation, and the archived Artifacts/ttt2/ validation trail together with replay instructions. Separately, scripts/stage_tpdb_submission.py stages, but does not perform, a future TPDB submission in that same Artifacts/ tree.
Trusted-base note. Beyond ordinary elaboration, the intended replay path is lake build OperatorKO7, lake exe verifyTpdbExport, and #print axioms on headline declarations after import OperatorKO7. In the pinned environment, the following spot-check commands return only the standard Lean / Mathlib axioms propext, Classical.choice, and Quot.sound:
OperatorKO7.MetaCM.wf_SafeStepRev_c
OperatorKO7.PolyInterpretation.wf_StepRev_poly
OperatorKO7.MetaDM.dm_order_type_omega_omega
MetaSN_KO7.safeStepCtx_length_le_contextualExpBound
The active tree currently uses native_decide 19 times: 15 concrete finite sanity checks in the dependency-pair extraction infrastructure, 3 small controlled hierarchy checks in Meta/OrdinalHierarchy_Controlled.lean, and 1 literal TPDB-text equality check in Meta/TPDB_Export.lean. No generic CPF parser, external solver oracle, or user-declared axiom layer is part of the main theorem surface; the only noncomputable proof layer is the explicitly identified classical ordinal calibration infrastructure above.
The active artifact (dual-licensed) is at https://github.com/MosesRahnama/OperatorKO7.
8 Related work
Termination of duplicating systems.
The multiset ordering of Dershowitz and Manna [9] handles rules replacing one term by multiple smaller terms. Recursive path orderings [11] extend this to simplification orders. Duplication and termination are surveyed in [26, 4]. Zantema [31] established a hierarchy between termination method classes. Endrullis, Waldmann, and Zantema [12] proved that specific TRSs cannot be handled by matrix interpretations alone, the closest prior art to a formal impossibility result for a specific method.
Interpretation frameworks.
Yamada’s tuple interpretations [30] give a modern umbrella framework subsuming polynomial, matrix, arctic, and related interpretation methods. This is the right taxonomic backdrop for the present barrier stack. Our twelve blocked direct classes are best read as a certified boundary for direct scalar and tracked-tuple fragments inside that broader interpretation framework, while the successful escape mechanisms in this paper either move to transformed relations (dependency pairs) or to global witnesses that are not members of the blocked direct fragments.
Dependency pairs and modular termination.
The dependency pair method [2] was introduced to overcome limitations of direct simplification orders. Hirokawa and Middeldorp [15] introduced the subterm criterion. Thiemann and Giesl showed size-change termination for TRSs captures only simple termination, while DP goes beyond [27]. IsaFoR/CeTA [25] provides machine-checked formalizations of termination methods in Isabelle; CoLoR [5] and Coccinelle/CiME [8] in Coq. These projects certify positive termination proofs or their soundness conditions. Our artifact occupies a different proof-engineering niche: it internalizes impossibility theorems for direct classes on a fixed terminating system, adds executable witness extraction and classification around those theorems, and only then reconnects to the external TRS toolchain through certified artifacts.
Table 11: Position of the present artifact relative to nearby formal developments.
Artifact Assistant Primary focus Fixed-system impossibility theorems? External certificate checking?
This artifact Lean 4 Fixed-system barrier theorems, escape classification, certified witnesses, and KO7 certification Yes Bridge / replay layer only
AFP undecidability formalization [28] Isabelle/HOL Undecidability of orientation problems across all inputs No No
IsaFoR/CeTA [25] Isabelle/HOL Certification and checking of positive TRS proofs No Yes
CoLoR / Coccinelle-CiME [5, 8] Coq Certified checking / soundness for positive rewriting arguments No Yes
Lean Metatheory [23] Lean 4 Generic rewriting, confluence, and normalization infrastructure No No
Path orders and proof-theoretic strength.
Ferreira and Zantema [13] showed LPO well-foundedness depends on the subterm property. Buchholz [6] and Dershowitz-Okada [10] analyzed path orders as requiring Kruskal’s Tree Theorem. Rathjen and Weiermann [24] calibrated the resulting general path-order metatheory at $\Pi^{1}_{1}$-CA0. This concerns well-foundedness results for path orders over arbitrary signatures and trees. It does not imply that every fixed-signature specialization must internally reproduce that strength: our Lean MPO result for KO7 is a direct ordinal-ranking proof for the fixed KO7 signature, precedence, and same-head clause. The DP subterm criterion [15] does not require Kruskal-strength principles; its soundness is formalized in IsaFoR at strength below PA.
Confluence and mechanization platforms.
Newman’s lemma [22] has been mechanized in Isabelle/HOL [25] and Lean 4 [23]. Our application is to a root ARS with a discharged local-join hypothesis. The TRS certification ecosystem (IsaFoR/CeTA in Isabelle, CoLoR in Coq, the Ramos library [23] in Lean 4) certifies positive termination proofs but not impossibility theorems for measure classes. IsaFoR also certifies nontermination via loop witnesses [25]; Kim, Saito, Thiemann, and Yamada [17] formalized co-rewrite pairs for non-reachability. Both certify properties of specific systems rather than statements about measure classes. Our contribution is therefore not another certificate checker or soundness layer, but a theorem-level boundary analysis with computed outputs that still connect back to the standard TRS toolchain. The distinction from [20, 28] is noted in §1.
Proof-assistant termination checkers.
The same obstruction is also relevant to proof assistants, even though the artifact does not formalize Lean’s termination_by checker, Agda’s Foetus checker, or Coq’s guard condition as theorem objects. The textbook duplicating rule $f(x,s(y))\to g(x,f(x,y))$, now packaged in Meta/TextbookDupInstance.lean, gives a concrete recursive pattern whose direct additive, transparent-compositional, and affine orientation already fails by theorem. So any termination checker restricted to direct structural or lexicographic aggregation of recursive arguments needs an escape mechanism analogous to argument projection or a stronger global witness: otherwise the duplicated payload remains on the wrong side of the boundary for exactly the same reason it defeats the direct measure classes formalized here.
9 Conclusion
We formalized an orientation boundary for first-order rewrite systems with step-duplicating recursors. Twelve direct measure classes in the barrier package are excluded by schema-level impossibility theorems (Theorems 4.1 to 4.14); the escape side is classified by the DP clause, the explicit DP base-order boundary, transparency essentiality, the generalized bounded-polynomial escape condition, the escape trichotomy, and the two SCC synchronization theorems; and the concrete witness calculus KO7 carries a guarded certification chain together with full-step polynomial and MPO witnesses.
The barrier stack separates baseline additive/compositional failure, widened scalar families, vector / pair projection layers, and the symbolic variable-condition side; the escape results track that organization closely. The paper therefore marks a boundary for explicit direct families, not a claim that KO7 defeats termination methods wholesale. Dependency pairs, LPO, the nonlinear witness, and the KO7-specialized MPO remain on the successful side of that line, and the repository states those imports at the theorem level.
The repository is a pinned Lean 4 snapshot for the explicit first-order setting treated here. It contains the schema-level impossibility theorems, the escape classification, the guarded certification chain, the full-step termination witnesses, and the external TRS validation trail.
Appendix A Addendum
A.1 Lean kernel location
The mechanized kernel is OperatorKO7/Kernel.lean at https://github.com/MosesRahnama/OperatorKO7. Under OperatorKO7/Meta/ lie the safe fragment and its proofs.
Path convention: manuscript references of the form Meta/... name Lean module paths in the OperatorKO7.Meta.* namespace; the corresponding repository-root files live under OperatorKO7/Meta/...lean. Literal filesystem paths are written repo-root-relative.
A.2 Module map
Table LABEL:tab:module-map lists the theorem-bearing kernel and meta modules in the active artifact. Lightweight test files under OperatorKO7/Test/ are intentionally omitted.
Table 12: Lean module map. Key theorem identifiers are shown in parentheses.
Module Contents
Kernel.lean Trace (7 constructors), Step (8 rules), StepStar
Meta/SafeStep_Core weight, kappaM, deltaFlag, guarded SafeStep
Meta/ComputableMeasure tau, Lex3c, per-rule decrease (wf_SafeStepRev_c)
Meta/ComputableMeasure_Verification Sanity checks (deltaFlag_binary, no_infinite_safestep_chain)
Meta/Normalize_Safe Certified normalizer; totality, soundness (normalizeSafe_sound)
Meta/NormalizeSafe_LowerBound Auxiliary step counter for the certified normalizer; exact merge-spine cost family (normalizeSafeSteps_mergeVoidChain)
Meta/SafeStep_Ctx Context closure SafeStepCtx
Meta/ContextClosed_SN ctxFuel measure for the guarded partial context closure (wf_SafeStepCtxRev)
Meta/ContextClosed_SN_Full Full unguarded context closure; polynomial witness lifted through all constructor positions (wf_StepCtxFullRev_poly)
Meta/ContextClosedBarrier Root-to-context barrier survival corollaries for StepCtxFull (stepCtxFull_orientation_implies_root, representative entries no_stepCtxFull_orientation_affine_with_pump,
no_stepCtxFull_orientation_matrixFunctional_with_projected_affine_pump; see the module for the complete no_stepCtxFull_orientation_* family list)
Meta/SafeStep_Complexity Exact-length contextual derivations; global $W$-based derivation-length bound (stepCtxFullPow_length_le_W)
Meta/SafeStep_Complexity_Ordinal Legacy ctxFuel-based SafeStepCtx derivation-length bound; size-indexed tower-exponential corollary (safestep_length_bounded_by_size)
Meta/SafeStep_Complexity_FastGrowing Explicit finite fast-growing envelope majorizing the same coarse SafeStepCtx tower bound (fgOmegaEnvelope, safestep_length_bounded_by_fgOmegaEnvelope)
Meta/ContextualCopyBudget Positive position-aware contextual copy-budget analysis and strict descent proof for ctxDupPotential (ctxDupPotential_decreases_ctx, safeStepCtx_length_le_two_pow_double_termSize)
Meta/ContextualCopyBudget_NoGo Auxiliary contextual coordinates, obstruction witnesses, and the no-go theorem for the monotone arithmetic candidate class (coordExpr_payloadFamily_nondec, not_coordExprStackCertifies_payloadFamily)
Meta/SafeStepCtx_Complexity_Exponential Public single-exponential SafeStepCtx envelope refining the old tower bound (contextualExpBound, contextualExpBound_le_complexity_bound, safeStepCtx_length_le_contextualExpBound)
Meta/SafeStepCtx_Complexity_Cichon Finite-note cichon wrapper around the sharpened contextual exponential theorem (ctxExpNote, ctxExpCichonBound, safeStepCtx_length_le_ctxExpCichonBound)
Meta/SafeStepCtx_Complexity_LowerBound Explicit contextual lower family with exact lengths, linear source size, and exponential normal-form output size (ctxLowerFamily_pow_nf, safeStepCtx_has_singleExponential_lower_family,
safeStepCtx_has_singleExponential_output_family)
Meta/SafeStep_Complexity_MW_Root Notation bridge from the calibrated root measure to NONote, together with the genuine root cichon theorem (mwRootNote, safeStepPow_length_le_mwRootBound)
Meta/SafeStep_Complexity_MW_Ctx Conservative contextual cichon lift plus explicit obstruction to direct contextual reuse of the root note (safestepctx_length_le_mwCtxBound, not_all_safeStepCtx_rootNote_drop)
Meta/SafeStep_Complexity_MW_CtxExact Direct contextual exact controlled-descent package on a separate finite note family (ctxExactNote, safestepctx_length_le_ctxExactBound)
Meta/SafeRoot_Complexity Exact-path realization and explicit linear / tower / root-cichon envelopes for the certified root normalizer (normalizeSafeSteps_realized,
normalizeSafeSteps_le_linear_termSize, normalizeSafeSteps_le_complexity_bound, normalizeSafeSteps_le_mwRootBound)
Meta/Reachability_Complexity Explicit decision-cost envelope for safe-normal-form reachability, including a linear size upper bound and matching linear families (reachabilityDecisionCost_le_linear_termSize,
reachabilityDecision_has_linear_size_family)
Meta/SafeStepCtx_Confluence Discharged global contextual local-join obligation and unconditional SafeStepCtx confluence, retaining the Newman characterization as a corollary (localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll)
Meta/EqGuardedConfluence Intermediate root fragment keeping all root rules except the reflexive eqW-diff overlap; root confluence restored (confluentEqGuarded)
Meta/Confluence_Safe Local-join lemmas (localJoin_all_safe)
Meta/EqW_Guard_Barrier Full-step eqW overlap obstruction; guard-necessity theorem (not_localJoinStep_eqW_refl)
Meta/Newman_Safe Newman instantiation; confluentSafe; unique NFs; reachability_decidable
Meta/Impossibility_Lemmas Failed-measure catalog; historical dead-end notes moved to Docs/Impossibility_DeadEnds.md
Meta/Conjecture_Boundary No-go theorems, GlobalOrients framework
Meta/StepDuplicatingSchema Parametric schema; generic impossibility; affine/linear barrier; projection-rank escape (Thms. 4.14.2;
no_additive_orients_dup_step)
Meta/BarrierWitness Computable certificate extractors: additive_witness, compositional_witness, affine_witness
Meta/BarrierWitness_Extended Extended witness extractors for pumped restricted-quadratic, max-plus, and projected matrix-side families (quadratic_with_pump_witness,
max_with_pump_witness, matrixFunctional_with_projected_affine_pump_witness)
Meta/BarrierWitness_Budgets Canonical witness-budget packaging for the extractor layer (WitnessStepShape, affine_with_pump_witness_shape,
quadratic_with_pump_witness_shape, max_with_pump_witness_shape)
Meta/AffineThresholdSharpness Canonical affine family showing exactness of the generic affine pump bound (affineThresholdMeasure_bound,
affineThresholdMeasure_exact_cutoff)
Meta/SynthesisOracle Small synthesis-oracle wrapper around the barrier certificates, with a local refinement-change invariant (SynthesisOracle, successful_refinement_changes_certificate)
Meta/TTT2_CertificateReplay Narrow Lean-side replay of the FAST certificate core: SCC count, projected argument, and resulting DP-pair well-foundedness (ko7FastReplay_sound, wf_DPPairRev_replayed)
Meta/SharingBarrierLift Sharing-aware surrogate showing exactly where the tree-style direct barrier stops applying (sharedCounter_orients_step, sharing_breaks_tree_barrier)
Meta/TPDB_Export Lean-side TPDB export for the full KO7 root TRS; exporter output equals the embedded artifact text literal (ko7_full_step_tpdb_matches_artifact_text)
Meta/TextbookDupInstance Second named schema instance for the textbook rule $f(x,s(y))\to g(x,f(x,y))$, with direct additive/compositional/affine barrier corollaries and witness aliases (no_global_textbook_orientation_additive, textbook_additive_witness)
Meta/CompositionalMeasure_Impossibility KO7 instantiation; KO7 DP witness; KO7 affine barrier (Thms. 4.194.22;
no_additive_compositional_orients_rec_succ)
Meta/QuadraticBarrier Restricted degree-$2$ barrier without step-counter coupling; KO7 specializations (no_quadratic_counter_orients_dup_step_of_unbounded, no_global_step_orientation_quadratic_of_succ_pump)
Meta/QuadraticCrossTermBarrier Bounded step-counter cross-term quadratic barrier; KO7 specializations under the wrapper-dominance side condition (CrossTermBoundedAtBase,
no_cross_quadratic_orients_dup_step_of_unbounded,
no_global_step_orientation_cross_quadratic_of_succ_pump)
Meta/MultilinearBarrier Finite multilinear direct-polynomial barrier with frozen-coefficient dominance; successor- and wrap-pump corollaries (MultilinearDominatedAtBase, no_multilinear_orients_dup_step_of_unbounded)
Meta/PolynomialBarrierGeneral Generalized bounded direct-polynomial barrier with repeated variables and a necessary frozen base-dominance failure condition (no_polynomial_orients_dup_step_of_unbounded,
polynomial_escape_requires_failure_of_base_dominance)
Meta/WPO_PolynomialBarrier WPO-facing direct polynomial-branch corollary lifting the bounded-polynomial barrier to direct orders sound for that algebra (no_wpoPolynomialDirect_orients_dup_step_of_unbounded,
no_global_step_orientation_wpoPolynomialDirect_of_succ_pump)
Meta/MaxBarrier Schema-level max-plus barrier with unbounded, successor-pump, and wrap-pump forms (no_max_orients_dup_step_of_unbounded, no_global_step_orientation_max_of_succ_pump)
Meta/ArcticBarrier Arctic-style primary-projection corollary of the max barrier (no_arctic_primary_orients_dup_step_of_unbounded,
no_global_step_orientation_arctic_primary_of_unbounded)
Meta/TropicalBarrier Broader tropical primary-projection barrier reducing finite tropical coordinates to the same scalar obstruction (no_tropical_primary_orients_dup_step_of_unbounded,
no_global_step_orientation_tropical_primary_of_unbounded)
Meta/PumpedBarrierClasses Strengthened pumped subclasses yielding unconditional affine / quadratic / bounded cross-term / generalized bounded-polynomial / multilinear / max-plus / projection-based matrix corollaries (PolynomialMeasureWithPump,
MultilinearMeasureWithPump, MaxMeasureWithPump,
MatrixFunctionalMeasureWithProjectedAffinePump)
Meta/StandardPumpLemmas Reusable successor-/wrapper-growth lemmas and subclass constructors bridging conditional barriers to pumped subclasses (affine_hasUnboundedRange_of_succ_pump, QuadraticCounterMeasure.withWrapPump, MatrixMeasure2.withPrimarySuccPump)
Meta/EscapeTrichotomy Escape trichotomy for the explicit KO7 direct universe, plus the projection-based extension for weighted functional and balanced mixed-coordinate families (ko7_nat_direct_escape_trichotomy, ko7_direct_escape_trichotomy_extended,
ko7_projection_escape_trichotomy)
Meta/DepthBarrier KO7-specific max-aggregative depth family; dedicated depth barrier theorem (MaxDepthMeasure, no_global_step_orientation_maxDepth)
Meta/PrecedenceBarrier KO7-specific pure head-precedence family; dedicated precedence barrier theorem (HeadPrecedenceFamily, no_global_step_orientation_headPrecedenceFamily)
Meta/ObjectAxiom_Ablation Directed wrapper collapse rule $\mathrm{app}(t,t)\to t$; collapse sanity and preserved direct barriers (wf_StepCollapseRev, no_global_stepCollapse_orientation_additive)
Meta/MutualDuplication_General Bounded two-node SCC theorem; delayed duplication with additive and affine composite-step barriers (alternating_dup2_realized, no_affine_orients_alternating_dup2_composite_of_unbounded)
Meta/MutualDuplication_CycleFlow Abstract delayed-duplication one-cycle metatheorems, packaging additive, affine, transparent-compositional, and scalar-projection contextual barriers from a certified cycle witness (no_global_orients_ctx_additive, no_global_orients_ctx_of_scalar_projection_affine_of_unbounded)
Meta/MutualDuplication_KNode Finite cyclic ‘k+1‘-node delayed-duplication generalization with exact cycle realization and additive / affine minimal-context barriers (cycle_realized, no_global_orients_ctx_affine_of_unbounded)
Meta/MutualDuplication_KNode_Abstract Bridge module showing that the finite delayed-duplication ‘k+1‘-node SCC factors through the abstract delayed-cycle layer, including transparent-compositional and projected matrix-style corollaries (no_global_orients_ctx_compositional_transparent_via_cycleFlow, no_global_orients_ctx_matrixFunctional_of_projected_unbounded_via_cycleFlow)
Meta/GraphPathExtraction Generic extraction of explicit edge paths from nonempty transitive-closure proofs, used to turn closed-path and round-trip SCC witnesses into concrete cycle data automatically (EdgePath, ofTransGen, ofRoundTrip)
Meta/FiniteGraphReachability Bounded successor-closure reachability for finite decidable directed graphs, proving that Fintype.card rounds capture reflexive transitive closure and supporting finite-round-trip SCC wrappers without explicit path proofs (Reachable, reachable_iff_reflTransGen, transGen_of_reachable_of_ne)
Meta/FiniteGraphSCC Finite-graph SCC witness packaging and search layer: define the finite pair-search surface and prove its equivalence with nontrivial SCC existence, then recover a concrete distinct pair together with both reachability directions internally (sccPairs, HasNontrivialSCC, witnessPair, reachable_witnessSrc_witnessDst)
Meta/MutualDuplication_GraphCycle Raw-graph delayed-duplication metatheorems for arbitrary node / edge systems equipped with a certified closed cycle witness, packaging additive, affine, transparent-compositional, and scalar-projection contextual barriers (CyclePath, no_global_orients_ctx_of_scalar_projection_affine_of_unbounded)
Meta/MutualDuplication_Case Concrete alternating worked instance of the bounded SCC theorem (alternating_dup2_realized, no_additive_orients_alternating_dup2_step)
Meta/MutualDuplication_Preserving Multiplicity-preserving synchronized SCC barrier; exact payload-count preservation with additive and conditional affine synchronized-exposure obstructions (step_preserves_payloadCount, no_additive_orients_synchronized_cycle, no_affine_orients_synchronized_cycle_of_wrapper_dominance)
Meta/MutualDuplication_PayloadFlow Abstract payload-flow metatheorems for one-cycle synchronized exposure, packaging additive and affine global contextual barriers independently of the concrete packet syntax (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_wrapper_dominance)
Meta/MutualDuplication_Preserving_KNode Finite cyclic synchronized-packet preserving generalization with exact payload preservation, one-cycle exposure, additive failure, and conditional affine wrapper-dominance failure (synchronized_cycle_realized, no_global_orients_ctx_affine_of_wrapper_dominance)
Meta/MutualDuplication_Preserving_Abstract Bridge module showing that the finite synchronized-packet SCC factors through the abstract payload-flow layer, including additive, affine, packet-transparent, and projected packet-transparent corollaries (no_global_orients_ctx_additive_via_payloadFlow, no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent_via_payloadFlow)
Meta/MutualDuplication_Preserving_Transparent Packet-transparent compositional and scalar-projection extension of the finite synchronized-packet preserving SCC barrier (no_global_orients_ctx_packetTransparent, no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent)
Meta/MutualDuplication_PacketGraph Raw-graph synchronized-packet metatheorems for arbitrary node / edge systems equipped with a certified closed cycle witness, together with closed-path / round-trip extraction wrappers, packaging additive, affine, transparent, and scalar-projection contextual barriers (CyclePath, no_global_orients_ctx_of_scalar_projection_transparent_of_roundTrip)
Meta/MutualDuplication_RelationalGraph Relation-level construction layer building the delayed or preserving raw-graph SCC system automatically from a node type, edge relation, shared constructor interface, and local edge-realization theorem, then reexporting the round-trip, finite-round-trip, finite-SCC, and finite-search-result contextual barrier wrappers (no_global_orients_ctx_additive_of_hasNontrivialSCC, search-result wrappers for successful finite-SCC discovery)
Meta/MutualDuplication_CallGraph Call-graph construction layer building the delayed or preserving SCC system automatically from finite extracted ‘nodeKey‘ / ‘succKeys‘ data on a node type, then reexporting the same finite-search-result contextual barrier wrappers (no_global_orients_ctx_additive_of_hasNontrivialSCC, call-graph wrappers for successful finite-SCC discovery)
Meta/MutualDuplication_ExtractedCallGraph Array-backed extracted-data construction layer building the delayed or preserving call graph and SCC barrier interface automatically from raw extracted node data indexed by Finnodes.size (finite SCC search on the extracted graph, then additive / affine / transparent wrapper export)
Meta/MutualDuplication_Transparent Transparent-compositional and projected matrix-style extension of the alternating SCC barrier (no_global_orients_ctx_compositional_transparent, no_global_orients_ctx_matrixFunctional_of_projected_unbounded)
Meta/MatrixBarrierD Fixed-dimension tracked componentwise vector barrier; KO7 specialization for arbitrary finite dimension (no_matrixD_orients_dup_step_of_componentwise_pump, no_global_step_orientation_matrixD_of_componentwise_pump)
Meta/MatrixBarrier2 Dimension-22 worked instance of the tracked componentwise barrier, plus symmetric second-component lemmas (no_matrix2_orients_dup_step_of_componentwise_pump, no_matrix2_orients_dup_step_of_componentwise_pump_snd)
Meta/MatrixBarrierLex Dimension-22 lexicographic affine pair barrier; tracked primary-component KO7 specialization (no_matrix2_lex_orients_dup_step_of_unbounded_primary, no_global_step_orientation_matrix2_lex_of_componentwise_pump)
Meta/MatrixBarrierMix2 Balanced mixed-coordinate dimension-$2$ pair barrier via aggregate-sum projection (no_matrixMix2_orients_dup_step_of_sum_pump, no_global_step_orientation_matrixMix2_of_sum_pump)
Meta/MatrixBarrierFunctional Weighted scalar-projection componentwise barrier; tracked-coordinate and aggregate-sum projection unification (no_matrixFunctional_orients_dup_step_of_componentwise_pump,
no_global_step_orientation_matrixFunctional_of_componentwise_pump)
Meta/MatrixProjectionCoverage Explicit fixed-row and row-sum corollaries exposing matrix-side projection coverage for tool-facing discussion (no_global_step_orientation_matrix_fixed_row_of_componentwise_pump,
no_global_step_orientation_matrix_row_sum_of_componentwise_pump)
Meta/ScalarProjectionBarrier Generic scalar-projection lift subsuming several projection-based matrix barriers (no_orients_dup_step_of_scalar_projection,
no_matrixMix2_orients_dup_step_of_sum_pump_via_projection)
Meta/SymbolicComparatorBarrier Symbolic variable-condition obstruction for direct KBO-style comparators on the duplicating rule schema (not_orients_dup_rule,
no_symbolic_variable_condition_orients_dup_step)
Meta/KBO_Impossible Paper-facing renaming layer over the symbolic variable-condition barrier in Meta/SymbolicComparatorBarrier.lean: re-exports the abstract obstruction under KBO-facing names and adds one genuinely new concrete-trace bridge theorem for rec_succ. The substantive proof lives in the symbolic barrier file, not here (no_kbo_orients_ko7_rec_succ, no_kbo_orients_ko7_rec_succ_trace)
Meta/BarrierClass_Classifier Decidable coefficient-table classifier for the formalized scalar and tracked-primary pair barrier families (ScalarCoeffTable.structuralClasses, PrimaryPairCoeffTable.barrierClass)
Meta/TypedBarrierSurvival Simply-typed first-order recursor fragment; additive survival and affine survival under a typed step pump (no_additive_orients_typed_recSucc,
no_affine_orients_typed_recSucc_of_stepPump)
Meta/ManySortedBarrierSurvival Specialized many-sorted first-order repackaging of the typed recursor fragment; additive and affine barrier survival carry over unchanged (no_additive_orients_manySorted_recSucc,
no_affine_orients_manySorted_recSucc_of_stepPump)
Meta/DependencyPairs_Fragment Reusable DP fragment: rank-based pair problems and SCC-path descent (wfRev in namespace DPProjection, target_lt_source in namespace SCCCycle)
Meta/DependencyPairs_FiniteGraph Finite dependency-pair graph interface reusing the finite-SCC search layer and repackaging a discovered SCC as the fragment’s standard cycle witness (FiniteDPGraph, HasNontrivialSCC, toSCCCycle)
Meta/DependencyPairs_CallGraph Finite dependency-pair call-graph interface building the pair relation automatically from extracted ‘nodeKey‘ / ‘succKeys‘ data and reusing the same finite-SCC search / cycle packaging surface (FiniteCallGraph, HasNontrivialSCC, toSCCCycle)
Meta/DependencyPairs_ExtractedCallGraph Array-backed extracted dependency-pair call-graph interface generating the node type as Finnodes.size from raw extracted node data before reusing the same finite-SCC search / contradiction surface (FiniteExtractedCallGraph, ofArrayMap, HasNontrivialSCC)
Meta/DependencyPairs_TPDBExtraction Concrete TPDB-side dependency-pair extraction layer computing defined heads and right-hand-side call heads from a finite rule list, then building the array-backed extracted call graph automatically (tpdbDefinedHeads, extractTpdbNodes, tpdbExtractedCallGraph)
Meta/DependencyPairs_FirstOrderExtraction Generic first-order dependency-pair extraction layer over arbitrary symbols and variables, computing defined heads, extracted call-head nodes, and the array-backed call graph for any finite first-order TRS (FOTerm, FORule, foDefinedHeads, foExtractedCallGraph)
Meta/DependencyPairs_FirstOrderFrontend Rule-record frontend for the generic first-order extractor, taking a finite rule array plus lhs / rhs extractors directly to the extracted call graph and the same SCC-search / contradiction surface (extractRuleFrontendNodes, foExtractedCallGraphOf, HasNontrivialSCCOf)
Meta/DependencyPairs_FirstOrderEngine Consolidated engine-level frontend packaging an internal first-order TRS / DP engine as one object and reexporting the same extracted-call-graph and SCC-search surface; the same module now also contains the typeclass-based internal-engine view and the raw-term bridge for engines with their own syntax (FiniteFirstOrderEngine, HasFiniteFirstOrderView, HasFirstOrderTermView, HasFiniteRawFirstOrderView, extractedCallGraph)
Meta/DependencyPairs_HeadView Minimal head/call-head bridge for internal engines that only expose root-head and recursive-call-head data, rederiving the extracted call graph without a full term conversion, exposing a typeclass-level head-view engine surface, and adding explicit adapters from the first-order and raw-term engine layers (HasCallHeadView, FiniteHeadRuleEngine, HasFiniteHeadRuleView, finiteHeadRuleViewOfRaw, extractedCallGraph)
Meta/DependencyPairs_FiniteCarrierView Consolidated finite-carrier bridge for internal engines whose rules already form a finite type, rederiving the canonical first-order, raw-term, head-view, and extracted-data surfaces from one module (FiniteCarrierFirstOrderEngine, HasFiniteCarrierFirstOrderView, HasFiniteCarrierRawFirstOrderView, FiniteCarrierHeadEngine, HasFiniteCarrierHeadView, FiniteCarrierExtractedEngine, HasFiniteCarrierExtractedView, finiteCarrierExtractedViewOfHeadCarrier)
Meta/DependencyPairs_KernelFirstOrder Internal KO7-symbol frontend for the generic first-order extractor, packaging the eight full KO7 rules over an internal symbol type, then exposing them through HasFiniteRawFirstOrderView, deriving the smaller head-view engine through the generic adapter layer, and feeding the same system through the consolidated finite-carrier stack before rederiving the extracted-node / call-graph surfaces without TPDB syntax (Symbol, EngineTag, RuleId, ko7Engine, ko7HeadEngine, ko7CarrierFirstOrderEngine, ko7CarrierHeadOnlyEngine, ko7CarrierExtractedGraph, ko7FullStepExtractedCallGraph)
Meta/DependencyPairs_Works Extracted DP pair; projection rank (wf_DPPairRev)
Meta/DP_BaseOrder_Boundary Boundary theorem showing the extracted DP pair already admits a simple linear polynomial-style base order (counterHeight_orients_dpPair,
no_blanket_dp_poly_base_barrier)
[…]
Meta/DM_OrderType DM-to-ordinal embedding; $\varepsilon_{0}$ bridge (Thm. 5.9; dmOrdEmbed_lt_opow_omega)
[…]
Meta/Mu3c_Image_LowerBound Explicit realized lower-bound families for the full triple-lex image in both $\delta$-blocks (mu3c_lower_family_block0,
mu3c_lower_family_block1)
Meta/LinearRec_Ablation Linear recursor; simpleSize orientation (§4.6; wf_linearStep)
Meta/MPO_FullStep KO7-specialized MPO: per-rule orientation; reverse-MPO WF (wf_StepRev_mpo)
Meta/MPO_ProofTheoreticBound Explicit ordinal envelope for the KO7-specialized MPO ranking (mpoOrd_lt_mpoVeblenBound, wf_StepRev_mpo_below_veblen7)
Meta/MPO_Precedence_Barrier Explicit bad-precedence obstruction for the KO7 MPO shape (not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec)
Meta/PolyInterpretation_FullStep Nonlinear polynomial $W$; axiom-violation witnesses (Rem. 4.21;
wf_StepRev_poly)
A.3 Relation/property status matrix
Table 13: Certified status matrix for the main relations and transformed systems.
Relation / layer Termination Confluence / normalization Complexity / ordinal Key Lean anchors
Step Root termination: yes Root confluence: no (witness); normalizer: n/a nonlinear polynomial and MPO witnesses; MPO ranking bounded above by $\varphi_{7}(0)$ wf_StepRev_poly, wf_StepRev_mpo, wf_StepRev_mpo_below_veblen7, not_localJoinStep_eqW_refl
SafeStep Strong normalization: yes Root confluence + certified normalizer + reachability to safe normal-form targets triple-lex ordinal upper-bound calibration, explicit two-block lower families in the realized mu3c image, explicit linear root-normalizer lower/upper envelope, plus the root cichon corollary wf_SafeStepRev_c, confluentSafe, normalizeSafe_sound, reachability_decidable, mu3c_lower_family_block0, mu3c_lower_family_block1, normalizeSafeSteps_le_linear_termSize, normalizeSafeSteps_le_mwRootBound, safeMeasure_below_epsilon0
SafeStepCtx Partial-context SN: yes Confluence: yes; Newman characterization retained single-exponential structural bound with matching explicit lower family, retained tower / fast-growing corollaries, conservative calibrated cichon lift, and direct exact contextual cichon lift on a separate note family wf_SafeStepCtxRev, localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, safeStepCtx_length_le_contextualExpBound, safestep_length_bounded_by_size, safestep_length_bounded_by_fgOmegaEnvelope, safestepctx_length_le_mwCtxBound, safestepctx_length_le_ctxExactBound
StepCtxFull Full-context SN: yes Confluence: not claimed; normalizer: n/a explicit $W$-based length bound wf_StepCtxFullRev_poly, stepCtxFullPow_length_le_W
DP transform Extracted DP problem terminates n/a projection-rank descent only dp_projection_orients_rec_succ, dp_projection_violates_sensitivity
A.4 Claim-to-code index
Tables LABEL:tab:claim-code-barrier and LABEL:tab:claim-code-certification index the numbered paper results by Lean identifier, abstraction level, and side conditions. This complements the module map: the tables below are organized by paper statement rather than by file. When a paper result is backed by a bridge theorem together with a family of repetitive class-specific corollaries, the tables list the bridge and the primary anchors rather than exhaustively enumerating every specialization.
Table 14: Barrier, escape, and frontier results indexed by paper label.
Paper label Lean identifier(s) Level Conditions
Thm. 4.1 no_additive_orients_dup_step, no_compositional_orients_dup_step_transparent_succ Schema Tier 1: positive wrap weight. Tier 2: wrap-subterm axioms and transparency at base.
Thm. 4.2 no_affine_orients_dup_step_of_unbounded, no_affine_orients_dup_step_of_succ_pump, no_affine_orients_dup_step_of_wrap_pump Schema Unbounded range, or certified successor / wrap pump.
Thm. 4.3 no_quadratic_counter_orients_dup_step_of_unbounded Schema Unbounded range, or certified successor / wrap pump.
Thm. 4.4 CrossTermBoundedAtBase, no_cross_quadratic_orients_dup_step_of_unbounded Schema Unbounded range plus bounded base coupling, or certified successor / wrap pump.
Thm. 4.5 MultilinearDominatedAtBase, no_multilinear_orients_dup_step_of_unbounded, no_global_step_orientation_multilinear_of_succ_pump Schema / KO7 Finite multilinear monomial table, unbounded range, and frozen pumped-variable coefficient dominated at the base point.
Thm. 4.6 no_polynomial_orients_dup_step_of_unbounded, no_global_step_orientation_polynomial_of_succ_pump,
polynomial_escape_requires_failure_of_base_dominance
Schema / KO7 Finite degree-bounded polynomial table with arbitrary exponents; unbounded or pumped growth; eventual frozen target dominance at the base point.
Cor. 4.7 no_wpoPolynomialDirect_orients_dup_step_of_unbounded, no_global_step_orientation_wpoPolynomialDirect_of_succ_pump,
wpoPolynomialDirect_escape_requires_failure_of_base_dominance
Schema / KO7 interface corollary A direct order whose strict comparison is sound for the bounded polynomial algebra of Thm. 4.6.
Thm. 4.8 no_max_orients_dup_step_of_unbounded, no_global_step_orientation_max_of_succ_pump Schema / KO7 Max-plus constructor-local interface with unbounded range, or positive successor / left-wrapper drift.
Cor. 4.9 no_arctic_primary_orients_dup_step_of_unbounded, no_global_step_orientation_arctic_primary_of_unbounded Schema / KO7 Distinguished finite arctic primary projection satisfying the max-plus interface.
Thm. 4.10 no_matrixD_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with one tracked affine coordinate and growth in that coordinate.
Thm. 4.11 no_matrix2_lex_orients_dup_step_of_unbounded_primary Schema / KO7 Lexicographic order by the first component; unbounded or pumped primary coordinate.
Thm. 4.12 no_matrixMix2_orients_dup_step_of_sum_pump Schema / KO7 Balanced mixed $2\times 2$ constructor maps and an unbounded aggregate-sum pump.
Thm. 4.13 no_matrixFunctional_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with a fixed nonzero natural-weighted scalar projection satisfying the affine interface.
Thm. 4.14 no_orients_dup_step_of_scalar_projection,
no_matrix2_orients_dup_step_of_componentwise_pump_via_projection,
no_matrixMix2_orients_dup_step_of_sum_pump_via_projection
Schema meta-layer A codomain order whose strict decrease forces strict decrease of an already-blocked scalar projection.
Thm. 4.15 not_orients_dup_rule, no_symbolic_variable_condition_orients_dup_step Symbolic schema Standard variable condition on the strict symbolic comparator.
Cor. 4.16 no_kbo_orients_ko7_rec_succ, no_kbo_orients_ko7_rec_succ_trace (paper-facing renaming layer over Meta/SymbolicComparatorBarrier.lean; substantive proof: not_orients_dup_rule) KO7 symbolic corollary KBO-style order satisfying the standard variable condition.
Cor. 4.17 no_affine_with_pump_orients_dup_step,
no_quadratic_with_pump_orients_dup_step,
no_cross_quadratic_with_pump_orients_dup_step,
no_matrix2_with_primary_pump_componentwise_orients_dup_step,
no_matrix2_with_primary_pump_lex_orients_dup_step
Strengthened subclasses Unconditional inside the subclass definitions.
Cor. 4.18 stepCtxFull_orientation_implies_root, no_stepCtxFull_orientation_additive_compositional,
no_stepCtxFull_orientation_compositional_transparent_delta,
no_stepCtxFull_orientation_affine_with_pump, no_stepCtxFull_orientation_quadratic_with_pump,
no_stepCtxFull_orientation_cross_quadratic_with_pump, no_stepCtxFull_orientation_multilinear_with_pump,
no_stepCtxFull_orientation_max_with_pump, no_stepCtxFull_orientation_matrix2_with_primary_pump,
no_stepCtxFull_orientation_matrix2_lex_with_primary_pump,
no_stepCtxFull_orientation_matrixFunctional_with_projected_affine_pump,
no_stepCtxFull_orientation_matrixMix2_with_sum_pump
Context-closed corollary layer Identity-context embedding of root steps into StepCtxFull.
Thm. 4.19 no_additive_compositional_orients_rec_succ, no_compositional_orients_rec_succ_transparent_delta KO7 Tier 1: positive app weight. Tier 2: app-subterm axioms and delta-transparency at void.
[…]
Thm. 4.22 dp_projection_orients_rec_succ, dp_projection_violates_sensitivity KO7 / DP None.
Prop. 4.23 counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order,
no_blanket_dp_poly_base_barrier
KO7 / DP Extracted DP pair problem; linear polynomial-style counter-height base order.
Thm. 4.24 alternating_dup2_realized,
no_additive_orients_alternating_dup2_composite,
no_affine_orients_alternating_dup2_composite_of_unbounded,
cycle_realized,
no_global_orients_ctx_affine_of_unbounded
Alternating SCC schema Additive branch: positive wrap weight inherited from Thm. 4.1. Affine branch: positive wrap weight + unbounded pump.
Thm. 4.25 step_preserves_payloadCount,
synchronized_cycle_realized,
synchronized_cycle_exposes_payload,
no_additive_orients_synchronized_cycle,
no_affine_orients_synchronized_cycle_of_wrapper_dominance,
no_global_orients_ctx_affine_of_wrapper_dominance
Synchronized SCC Additive: synchronized tagged payload inputs. Affine: wrapper-dominance condition + unbounded pump.
Cor. 4.26 no_global_orients_additive, no_global_orients_compositional_transparent_succ, no_global_orients_affine_of_unbounded Schema / global Affine branch inherits the unbounded-pump condition.
Prop. 4.27 W_orients_step, wf_StepRev_poly KO7 root Step None.
[…]
Prop. 4.29 mpoOrd_lt_mpoVeblenBound, wf_StepRev_mpo_below_veblen7 KO7 root Step Fixed signature; ordinal envelope below Ordinal.veblen70.
Prop. 4.30 not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec KO7 root Step Bad-precedence hypothesis.
Thm. 4.31 ko7_direct_escape_trichotomy_extended (preliminary: ko7_nat_direct_escape_trichotomy; projection extension: ko7_projection_escape_trichotomy) KO7 direct universe Successful orienter must violate wrapper sensitivity, violate transparency, or leave the twelve explicit representable families.
Prop. 4.32 not_localJoinStep_eqW_refl KO7 root Step None.
Table 15: Certification and ordinal-calibration results indexed by paper label.
Paper label Lean identifier(s) Level Conditions
Thm. 5.1 measure_decreases_safe_c SafeStep Guarded root step.
Cor. 5.2 wf_SafeStepRev_c SafeStep None.
Prop. 5.3 localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, newman_safeCtx SafeStepCtx Strong normalization already available; the contextual local-join obligation is discharged, yielding unconditional confluence.
Thm. 5.4 wf_StepCtxFullRev_poly StepCtxFull None.
Prop. 5.5 stepCtxFullPow_length_le_W, stepCtxFullPow_length_lt_W StepCtxFull Exact-length contextual reduction hypothesis.
Prop. 5.14 ctxDupPotential, ctxDupPotential_decreases_ctx,
copyBudget_le_termSize, ctxDupPotential_add_one_le_two_pow_double_termSize,
safeStepCtx_length_le_two_pow_double_termSize, contextualExpBound,
contextualExpBound_le_complexity_bound, safeStepCtx_length_le_contextualExpBound
SafeStepCtx Single-exponential structural bound via the position-aware multiplicity potential ctxDupPotential; the older ctxFuel/towerBound estimate survives only as a coarser corollary.
Prop. 5.15 ctxLowerFamily, ctxLowerNF, ctxLowerLen,
termSize_ctxLowerFamily, termSize_ctxLowerNF, ctxLowerFamily_pow_nf,
two_pow_le_ctxLowerLen, safeStepCtx_has_singleExponential_lower_family,
safeStepCtx_has_singleExponential_output_family
SafeStepCtx Explicit linear-size contextual lower family with exponentially long exact chains and exponential normal-form output size, matching the new guarded-context upper theorem up to constants in the exponent and giving a direct duplication-growth witness.
Prop. 5.16 fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope,
safestep_length_bounded_by_fgOmegaEnvelope
SafeStepCtx Explicit finite fast-growing-family envelope majorizing the older coarse contextual tower bound.
Prop. 5.17 mwCtxNote, mwCtxBound, mwCtxNote_lt_opow_omega_mul_two,
ctxFuel_le_mwCtxBound, safestepctx_length_le_mwCtxBound,
not_all_safeStepCtx_rootNote_drop
SafeStepCtx Conservative ctxFuel-to-lex3Note lift; direct contextual reuse of the root note is explicitly obstructed.
Prop. 5.18 ctxExactNote, ctxExactBound, ctxExactNote_lt_omega,
safeStepCtx_exact_step, safeStepCtxPow_exact_drop,
safestepctx_length_le_ctxExactBound
SafeStepCtx Direct exact contextual fundamental-sequence extraction on a separate finite note family built from ctxFuel.
Prop. 5.6 mergeVoidChain_pow_void, normalizeSafe_mergeVoidChain, normalizeSafeSteps_mergeVoidChain SafeStep normalizer Merge-spine family.
Prop. 5.7 tau_add_two_le_two_mul_termSize, normalizeSafeSteps_realized, normalizeSafeSteps_le_linear_termSize,
normalizeSafeSteps_le_ctxFuel, normalizeSafeSteps_le_complexity_bound, normalizeSafeSteps_le_mwRootBound,
termSize_mergeVoidChain, normalizeSafeSteps_has_linear_lower_family, normalizeSafeSteps_has_linear_size_family
SafeStep normalizer Exact guarded root path; linear size bound via $\tau$, plus coarser bounds factored through ctxFuel / complexity_bound and the root mwRootBound.
Reachability decision layer reachabilityDecisionCost_le_linear_termSize, reachabilityDecisionCost_le_complexity_bound,
reachabilityDecision_has_linear_lower_family, reachabilityDecision_has_linear_size_family
Guarded decision-cost envelope Safe normal-form targets; runtime modeled as root normalization plus one equality check.
Thm. 5.8 localJoin_all_safe, confluentSafe SafeStep None.
Thm. 5.9 dmOrdEmbed_lt_opow_omega, safeMeasure_strictMono_explicit, safeMeasure_below_epsilon0 Ordinal calibration Guarded step for the strict-decrease clause.
Thm. 5.11 phaseB_cnf_scaffold_exact_order_type, phaseB_cnf_scaffold_cofinal,
dmOrdEmbed_eq_dmRankOrd, lex3c_order_type_bound
DM / ordinal scaffold None.
Prop. 5.12 payloadTower_mu3c_lower, payloadTower_mu3c_lt_block_cap,
flaggedTower_mu3c_lower, flaggedTower_mu3c_lt_two_block_cap,
mu3c_lower_family_block0, mu3c_lower_family_block1
Full mu3c image Explicit realized families only; no exact order-type claim for the full trace image.
A.5 Local-join report (confluence)
Table 16: Local-join witnesses for the safe relation. Table LABEL:tab:local-join-report collapses the symmetric merge void void instance to the worked concrete case localJoin_merge_void_void; the parameterized left/right family appears earlier in Table 8.
Type Shape Join lemma name Notes
Root merge void void localJoin_merge_void_void both branches to void
Root merge t t localJoin_merge_tt joins at $t$
Root integrate (delta t) localJoin_int_delta unique target void
[…]
Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
Root eqW a a, $\kappa^{M}(a)\neq 0$ localJoin_eqW_refl_guard_ne vacuous join
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root merge
Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
Ctx eqW a a, $\kappa^{M}(a)\neq 0$ localJoin_ctx_eqW_refl_guard_ne vacuous join
Ctx eqW a a (norm + guards) localJoin_ctx_eqW_refl_if_normalizes_to_delta via normalizeSafe